A Complete Bibliography of Publications in Isis, 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.06

Title word cross-reference

1 [310, 411, 940]. 1249 [239]. 1279 [51]. 1299 [557]. 1399 [113]. 13th [405]. 14436 [753]. 1494 [853]. 14th [911]. 1577 [617]. 1665 [17]. 1727 [41]. 1766 [613]. 1780 [307]. 1783 [125]. 1801 [726]. 1842 [228]. 1850 [208]. 1853 [780]. 1860 [520]. 18e [42]. 1900 [737]. 1915 [594]. 1920 [22]. 1926 [92]. 1938 [32]. 1939 [44, 516]. 1941 [316]. 1944 [732]. 1947 [878]. 1949 [958]. 1950 [959]. 19e [42]. 1a [516].

2 [940].

44 [364]. 449 [537].

7 [459]. 75th [878].

=Thomas [649]. âge [265, 754].

A.D [537]. A.D. [239, 51, 113]. Abano [278]. Abarbanel [15]. Abdulhak [85, 925]. Abel [76, 864, 531]. Abendland [883]. Abi [183]. Abraham [63, 484, 13, 652, 848, 539]. Abrahamismus [849]. absolu [34]. Absorption [425]. Academic [460]. Académie [26]. Academy [688]. Acclimatization [942]. according [884]. Account [835]. Accounts [726]. Achievement [285]. Acre [671]. Act [38]. action [225]. Adams [96, 60]. Additional [806]. Addresses [97]. Adelmann [450]. Adivar [730, 925]. Administrative [236, 493, 294, 465, 553, 623, 699, 831, 900, 174, 260, 276]. Admiral [337]. Adnan [730, 85, 925]. Adnan-adivar [925]. Adolphe [611]. Advance [377]. Advances [768]. Adventures [736, 210]. Adventurous [549]. Adversaries [750]. Advocate [222]. Aeronautics [639]. Aesch [338]. Aesculapius [575]. African [349]. after [408]. Again [284, 603]. against [152, 122]. Agassiz [129]. Age [607, 234, 745, 419, 868, 395, 490, 376]. Ages [789, 265, 754, 346]. Aggression [942]. Agricultural [597]. Agriculture [364, 171, 457, 649, 772, 671]. Agus [809]. Aimé [423]. Airy [501]. Ak [119]. Akademien [153]. Akkoyunlu [119]. al [184, 592, 711, 469, 50, 184, 362, 362]. al-‘arab [592]. al-din [184]. Al-Fadil [469]. al-firasa [184]. al-Haitham [362]. Al-Hasan [362]. al-‘ilmi [592]. Al-Kāshī’s [711]. al-Razi [184]. Albert [102, 347, 919]. Alberti [53]. Albertus [111, 540]. Albright [234, 454]. Alchemical [140, 540, 756]. Alchemy [713, 790]. alchimiques [93]. Aldo [651]. Alexander [941, 663, 338, 764, 500, 746]. Alexandre [648]. Alexandria [751]. Alfonso [341, 729]. Alfred [23, 887, 103, 574, 254]. alia [198]. Alice [513]. Alimentary [244, 331]. Alkimia [111]. Allen [681, 366, 396, 416]. Allergy [630]. Allgemeiner [287]. Allott [409]. Allyne [811]. Almageste [611]. almanach [635]. Alone [782]. Alpert [130]. Alphabet [882]. Alwin [142]. Amand [602]. Amateur [497]. Amé [430]. Amelia [20]. America [208, 940, 687, 500, 639, 569, 575, 682, 171, 670]. American [676, 32, 726, 170, 95, 102, 251, 594, 316, 419, 673, 233, 407, 649, 671, 505, 472, 385, 606, 471, 218, 829, 653]. Amerindian [681]. among [730, 241]. Amos [478, 228]. Amsterdam [959]. Anadolu [119]. Analysis [146]. Analytical [375]. Anatolian [119]. Anatomicae [721]. Anatomy [775, 672, 458]. Ancestry [170]. ancienne [200]. anciens [754]. Ancient [750, 39, 844, 196, 197, 560, 634, 907, 866, 241, 472, 754, 920]. Ancients [179]. Andean [942]. Anderson [209]. Andes [942]. Andreas [545]. Andrew [407, 590, 653]. Anecdota [198]. Anecdotal [138]. Anesthesia [737, 679]. Angelo [760]. Angie [371]. anglais [66]. Anglicus [441]. Angus [810]. animalium [146]. Animals [508]. Ann [212]. Anniversary [878]. Anno [516]. annotées [652]. Annual [213]. Anquetil [159]. Anquetil-Duperron [159]. Anshen [870]. Answers [226, 641, 836, 403, 181, 527, 506, 543, 564, 304, 162, 446, 715, 359, 475]. Antarctica [95]. ante [110]. Antécédents [62]. antérieurs [93].

Anthemius [273]. Anthropological [102]. Anthropology [893, 251]. antico [621]. Antilles [14]. Antiquity [57, 6, 658]. Antoine [406]. Anton [863]. Antoni [124]. Antonina [64]. Antonio [451]. Antony [157]. Ape [57]. Apollonia [845]. Apology [292]. Appeal [825, 843]. Appendice [760]. Application [904]. Applied [726]. Appreciation [72]. Approach [731]. April [613, 317]. Aquapendente [450]. arab [592]. arabe [184]. arabes [62, 309]. arabi [621]. Arabia [341]. Arabic [587, 112]. Arabs [592, 486]. Aramaic [263]. Arang [77]. Archaeology [454]. Archer [92]. Archibald [370, 32, 929]. Archimède [709]. Archimedes [299, 264]. Architectural [751]. Architecture [274, 370]. Arctic [94, 147]. Aristarchos [264]. Aristarchus [299]. Aristides [575]. Aristotelian [31, 581]. Aristotelianism [55]. Aristotle [775, 672, 508]. Aristotle. [61]. Arithmetic [637]. Armand [308, 198, 865]. Armando [722]. Armitage [810]. Army [312]. Arranged [884]. Ars [892]. Art [875, 825, 510, 678, 778]. Arthur [380, 523, 491, 20, 146, 169, 881, 197, 135, 574, 657, 453, 361]. Article [947]. artificum [189]. Arts [922, 934]. Artur [549]. Arturo [736]. Arzt [104]. Aschoff [144]. Asclepius. [660]. Ascribed [111]. Ashley [313, 433, 732, 458]. Ashton [646]. Asin [252]. Aspect [180]. Aspects [789, 72, 872, 943]. Association [673]. Astāṅgahrdayasamhita [340]. Astrolabe [907, 135]. astrolabes [303]. Astrological [63, 539]. Astrology [416, 875, 875]. astronom [125]. Astronomer [177, 848, 810]. astronómica [848]. Astronomical [904, 300, 752]. astronomischen [616]. Astronomy [560, 634, 907, 592, 526, 322, 13]. Athanasius [834]. Athelstan [859]. atheniensia [198]. Atlantica [165]. Atomic [869, 330, 281]. Atomists [748]. Attempt [638]. Attraction [66]. Attributed [335, 540]. Aufklärung [307]. August [959, 732, 537]. Auguste [127]. Augustin [34, 35]. Augustus [920]. Aurel [375]. Ausbreitung [616]. autenticidad [848]. authenticity [848]. Authorized [591]. Authors [293]. Autobiographical [650]. Autrecourt [911]. Auzels [663]. Avenezra [539]. Avery [249]. Aviation [943]. Avicennian [224]. Avoidance [326]. Aydin [926]. Ayurveda [340].

B [91, 768, 731, 646, 74, 878, 210, 450, 761, 316, 515, 462, 862, 171]. bâtisseurs [265]. Babel [787]. Babylonian [9, 58, 560, 634, 8]. Bachelard [924]. Back [46, 318, 351, 383, 421, 438, 467, 577, 598, 625, 683, 701, 739, 784, 799, 888, 901, 945, 961]. Background [501, 81, 27]. Bacteriologist [74]. Bacteriology [37]. Badianus [185, 185]. Baer [628]. Baitsell [75]. Baker [827, 517]. Bakr [183]. Ballard [532]. Bandelier [251]. Bandelier-Morgan [251]. Banting [917]. Baptist [267]. Barbara [737]. Barbegal [754]. Barbour [548]. Barclay [120]. Barnes [40]. Barnette [336]. Barney [448]. Barnouw [515]. Barometric [479]. Baron [485, 248]. Barry [397]. Bartram [250, 613, 613]. Barzun [682]. Basel [112]. Basis [934]. Batchelor [572]. Bates [738]. Bateson [256]. Bats [326]. Battista [53]. Baud [166]. Bayle [18]. Beast [255]. Beast-Machine [255]. Beaumont [131]. Becker [40]. Beeckman [762]. Beer [938]. before [596, 649].

Beginning [63, 300]. Beginnings [713]. Behind [463]. Being [132, 20, 721]. Beirut [316]. Beitrag [763, 723]. Belgian [161, 843]. Belgique [180]. Beliefs [867]. Bell [207, 914]. ben [248]. Benedict [215]. Bengt [153]. Benjamin [650, 271, 196, 612, 159, 650, 726, 271]. Bennett [413]. Bentwich [83]. Bequaert [483]. Bernard [549, 271, 569, 197, 376, 816, 868]. Bernardini [189]. Bernardino [189]. Bernhard [940]. Berryman [193]. Bert [479]. Bertrand [291, 367, 657, 779]. Beryl [781]. Berzelius [589, 589]. Best [607]. Bestiary [405]. Betts [649]. between [242, 343]. Bey [362]. Beylikleri [119]. Bhagavad [568]. bi [80]. Bible [485]. Bibliografia [894]. Bibliographic [940]. Bibliographical [53, 3]. Bibliographie [616]. Bibliographies [587]. Bibliography [632, 256, 154, 188, 44, 173, 235, 317, 437, 382, 494, 535, 552, 563, 608, 698, 798, 839, 957, 899, 545, 894, 293, 208]. Biblioteca [452, 621, 118]. Bibliothèque [773]. bibliothèques [93]. Bicentennial [649]. Bidez [25]. Bigelow [168]. Bigger [152]. Billion [671]. Bio [545, 293]. Bio-Bibliography [545, 293]. Biochemistry [455]. Biography [591, 83, 656]. Biological [801, 456, 459, 267, 618, 432]. Biologist [210]. Biology [168]. Birds [394]. Biringuccio [449, 449]. Birth [96]. Birthday [732]. Biruni [306]. Black [41]. Blackley [630]. Blades [772]. Blake [920]. Blasius [636]. Blood [788]. Bloomfield [291]. Boardman [655]. Bobbio [273]. Boerhaave [71, 71]. Boethius [424]. Bohr [291, 823]. Boise [920]. Bonpland [423]. Book [934, 15, 27, 29, 40, 380, 783, 435, 676, 523, 617, 875, 769, 928, 821, 827, 855, 924, 191, 375, 763, 121, 335, 911, 21, 31, 32, 65, 91, 92, 132, 136, 137, 138, 168, 186, 207, 209, 232, 250, 272, 291, 292, 342, 364, 417, 533, 591, 613, 650, 651, 664, 725, 726, 765, 768, 825, 856, 944, 775, 549, 430, 271, 72, 88, 139, 140, 449, 771, 659, 39, 76, 133, 273, 514, 660, 661, 662, 908, 941, 942, 943, 936, 164, 334, 455, 510, 663, 142, 170, 310, 429, 666, 883, 491, 206, 730, 926]. Book [546, 731, 817, 534, 63, 84, 201, 202, 234, 248, 314, 369, 372, 412, 434, 454, 484, 485, 486, 550, 809, 819, 826, 867, 34, 35, 89, 99, 308, 95, 313, 371, 381, 408, 436, 482, 566, 674, 677, 681, 736, 772, 777, 870, 914, 916, 927, 69, 365, 406, 593, 339, 77, 811, 96, 19, 149, 880, 938, 94, 338, 416, 646, 810, 16, 64, 249, 858, 307, 431, 433, 141, 12, 17, 20, 23, 36, 60, 74, 98, 120, 144, 146, 147, 167, 213, 228, 259, 379, 410, 432, 448, 456, 459, 481, 738]. Book [165, 547, 569, 864, 872, 463, 575, 737, 886, 887, 923, 878, 37, 152, 219, 257, 275, 315, 377, 781, 30, 90, 169, 343, 665, 733, 734, 823, 860, 873, 876, 937, 881, 512, 28, 38, 192, 215, 413, 479, 530, 531, 567, 595, 678, 682, 857, 492, 852, 102, 127, 529, 24, 81, 82, 83, 75, 78, 100, 103, 128, 129, 130, 131, 143, 145, 150, 151, 172, 187, 190, 194, 196, 197, 203, 205, 210, 214, 216, 217, 229, 230, 231, 251, 253, 255, 256, 274, 288, 345, 349, 350, 366, 367, 373, 409, 411]. Book [450, 460, 480, 483, 487, 490, 517, 594, 647, 655, 672, 675, 778, 782, 469, 879, 863, 13, 122, 135, 252, 939, 70, 104, 622, 761, 854, 922, 874, 882, 86, 341, 764, 869, 574, 620, 648, 657, 779, 828, 97, 193, 376, 616, 909, 770, 861, 610, 117, 188, 200, 290, 347, 370, 414, 568, 612, 615, 722, 80, 14, 18, 26, 33, 41, 42, 43, 61, 62, 68, 79, 85, 87, 93, 123, 124, 125, 134, 153, 166, 183, 184, 189, 195, 198, 199, 204, 208, 247, 258, 287, 289, 293, 306, 309, 312, 316]. Book [340, 346, 362, 363, 368, 418, 420, 452, 453, 508, 509, 511, 515, 516, 532, 570, 573, 589, 592, 611, 614, 621, 652, 728, 729, 732, 735, 762, 766, 773, 774, 812, 814, 815, 816, 818, 820, 830, 865, 866, 885, 910, 913, 918, 919, 920, 921, 929, 843, 890, 71, 218, 348, 361, 419, 428, 462, 464, 513, 545, 551, 654, 656, 673, 679, 680, 721, 780, 829, 853, 862, 884, 917, 119, 336, 458, 596, 461, 619, 760, 915, 478, 813, 912, 724, 254, 374, 767, 25, 925, 477, 776, 935, 723, 940]. Book [822, 22, 415, 118, 337, 451, 576, 727, 148, 211, 378, 73, 101, 171, 185, 212, 233, 311, 344, 407, 457, 488, 489, 518, 519, 520, 521, 522, 548, 571, 572, 590, 597, 618, 649, 653, 658, 667, 668, 669, 670, 671, 824, 859, 868, 871, 877, 930, 931, 932, 933, 66, 67, 126, 405, 649]. Books [948, 545]. Bosch [854]. Bosmans [843]. Boston [958, 532, 327]. Botanic [223]. Botanik [877]. Botanists [250]. Botany [572, 653, 407]. Bottle [259]. Bouchard [69]. Bowditch [327]. Boyden [202]. Boyer [91]. Boyle [546]. Boyunca [925]. Bradford [614]. Brahe [811]. Brant [530]. Brave [262]. Bread [782]. Breasted [385]. Breeding [572]. Brett [585]. Brev [589]. Brief [45, 105]. Briefwechsel [287]. Brinton [880, 288]. British [677, 464]. Brodbeck [431]. Brodbeck-Sandreuter [431]. Brodie [376]. Broglie [136, 823]. Bronislaw [329]. Brothers [591]. Brown [942, 817]. Browne [649, 126, 597]. Brunetto [910]. Bruno [104, 760, 176]. Bruyne [818]. Bryan [764]. Buch [908]. Buchler [148]. Buess [763, 723]. builders [265]. Bulloch [37]. Burma [483]. Buros [415]. Burr [502]. Burrows [434]. Busbecq [262]. Butler [442]. Byron [869].

C [380, 523, 617, 821, 138, 591, 99, 436, 456, 872, 886, 377, 873, 192, 150, 414, 61, 123, 208, 721, 884, 212, 387]. ca [49]. Cable [567]. Cadiz [14]. Caillet [547]. Calculus [91]. Calendar [41]. California [36]. called [663]. Calvin [706]. Cambridge [256, 411]. Cameron [416, 370]. Camille [531]. Campbell [97]. Canada [140, 290]. Canby [944]. Cannon [614]. Canonicus [953]. Cantera [63]. Capitolare [118]. Cardiology [462]. Care [940]. Cari [448]. Carl [91, 364, 650, 499]. Carlo [341, 729]. Carlos [942]. Carmody [910]. Carnap [291, 374]. Carnegie [88, 10]. Carnot [365]. Carolina [746]. Carolinas [613]. Carre [19]. Carroll [99]. cartes [773]. Cartographie [773]. Case [367]. Cases [41]. Casey [510]. Cassadors [663]. Cassius [769]. Casson [214]. Castiglioni [736]. Catalog [269]. Catalogue [532, 140, 93, 293, 773]. Catedral [452]. Cattell [432]. Caullery [98]. Cave [189]. Cayenne [324]. Cecilia [884]. Celebrated [443]. Celebration [878, 649]. Celestial [375, 538]. Cell [504, 629]. Cellular [504]. Census [290]. cent [44]. Centuries [771, 42, 843]. Century [121, 911, 725, 817, 16, 753, 293, 654, 405, 393, 266, 638, 394, 600, 21, 342, 775, 914, 852, 672]. Cerebrum [190]. Ch [209]. Challenge [886]. Change [767]. Chapter [186, 630]. Charges [330, 281]. Charlatan [151]. Charles [380, 855, 291, 364, 38, 192, 253, 573, 551, 721, 451, 489, 597, 649, 824, 280, 630, 442, 817, 148]. Charles-Édouard [817]. Charles-Edward [551]. Chase [661]. chasse [663]. Chemical [922]. Chemie [666]. Chemist [170]. Chemistry [168, 310, 400, 948, 139, 597]. Chemists [771]. Cheney [89]. Cherbury [19]. chez [127, 754, 85, 66]. chi [835]. Ch’iao [922]. Ch’iao-P’ing [922]. Chiashêng [921]. chiens [663]. Child [170]. chimiste [69, 430]. Chin [549]. China [426, 586, 695, 850, 922, 866]. Chinese [199, 453, 921, 51, 601, 756]. Ch’ing [453, 835, 835]. Christian [914]. Christianity [234]. Christopher [734, 14, 337, 126, 55]. Chromatography [584]. Chronology [109, 263, 940]. church [265]. Churchman [873]. Chypre [117]. Científico [727]. Cifris [53]. Cincinnati [688]. Circulation [240]. Circumference [445]. Circumstances [721]. Citizen [29]. Citrus [572]. City [938]. Civil [596, 401]. Civilization [154, 44, 173, 235, 317, 437, 382, 494, 535, 552, 563, 608, 698, 798, 839, 957, 899, 870, 923, 150, 879, 346, 534]. Clagett [334]. Clapesattle [428]. Clapham [411]. Clara [192]. Clare [32, 929]. Clark [149, 451]. Clarke [523]. Clarkson [436]. Classical [244, 331, 805, 658]. Classics [195]. classique [34]. Claudius [293]. Clement [549]. Cleves [694]. Clifford [767]. clima [272]. Climatic [942]. Clock [634]. Clue [769]. Co [753]. Co-ordinates [753]. Cochran [436]. Code [202]. Codex [118]. Codicum [728]. Coffee [932]. Cohen [271, 255, 918, 868]. Cohn [887]. Colbert [483]. Cole [775, 672]. Collected [124]. Collection [660, 545, 327, 368]. Collis [482]. Color [425, 664]. Colours [209]. Columbus [14, 337]. Columns [298]. comacini [265]. Combe [854]. Comenius [478, 813]. Comet [617, 806, 557]. Commentaires [908, 611]. Commentariolus [122]. Commentary [412]. commentateur [847]. commentateurs [66]. commentator [847]. Commentators [909]. Comments [743]. Committee [43]. Community [435, 435, 485]. Como [265]. Compagnie [773]. Company [677]. Comparative [775, 672, 458]. Comparison [301]. Compendium [280, 451]. complètes [67]. Compositiones [118]. comprenant [652]. Comprising [726]. Computation [526]. Computus [636]. Comte [127]. con [760]. Conant [731, 477]. Concentric [243, 694]. Concept [474, 267]. Concepts [91]. Concerning [282, 953]. Concise [928]. Concordiae [249]. Conduit [931]. Conference [878]. Configuration [674]. Confirmations [942]. Confiteor [580]. Conflicting [9]. Congress [959, 269]. Conjunctions [711]. Connected [390]. Conqueror [336]. Conquest [147, 551]. Conservation [744]. constructeurs [303]. Construction [920]. constructors [303]. Consultative [43]. Containing [749]. Contents [45, 105]. Contribution [753, 529, 515, 570]. Control [179, 869]. conventionnel [69]. Cook [197]. Coolidge [232]. Cooperation [88]. Copernic [356]. Copernican [122]. Copernicus [810, 122, 710, 264, 646]. coppernicanischen [616]. Coptic [691]. Corbett [93]. Cordial [443]. Cornelis [762, 812]. Corps [312]. Correlative [884]. Correspondance [134, 652, 812]. Correspondence [11, 182, 333, 544, 163, 391, 837, 896, 404, 528, 476, 565, 609, 227, 246, 270, 286, 447, 717, 59, 115, 588, 696, 757, 507, 305, 642, 793, 954, 427, 360]. Cortes [727]. Cortesão [722]. Cosmic [138]. cosmologiques [562].

Cosmology [176]. Cotton [426, 188, 694, 188]. Coulter [590]. Courant [533]. courants [874]. Course [925]. Courtines [18]. Cow [242]. Craig [724]. Crane [880, 288]. Creswell [370]. Crew [186]. Crime [102]. Crimes [941]. Criminal [102]. Crisis [490]. Critical [949, 154, 44, 173, 235, 317, 437, 382, 494, 535, 552, 563, 608, 698, 798, 839, 957, 899, 389]. Critique [570, 774]. Crop [179]. Cross [242]. Crossed [416]. Crowther [231]. Crucible [406]. Culture [769, 674, 133, 514, 662]. Culturhistorischer [877]. cum [266]. Cummings [233]. Cumont [25]. Cuneiform [659]. Cushing [545, 656, 545]. Cylinder [78]. Cyril [449].

D [817, 436, 654]. D’Abro [137]. Dainville [936]. d’Alexandrie [611, 611]. dalla [516]. Dalmatius [603]. Dangerous [433, 513]. Daniel [916, 915]. dans [529, 200, 68, 935]. Dantzig [33]. d’après [773, 25]. D’Arcy [195, 766, 378]. d’Aristote [164]. Darwin [223, 630, 631]. Darwinism [594, 229]. d’astrolabes [303]. Data [690, 940]. Date [893, 110]. Dated [906]. Daude [663]. Dauzat [347]. David [149, 28, 882, 764, 123, 73, 519, 666]. Davies [573]. Davis [310]. Dawn [504]. Dawson [96]. Daylight [526]. Dayton [138]. Deasmumhan [771]. Death [406]. Debo [371]. Decades [653]. December [958, 494, 608]. Decline [137, 55]. Deductive [254]. Defence [440]. Defense [890]. Defries [20]. Degenerate [394]. degli [516]. d’églises [265]. degrees [516]. del [252, 621, 760]. Delatte [198, 865]. della [209, 249, 621]. Dels [663]. demain [874]. Deman [920]. Demands [725]. Democritus [845]. Demonstration [240]. Denis [788]. Denisovich [668]. Denny [407, 590, 653]. Dental [940]. Dentistry [940]. Department [772, 457]. Descartes [209, 498, 648]. descripsit [728]. Description [451]. Desert [567]. Design [51]. Desmond [862]. dessins [166]. d’esthétique [818]. Destiny [381, 681, 218]. Destombes [773]. Determination [5]. DeTurk [933]. d’Euclide [908]. Deutchland [616]. Deux [919]. Development [390, 503, 207, 639, 942, 586, 914, 96, 737, 619]. Developments [772]. Deventer [814]. Devletleri [119]. Dewey [291, 367]. Dexter [572]. d’Hiéroklès [117]. d’Hippocrate [864]. Dhirendra [340]. Diadochus [908]. Diagnosis [6]. Diary [613]. diatriba [189]. Dichtung [104]. Dickes [64]. Diction [114]. Diderot [178]. Didyme [273]. Dies [134]. Diffusion [413]. Dignity [827]. Dijksterhuis [912]. Din [925, 184]. Dingler [90]. Diogenes [845]. Diophane [273]. Diringer [882]. Dirk [928, 857]. Disappearance [371]. Disciple [746]. Discours [26]. Discoveries [209, 95]. Discovery [283, 905, 214, 14, 141]. Discussion [721]. Disease [534, 551]. Diseases [189]. Divine [856, 756]. Division [10]. Djin [549]. Do [607]. Dobzhansky [432, 344, 668]. Docks [600]. Doctor [420, 814, 853]. Doctors [218, 428, 941]. documenti [760]. Documents [236, 493, 294, 465, 553, 623, 699, 831, 900, 174, 260, 276, 783, 95]. Dodo [930]. Doe [492]. Don [416]. Donald [942, 733]. Doren [650]. Doris [617]. Dorothy [410, 939, 853]. dou [910]. Douglas [123, 859]. Dr [613, 131, 443].

Dr. [54]. Drabkin [918]. Drake [915]. Drinking [932]. Drinks [468]. Drugs [463]. Drummond [212]. Dublin [23]. Ducassé [127]. Dudley [726]. Duhem [308]. Dulles [249]. d’un [430]. Duncan [854]. Duncum [737]. Duperron [159]. Dupre [365]. Durham [177]. Durkheim [130]. Dury [813]. Dutrochet [629].

ea [272]. Earliest [749, 444, 633]. Early [396, 160, 690, 726, 586, 630, 948, 483, 907, 601, 742, 780, 387, 647, 370]. Earnest [102, 250]. Earth [844, 445]. East [677, 677, 519]. Eastern [402]. Eaton [228]. ébauchés [166]. eclipses [560]. Economic [339, 216, 411, 869]. Eddington [169]. Edelstein [660, 660]. Eder [664]. Edgar [523, 312, 818]. Edgerton [568]. Edib [84]. Editi [729, 341]. Editing [643]. Edmund [209, 194, 694]. Édouard [817]. Eduard [307, 87]. Education [43, 596, 935, 43, 936]. Educational [783]. Educator [228]. Edward [292, 664, 202, 89, 149, 219, 647, 122, 199, 511, 735, 551, 410, 458]. Edwards [856, 649]. Edwin [530, 483, 649]. Eecke [273]. eeuw [42]. Effect [681]. Egerton [549]. Egleston [289]. Egypt [7]. Egyptology [385]. eight [703]. Eighteen [906]. Eighteenth [21, 725, 775, 672, 42, 394]. Eighteenth-Century [394]. Eighth [833, 44, 563]. Eileen [411]. Einstein [765]. Eisler [875]. Elaine [566]. Electricity [138]. Electrolysis [605]. Elementen [908]. Elements [141, 908]. Elenco [621]. Elephantine [263]. Eleventh [753]. Eliot [825, 14, 337]. Elizabeth [726, 920]. Ellery [448]. Ellipse [710]. Ellis [128]. Elmer [40, 22]. Elvira [141]. Embryological [450]. embryologie [98]. Embryology [146]. Emergent [870]. Emerson [36]. Emery [826]. Emil [209, 776]. Emile [547, 130]. Emily [185]. Eminent [453]. Emma [660]. Emmanuel [190, 570, 774]. Emmart [185]. Emmet [915]. Empire [677, 166]. Empiricism [148]. Encyclopedia [291, 875]. End [161]. Ende [773]. Endocrinologist [817]. Energy [744, 869]. Engineering [120]. Engineers [120]. England [18, 387, 416, 16, 458]. English [229, 256, 405, 637, 18, 123, 824]. Englishman [212]. Enlightenment [855]. Ensayo [727]. enseignements [323]. entre [652]. Entretiens [648]. Entstehung [616]. entwickelt [209]. Entwickelung [877]. Entwicklung [776]. Entwicklungsgeschichte [666]. Epicycles [710]. Epidemic [551]. Epsom [357]. Epstean [664]. Epstein [485, 781]. Equations [58]. Equatorials [51]. Erasmus [223]. Erich [595, 620]. Erik [718]. Erinnerungen [860]. erkenntnispsychologisch [209]. Ernest [250, 291, 170, 117, 22]. Ernst [209, 616, 488, 640]. Erskine [312]. ersten [908]. especial [272]. Espinosa [451]. Esprit [924]. Essai [127]. Essay [301, 894, 879]. Essays [523, 923, 732, 380, 769, 887, 248, 881]. Esther [920, 921]. Estimates [160]. Ethel [228, 854]. Ethiques [164]. Ethnic [676]. Ethnology [358]. Etudes [818]. Euclide [919]. Eudoxus [109]. Eugène [68]. Euklids [908]. Europe [411, 41, 171]. European [251, 750]. Eve [132]. Everett [726, 649]. Evil [661, 890]. Evolution [473, 714, 628, 230, 225, 209, 170, 669, 518]. exakten [307]. Examples [906].

Excursion [930]. Excursus [587]. exemplaire [773]. Exhaustion [401]. Existence [937]. Expedition [324]. Experience [610]. Experimental [240, 285, 873, 330, 654]. expérimentale [98]. Experiments [425, 271]. Explanation [50]. Explorers [250]. Exploring [513]. Explosives [310]. Expression [89]. Extracts [251]. Eyer [677]. Ezo [80, 80]. Ezra [693, 13, 848, 159, 63, 63].

F [523, 821, 65, 168, 209, 291, 726, 942, 943, 510, 234, 454, 313, 408, 433, 456, 256, 672, 874, 41, 293, 532, 732, 866, 656, 458, 877]. F.R.S [458]. F.R.S. [470]. Fêng [921]. Faber [443]. Fabricius [450]. Fadil [469]. Fahrenheit [792]. Fairchild [73, 519]. Fairfield [871]. Faith [855, 81]. Fakhr [184]. falak [592]. Falconry [444, 510]. Fallacy [433, 394]. Falling [622]. Falsely [335]. Far [677]. Faraday [605, 2, 468]. Farber [828]. Farm [671]. Farrington [196]. Fate [661]. Father [385, 496, 843]. fauconnerie [663]. Faÿ [816]. Feminine [442]. Festschrift [431]. Ficino [477]. Field [915]. Fierz [666]. Fierz-David [666]. Fifteen [906]. Fifth [423, 494, 802]. Fifty [154, 44]. Fifty-eighth [44]. Fifty-ninth [154]. Fig [931]. fil [592]. fil-riyadiyyat [592]. Filii [183]. Find [313]. Finkelstein [81]. Finland [194]. Fir [859]. firasa [184]. Firearms [586, 695]. First [5, 726, 637, 235, 798]. Fisch [853]. Fischer [74]. Fish [601]. Fishers [632]. Fishes [284, 766]. Five [23, 878]. Flexner [419, 419]. Floating [600]. Florida [613]. Flying [326]. Folk [82]. fondo [621]. Food [522, 933, 212, 233, 571]. Fools [530]. Form [843, 378]. formación [651]. Formation [924]. formazione [621]. Formulas [400]. Forris [139]. Forscher [87]. Fortune [890]. Forty [353, 890, 947]. Forty-three [353]. Fossil [284, 601]. Fothergill [613]. Foundation [828]. Foundations [375, 665]. Fournival [847]. Fourteenth [852]. Fourth [385, 957, 437]. fragment [273]. fragmentaque [183]. franc [816]. franc-maconnerie [816]. francaise [936, 347]. France [26, 654, 753, 754]. Francesca [567]. Francesco [151]. Francis [775, 60, 612, 910, 921, 559]. Franciscan [636]. Francisco [63]. Franco [558]. Francois [936, 654]. Frank [765, 96, 289, 218, 618, 343, 665, 620]. Frankel [855]. Frankfort [78]. Franklin [191, 650, 271, 568, 159, 425, 271]. Frans [615, 670]. Franz [549, 25]. Fred [591, 941, 884]. Frederick [917, 596, 22, 397]. Freedom [686, 65, 778, 89, 933]. Fremont [576]. French [855, 52, 847, 323, 951, 406, 567, 303, 265, 754]. French/Italian [265]. Freud [949]. Friedenwald [550]. Friedman [664]. Front [1, 47, 107, 175, 220, 237, 261, 277, 296, 297, 320, 352, 384, 422, 439, 495, 524, 536, 555, 556, 579, 599, 626, 685, 702, 741, 786, 800, 841, 842, 889, 902, 946]. Fueter [307, 87]. Fulton [726, 943, 656]. Function [425]. Fundamental [503]. Fungi [377]. Funun [52]. Furnas [150]. Future [576, 655, 571]. Fyfe [510].

G [191, 292, 650, 491, 485, 410, 231, 289, 813]. Gade [811]. Gaetano [172]. Galdston [463, 257]. Galen [610, 278]. Galilean [279]. Galilée [356].

Galileo [389, 65]. Gaon [867]. Garden [649, 746, 73, 223, 519]. Garment [414]. Garnett [126]. Garrett [368]. Gaston [924]. Gauthier [62]. Gay [22]. Gaylord [669]. Geddes [655]. Gegenwart [877, 883]. General [789, 293]. Generation [508]. generatione [146]. Genetics [667, 344]. Genius [191]. Genuine [60]. Geo [222]. Geo-Heliocentric [222]. géographie [936]. géographique [117]. Geography [211]. Geological [96]. géologie [570, 774]. Geometrical [232]. géométrique [919]. Georg [791]. George [765, 167, 923, 219, 38, 852, 75, 349, 366, 669, 859, 585, 502, 222, 56]. Georges [69, 117, 363]. Georgia [613]. Gerald [145, 862]. German [338, 753]. Germs [275]. Gershom [485]. Gervase [497]. Geschichte [142, 209, 307, 616, 770]. geschichtlichen [776]. Ghatrif [663]. Ghosts [587]. Gibbs [726, 429]. Giedion [274]. Gihring [883]. Gilbert [133, 514, 662, 703]. Ginzberg [412]. Giordano [760, 176]. Giorgio [621]. Giovanni [335, 334]. Girolamo [832]. Gita [568]. Glass [468]. Glasse [582]. Gleanings [813]. Gleason [92]. Globe [538]. glossaire [247]. Glossary [766]. Glueck [201]. Gmunden [354]. Gnudi [449]. Gobi [567]. God [149, 414]. Gode [338]. Gode-von [338]. Goichon [224]. Golden [549]. Goldfrank [921]. Goldsmith [394]. Goldstein [143]. Good [661, 890]. Goodman [650, 191]. Gottfried [287]. Governments [119]. Gradation [473, 714]. Grammar [43]. grands [874]. Grant [203]. Graph [690]. Graphic [400]. Gras [171]. Grass [772]. Grassmann [542]. Graubard [522]. gravée [773]. Gravitation [950]. Great [783, 482, 916, 519]. Grèce [76]. gréco [62]. gréco-arabes [62]. grecque [864, 25, 919]. grecs [198, 865]. Greek [133, 514, 661, 662, 167, 766, 918, 767, 51, 703, 748, 278, 268]. Greene [661, 100]. Greenland [418]. Grégoire [708]. Gregory [350]. Grete [151]. Grew [301]. Grinnell [481, 481]. Griswold [86]. Grosse [863, 87]. Ground [733]. Groups [676]. Growth [168, 876, 674, 378]. Grundlagen [763, 776, 723]. Guide [926, 219]. Gunpowder [695]. Gunther [647, 17]. Gustave [346]. Guterman [646]. Guthrie [61]. Guyton [69]. Guyton-Morveau [69]. Gynecology [680].

H [136, 168, 291, 292, 726, 944, 663, 666, 314, 19, 938, 12, 30, 530, 640, 78, 205, 251, 411, 483, 70, 309, 592, 921, 829, 853, 813, 935, 667, 932]. Haak [4]. Haddon [480]. Haitham [362]. Hakan [663]. Hakikî [926]. Hakki [119]. Hakluyt [735]. Haldane [768, 210]. Hale [392]. Halide [84]. Hall [139, 213]. Hallam [483]. Halley [694]. Hallowes [396]. Hamilton [513]. Handbook [492]. Hans [777, 593]. Happiness [313]. Harding [772]. Hardy [292]. Harley [349]. Harness [638]. Harrison [630]. Harry [40, 550, 130]. Hartlib [813, 4, 813]. Hartner [135]. Hartshorne [211]. Harvey [146, 545, 859, 656]. Harveyan [240]. Hasan [362]. Haupt [405]. Havelock [128]. Hayatta [926]. Hayyan [529]. Head [480]. Head-Hunter [480]. Health [23, 830]. Heard [145]. Heat [425]. Heavens [61]. Hebrews [241]. Heidel [361]. Heinrich [763, 459, 816, 723]. Heinz [941]. Helen [428]. Hélène [66]. Heliocentric [222]. Helix [322]. Hell [496]. Hellénises. [25].

Hellenistic [339]. Hellman [617]. Hellmut [483, 405]. Helmont [267, 268, 761]. Hendrik [217, 814]. Henkin [229]. Henri [200, 822]. Henry [380, 186, 944, 534, 84, 56, 761, 258, 346, 385, 830, 419, 862]. Herbal [239, 612]. Herbert [533, 95, 19, 38, 572]. Heredity [668]. Heritage [592]. Herman [71]. Hermann [369, 646]. Hero [851]. Heroic [419]. Heron [110]. Herrick [462]. Herschel [827]. Herskovits [216]. Hesiod [893]. Hicks [199]. Hieronymous [450]. High [43]. Highet [133, 514, 662]. Hildebrand [153]. Hilgenberg [340]. Hill [302, 275]. Hindu [526]. Hingston [480]. Hipolito [272]. Hippocrates [60]. Hippocratic [361]. Hispaniola [14]. histoire [323, 529, 198, 570, 935]. Historia [727]. Historian [500, 704, 606, 172]. Historians [825]. Historical [10, 380, 209, 942, 731, 358, 474, 904, 80]. historisch [209]. Historischen [763, 723]. historischer [287]. History [396, 10, 958, 959, 380, 928, 21, 32, 138, 168, 209, 232, 342, 944, 549, 139, 444, 372, 772, 898, 339, 399, 632, 705, 753, 923, 37, 881, 561, 400, 852, 350, 411, 672, 879, 907, 267, 882, 716, 779, 330, 281, 154, 794, 44, 173, 235, 317, 437, 382, 494, 535, 552, 563, 608, 732, 627, 698, 798, 839, 929, 957, 899, 947, 462, 679, 721, 780, 471, 925, 940, 758, 101, 171, 311, 520, 572, 671, 932, 323, 664, 440, 168, 826, 12, 101, 92, 664, 825, 775, 219, 921, 680, 884]. Hitherto [726]. Hitti [486]. Hoary [331]. Hobbes [391]. Hobbs [95]. Hofstadter [594]. Hogben [29]. Hole [618]. Holinshed [712]. Holland [515, 3]. hollandaises [773]. Holmes [188]. Homage [732]. hombre [272]. homme [864]. Hon [132]. Honig [615]. Honigmann [117]. Honor [523]. Honourable [546]. Hooke [17, 158, 744, 950]. Hooton [102]. Hope [725, 927]. Hoppe [209]. Horace [367, 778]. Horine [915]. Hormones [456]. Horse [242]. Houdon [940]. Howard [40, 450, 311, 671]. Howells [205]. Howland [380, 380]. Hsi [549]. Hudson [667]. Huellas [252]. Hugo [90]. Human [258, 458, 101, 28]. Humana [727]. Humanism [373]. humanistes [936]. Humanities [821, 100]. Humboldt [500]. Hume [199, 312]. Hummel [453]. Hundert [144]. Hundred [726, 673]. Hunger [101, 205, 101]. Hunter [480]. Huntley [365]. Hurricanes [521]. Huxley [518]. Huygenian [846]. Huygens [914, 690]. hydrauliques [754]. Hystaspe [25].

Iago [463, 257]. Ibn [63, 693, 13, 538, 529, 362, 848, 238]. Iceland [418]. Idea [855, 705, 827]. Ideals [133, 514, 662]. Ideas [326, 268, 881]. idées [529]. Ideologues [547]. II [714, 198, 266, 9, 67]. III [58, 852, 611]. Ilhani [52]. Ilim [730, 925]. Ilimdir [926]. Illustrations [906]. ilmi [592]. Image [482]. Images [856]. Imagination [292]. Implications [869]. Impressi [266]. including [878]. inclusive [44]. Incomparable [478]. Incunabula [505, 803, 54, 532, 906]. Indes [773]. Index [293]. Index-Catalogue [293]. India [677]. Indies [615, 451]. Induction [733]. Industry [835, 878, 572]. Inedili [729]. inédite [652]. inediti [341]. Infamy [941]. Infeld [230]. Inference [873]. Influence [851, 416, 950, 101]. influencin [272]. Informal [671]. Information [106, 295, 319, 466, 554, 578, 624, 684, 700, 740, 785, 840, 960]. Inhalation [737]. Injektion [763, 723]. Inks [691]. Inquiry [89, 593, 937]. Inquisizione [760]. Inscriptions [86]. Insect [633]. Institut [26]. Institution [10, 88, 436]. Instrument [51]. Instruments [904]. intellectuelle [816]. Intelligencers [4]. International [959, 291]. Interpretation [660, 282]. Interpretations [9]. Intravenösen [763, 723]. Introduction [940, 762, 314, 852, 254, 374]. Introductory [118]. intuition [127]. intuitives [127]. Invention [392, 695, 323]. Inventions [726]. Inventor [584]. Investigator [614]. Involved [745]. Ira [931]. Ireland [780]. Irish [771]. Irma [511]. Irving [809]. Isaac [762, 187]. Isis [161, 643, 116]. Islam [252, 84]. islamici [621]. Island [80, 755]. Islands [519]. Isoperimetric [58]. Israel [454]. Italian [894, 265]. italiana [516]. italiani [516]. italiano [516]. Italy [920, 118]. Ivan [521]. Ives [405]. IX [456, 907].

J [928, 209, 232, 291, 768, 775, 660, 817, 777, 406, 23, 144, 857, 210, 216, 229, 231, 411, 483, 672, 290, 134, 309, 515, 814, 815, 910, 218, 654, 912, 776, 212, 931]. J. [951, 761]. J.-J. [951]. Jabir [529]. Jac. [589]. Jackson [769]. Jacob [15]. Jacques [431, 682, 854, 951]. Jaeger [133, 514, 662]. Jaffe [569]. Jahre [144]. Jahrhunderts [763, 723]. James [31, 292, 731, 213, 876, 512, 869, 188, 79, 93, 532, 419, 462, 520, 385]. Jan [42]. Janet [492]. January [563, 698, 839]. Japan [121]. Jaques [432]. Jasny [658]. Jean [823, 511, 788, 951, 324, 70]. Jean-Jacques [951]. Jeans [876]. Jefferson [649, 649, 649]. Jerome [140]. Jersey [364]. Jesse [436, 413, 131]. Jessen [877]. Jésuites [936]. Jew [24, 202]. Jewett [139]. Jewish [485, 82]. Jews [484, 550]. Jheronimus [854]. Jirka [218]. jo [121]. Joad [872]. Joannes [185]. John [291, 613, 726, 943, 302, 811, 531, 187, 367, 487, 517, 97, 414, 368, 573, 921, 656, 724, 572, 393, 747, 470, 328, 354, 267, 802, 602, 953, 694, 566, 250, 478, 407, 590]. Johnson [944, 764, 734, 678, 118]. Johnston [136]. Jonathan [856]. Jordan [201]. Jose [272, 452]. Josef [664, 77]. Joseph [934, 27, 209, 664, 455, 481, 152, 345, 13, 866, 25, 481]. Joshua [82]. Josua [773]. Journal [251, 762]. Journey [613]. jours [209]. Jr [523]. Jr. [612, 316]. Jules [708, 409]. Julia [830]. Julian [819, 518]. Julius [911]. July [613, 535]. Jumar [850, 332, 242]. June [437]. Jung [790]. Junior [802]. Jupiter [802]. Jurokuseiki [121]. Justice [675]. Justina [275]. Justus [148].

K. [628]. Kāshī’s [711]. kô [80]. Kahler [595]. Kallen [367, 778]. Kantian [126]. Kara [119]. Karakoyunlu [119]. Karl [431, 358, 921, 877, 72]. Karpinski [208]. Kasner [292]. Katalyse [142]. Keep [313]. Kelly [591]. Kelsen [593]. Kenneth [409]. Kentucky [915]. Kepler [952]. Keppel [370]. Kesten [646]. Key [882, 461]. Keys [679]. Keyser [769]. Khatai [52]. khi [866]. Kibat [549]. kills [388]. King [549]. Kingdom [866]. Kircher [834]. Kirfel [340]. Kisselgoff [921]. Kitab [184]. Klüver [459]. Knowledge [402, 57, 689, 204]. Known [749]. Koch [456]. Kommentar [908]. Korean [541]. Kosher [202]. Koyré [648]. Koyré’s [389]. Koyunlu [119]. Kraus [529, 183]. Kremers [219]. Krisen [415]. Kristeller [477]. kritische [858]. Kroeber [674]. Kuhn [549]. Kultur [883]. Kulturphilosophie [883]. Kungl [125]. Kungl. [153]. Kurt [883, 143]. Kurze [142]. Kusaka [765]. Kuster [144].

L [92, 232, 310, 674, 858, 379, 129, 373, 483, 316, 508, 619]. Laboratory [618]. Lagrangiana [894]. Lahire [710]. Lamers [814]. Lancelot [29]. Land [482, 566]. Landen [470]. Landheer [515]. Láng [346]. langage [951]. Langer [461]. Language [114, 747, 33, 951, 764]. Languages [42]. L’Antiquité [935]. Laplace [604, 221]. L’apogée [864]. Large [377, 548]. l’Astrolabe [822]. Late [186, 334]. Later [326]. Latin [140, 575, 670, 687, 500, 388, 693, 539]. Latini [910]. latins [93]. Laue [770]. Laurence [675]. Lavoisier [406, 651]. Law [581]. Laws [485]. Layard [487]. Layman [868]. Lazare [365]. Lea [380, 921]. Learning [852, 732]. Lee [413, 937]. Leeuwenhoek [124, 124]. Lehmann [405]. Lehmann-Haupt [405]. Lehre [616]. Leibniz [287]. Lejeune [919]. Length [526]. Lenoble [913]. Lenzen [291]. Leo [676, 229, 18]. Leon [572, 62]. Leonard [291, 448]. Leonardo [482, 511, 680, 64]. Leone [53]. Leonora [255]. Leopardi [395]. Leopold [230]. l’Escurial [309]. Leslie [251]. Lesquereux [321]. Lesser [14]. Lesson [179]. lessons [323]. Letter [391, 122, 388]. Letters [132, 640, 131, 647, 124, 251, 155, 531]. lettres [652]. Leven [814, 815]. LeVene [192]. Levi [621]. Levin [202]. L’évolution [823]. Levy [63]. Lewes [56]. Lewis [251]. Ley [930]. Leyes [727]. L’humanisme [874]. Lhwyd [647]. Li [922, 910]. Liao [921]. Libraries [505]. Library [492, 657, 41, 293, 269, 576, 327, 952, 79, 532]. Libri [266]. Lie [893]. Liebig [408, 408]. Life [435, 132, 765, 72, 926, 546, 406, 811, 810, 17, 20, 923, 129, 216, 647, 316, 915, 940, 337, 128, 517, 131]. Light [160, 136, 5, 209, 345]. Lightning [138]. Lillie [618]. Lima [272]. Limitation [401]. Limoges [847, 557]. Lincoln [502]. Lindsay [28]. Linnaeus [746]. Lionel [531]. Lionnais [874]. l’Islam [529]. List [41, 41]. Literary [825, 511, 245, 944]. Literatur [616]. Literature [238, 256, 203]. Littera [388]. litteraire [547]. liu [835]. Lived [315]. Livestock [520]. Living [657]. Livre [164, 908]. livres [910, 67]. Lloyd [435, 676, 417, 917]. Logan [460]. Logic [821, 99, 753, 254, 31, 206]. Loi [709]. London [204]. Looks [606]. Loon [217]. Lord [132]. Lore [40]. Lorenz [858]. Lost [391]. Lotka [23]. Lotus [549]. Louis [136, 726, 546, 412, 485, 23, 147, 823, 81, 129, 208, 134]. Lovage [244]. Lovejoy [881]. Lowie [640]. Lowinger [308]. Lucca [118]. luce [209]. Lucien [531, 531]. Luck [213]. Lucretii [448]. Ludwig [660]. Luise [340]. lumière [209]. lumine [186]. Lunar [7, 560]. Lungfish [930]. Lunt [435]. Luqa [538]. Lycie [908]. Lynam [735]. Lynde [726]. Lynn [909, 612, 285]. Lysenko [668].

M [523, 817, 485, 313, 339, 433, 228, 872, 737, 192, 100, 216, 367, 778, 13, 939, 732, 814, 654, 853, 458]. M. [132]. M.D [458]. Mabel [129]. MacCurdy [511]. Mach [209, 640]. Machine [255, 745, 376, 157, 395, 255]. maconnerie [816]. Macrobiusxo [804]. maestri [265]. Magendie [654]. Mages [25]. Magic [285, 82, 259]. Magnet [602]. Magnitudes [560]. Magnus [111, 540]. Maimon [248]. Maimonide [247]. Maimonides [469, 248]. Maistre [164]. Major [485]. Maker [655]. Making [857]. Malebranche [67]. Malekula [487]. Malinowski [329]. Malpighi [301]. Man [827, 777, 315, 483, 868, 942, 381, 275, 102, 214, 255, 460, 619, 433, 152, 595, 253, 522]. Man-Machine [255]. Manchester [245]. Manent [525]. Manetho [289, 289]. Mankind [882]. Manniparous [633]. manoscritti [621]. Mansions [7]. Manuscript [112, 386, 185]. Manuscriptis [266]. Manuscripts [140, 290, 532, 545]. manuscrites [773]. manuscritos [452]. manuscrits [93, 309]. Many [880]. Map [541]. Mappemonde [773]. Maps [121, 156, 95]. Marat [70]. Marcel [773]. Marcelle [166]. Marcello [301]. March [168, 878, 348, 957]. Marco [482]. Marcus [24]. Margerie [570, 774]. Marguerite [677]. Maria [664, 341, 452, 815]. Marie [708]. Marin [812]. Marine [632, 618]. Marion [920]. Marjorie [725, 510, 927]. Mark [522, 157]. Markham [497]. Markwart [77]. Marliani [334, 335]. Marriage [485]. Marrou [935]. Marsh [192]. Marshall [334]. Marsilio [477]. Martha [449]. Martin [74, 734, 678, 828]. Marvin [559]. Mary [141]. Mask [940]. Maspero [200]. Massachusetts [958]. Massé [84]. Masters [265]. Material [749]. Materials [380]. Mathematical [934, 32, 178, 560, 208, 951, 659, 206]. Mathematician [292, 470]. Mathematics [168, 533, 9, 58, 8, 592, 843, 292, 928, 207, 929, 769]. mathématique [273, 951, 874]. mathématiques [273, 864]. Mather [188, 188, 694]. matiere [247]. Matter [46, 1, 47, 107, 175, 220, 237, 261, 318, 351, 277, 296, 297, 320, 383, 421, 438, 467, 352, 384, 422, 439, 495, 524, 536, 555, 577, 556, 598, 625, 579, 599, 683, 701, 739, 626, 685, 702, 784, 799, 741, 786, 800, 841, 888, 901, 945, 961, 842, 889, 902, 946, 136]. Matthew [392, 478]. maturité [76]. Maurice [482, 98, 652]. Maurolycus [186]. Max [860, 939, 770, 861, 247, 464, 853, 398]. Maxson [328]. May [235, 316]. Mayer [293, 345, 293]. Mayo [428]. Mayr [488]. McAllister [228]. McColley [203]. McKeehan [726]. Mead [38]. Mean [434]. Meaning [734, 100]. Means [366]. Measure [595]. Measurements [844, 748]. mécanisme [913]. Mechanical [726, 745]. Mechanics [726, 375]. Mechanism [137]. Mechanisms [459]. médecine [885, 68]. Mediaeval [238]. Medical [505, 941, 492, 350, 474, 610, 293, 312, 532, 780, 829, 940, 596]. medicale [247]. Medici [71]. Medicinae [814]. Medicine [943, 241, 361, 464, 654, 884, 550, 257, 349, 939, 199, 419, 258, 830, 348]. Medieval [186, 334, 24, 290, 79, 532]. médiévale [818]. Medizingeschichte [763, 723]. Meeting [958, 933]. Meh [549, 549]. Mei [549]. Meir [809]. Mémoires [134]. Memorialia [71]. Memories [193, 573]. Men [549, 487, 569]. Mendeleyev [916]. Menges [921]. Menninger [886].

Mention [633]. Mercati [760]. Mercator [323]. Merian [815]. Merle [590]. Merriam [414, 97]. Mersenne [812, 391, 913]. Merton [16]. Metallurgical [893]. Metaphysics [224]. ’Method [392, 705, 358, 904, 756, 345, 361]. Methode [90, 127]. Methodology [308, 254, 415]. Methods [232]. Mettler [884, 884]. Metzger [66]. meunerie [754]. Meyer [146, 146]. Meyerhof [247]. Meyrick [19]. Miall [549]. Michael [420]. Michel [822]. Michigan [368]. Microbe [152]. Microbes [781]. Micrographia [17]. Micromerism [801]. Middle [789, 346, 265, 754]. Mieli [651]. Mielke [941]. Miguel [252]. Mikhail [584]. Mildred [567]. Millar [434]. Millas [13, 452]. Miller [396, 31, 138, 856, 869, 336]. milling [754]. Milton [259]. Mind [168, 22, 736]. Mineral [878]. Minime [812]. Minkin [15]. Minor [380, 111]. Mint [724]. Miracles [781]. Mirandola [249]. Missionary [590]. Mitscherlich [941]. Mittasch [142]. Mlle. [224]. Moamin [663]. Mode [669]. Modena [760]. Modern [27, 435, 137, 498, 99, 872, 721, 387, 691, 518, 39]. moderna [651]. Modes [103]. Mogens [858]. Mohammadi [183]. moindre [225]. Moira. [661]. Moll [575]. Monarchie [442]. Money [48]. Monge [942]. Mongol [51]. Montagu [313, 433, 732, 458]. Monument [903]. Moon [927]. Moore [139, 491]. Moravian [478]. morbis [189]. Morgan [251, 366, 251]. Morgenthau [777]. Morison [825, 14, 337]. Mörl [863]. Morley [86]. Morphogenesis [455]. Morris [291, 918, 649, 38]. Morrison [689]. Morrow [489]. Morse [410]. Morton [280]. Morveau [69]. Moses [248]. Moslem [51]. Most [433]. mot [635]. Motion [638, 157, 833]. Moulton [408]. Mountain [129]. Mourad [184]. mouvements [34]. Movements [325]. Movius [483]. moyen [265, 754]. MS [753]. Muhammad [336]. Munich [753]. Murex [632]. Muriel [429]. Murray [213, 703]. Mürsit [926]. Muse [725]. Music [346, 939, 346]. Musical [587]. Muslim [370, 303]. Mustafa [362]. Mustard [331]. musulman [303]. My [73, 128]. Myer [131]. Mysteries [94, 816]. Mysticism [485]. Myth [433]. Mythology [819].

N [125, 171]. Na-khi [866]. Nagel [291]. Nahm [523]. naissance [913]. Nallino [341, 341, 729]. Name [755]. Nanda [870]. Narratio [122]. Nat [42]. Nath [340]. Nathan [191, 650]. Nationale [773]. Native [102, 349]. natura [448]. Natural [632, 688, 204, 338, 12]. Naturalist [393, 548]. Naturalists [824]. Nature [593, 474, 211, 209, 864, 481, 253, 167]. naturelle [66]. Naum [658]. nautiques [773]. Navies [501]. Navin [187]. Nazi [941]. Nazif [362]. Neckam [824]. nécropole [166]. Nederland [42]. Need [471, 940]. Needham [27, 455]. Needs [933]. Neff [826]. Nehemiah [301]. nel [760]. Nelson [201]. Neo [55]. Neo-Aristotelianism [55]. Netherlands [615, 42]. Neuburger [464]. Neugebauer [659]. Neuman [484]. Neurath [291]. Neurologist [817]. Newman [869, 292]. Newton [725, 950, 187, 66, 725, 724]. Next [619]. Nicholas [875]. Nicholson [927]. Nicolaus [911, 853]. Nicole [164]. Nicolson [725]. Niels [291]. Nienhuis [815]. Nietzsche [366, 288]. Nihon [121]. Nine [756]. Nineteenth [342, 817, 42, 638]. ninth [154, 608]. Nixon [780]. no [121, 887]. Norbert [646]. Norden [941]. Nordenmark [125]. Nordenskiöld [718]. Norman [83, 573]. North [103, 574, 407, 49]. Northrop [821]. Norwood [596]. nos [209]. Note [396, 710, 690, 562, 705, 693, 53, 708, 952, 389, 691, 335]. Notebooks [511]. Notes [11, 182, 333, 544, 163, 364, 837, 632, 896, 948, 404, 528, 476, 565, 609, 227, 246, 270, 286, 447, 717, 59, 115, 588, 696, 225, 757, 3, 507, 305, 642, 793, 954, 427, 268, 360, 762]. Notice [718, 719]. Notices [26]. notion [823]. Notis [266]. nourrir [200]. nouvel [166]. Novel [229]. November [899]. Number [33]. Numerals [112]. Numeration [503].

O [132, 659, 881, 298, 771]. O. [192]. Obermann [819]. Obituaries [645, 697, 720, 759, 797, 807, 838, 895, 956]. Obituary [718, 719]. obra [848]. Observaciones [272]. Observations [846, 752]. Obstacles [326]. Occasion [732]. occident [303]. Occidit [388]. Ocean [337]. October [44, 173, 382, 798]. OEuvres [67]. Offered [732]. Office [293]. Ohio [638]. Ohm [727]. Okamoto [121]. Olaus [165]. Old [314, 922]. Oldenburg [4]. Older [830]. Oliver [373, 315]. Olivier [68]. Olmsted [817, 654]. Olschki [482]. One [880, 673]. oostersche [42]. Open [22]. Opera [183]. Opinions [867]. Optick [582]. Opticks [725]. Optics [209, 186]. Optik [209]. Optimi [71]. optique [919]. Opuscule [117]. opuscules [273]. Orach [805]. ordinates [753]. Oresme [164]. Organic [400]. Organism [143]. organizados [272]. Oriental [722, 42]. Orientales [773, 452]. orientali [621]. Origin [444, 58, 332, 344, 488]. origine [635]. origines [127]. Ornithologists [312]. Ornithology [394]. Orthodox [202]. Orville [591]. Osborn [233, 871]. Osborne [767]. Oscar [415]. Oskar [477]. Osmanli [730]. Ostanès [25]. Other [380, 769, 201, 887, 300]. Otis [937]. Otto [291, 549]. Ottoman [730]. Ottomans [85]. Our [490, 870, 871]. Outline [168, 932, 929]. Owsei [622]. Oxford [647]. Oyler [726].

P [523, 168, 89, 490, 309, 812, 22, 667]. Pachymère [363]. Pagel [27]. Paideia [133, 514, 662]. Pain [145]. Palace [336]. Palacios [252]. Palaeontologist [193]. Paleontology [192]. Palestinian [412]. Pamphlets [191]. Papanastassiou [209]. Papermaking [472]. Papers [97, 649, 813]. Pappus [611, 751]. Papyri [263, 368]. Paradijzen [815]. Parallels [156]. parchemin [773]. Parentage [102]. Paris [93, 788, 752, 773]. Parker [118]. Parmalee [101, 571]. Parsons [120, 878]. Part [17, 266]. partem [728]. Pasquier [134]. Past [829]. Pat [780]. Pathfinder [726]. Patrick [655]. Patriot [365]. Paul [435, 39, 273, 491, 479, 529, 879, 70, 574, 134, 346, 363, 511, 812, 477, 708]. Paulus [183]. Pays [68]. Pearson [72, 72, 358]. Peck [508]. Pehr [125]. Pei [952]. Pei-t’ang [952]. Peirce [148]. Pekin [952]. Pelseneer [823]. Penicillin [905]. Pennsylvania [948]. Penrose [316]. pensée [76, 874]. People [420]. Peoples [216]. Perfection [64]. Period [244, 331, 805, 453, 920]. Perpetual [638, 157]. Perplexities [628]. Perry [856]. Persistence [827]. Personalia [644, 795, 808, 897, 955]. Petén [86]. Peter [499, 847, 278, 557]. Petrarca [890]. Petrus [773]. Pfeiffer [314]. Ph.D [726]. Pharisees [81]. Pharmacy [219]. Phelps [726]. phénomène [823]. Phenomenology [828]. Philadelphia [440]. Philemon [3]. Philip [486, 456, 655]. Philipp [765, 343, 665]. Philo [903]. Philosophers [657]. philosophica [183]. Philosophical [209, 245, 649, 687, 523, 872, 620]. Philosophy [21, 704, 491, 481, 169, 38, 574, 657, 154, 44, 173, 235, 317, 437, 382, 494, 535, 552, 563, 608, 698, 798, 839, 957, 899, 477, 126, 343, 779, 461]. Phoenician [893]. Photismi [186]. Photography [664, 664]. Physicae [280]. Physical [209, 169, 876]. Physici [335]. Physician [862]. Physicist [686]. Physics [789, 168, 399, 343, 136, 137, 334, 30, 665, 417]. Physik [90, 770]. physikalischen [209]. Physiker [858]. physiognomonie [184]. Physiologist [862]. Physiology [654, 213]. Physique [35, 823, 209]. Pico [249]. Pictet [430]. Picture [306]. Pierre [308, 127, 18, 134, 847]. Pierson [767]. Pieter [615]. Pihl [858]. P’ing [549, 549, 922]. Pioneer [417, 192, 654, 915]. Pioneers [251]. Pires [722]. Pirotechnia [449]. Pissarro [531]. piu [621]. Pla [727]. Plancius [773]. Planck [860, 861]. Planet [871]. Plant [311, 670]. Plants [670]. Plate [711]. Plato [109, 893, 322]. Platon [864]. Platonism [39]. Plea [703]. Pledge [168]. Pliny [12]. Ploughs [364]. Plundered [871]. Pneumatica [851]. Poetic [114]. Poetry [826, 167]. Poets [725]. Pol [853]. Political [869]. Politicks [364]. Politics [196, 777, 215]. politischer [287]. Pollen [630]. Polo [482]. Polonia [558]. Polyak [379]. Polygraphy [834]. Poor [191]. Pope [135]. Population [707]. Portrait [158]. Portraitures [940]. portulans [865]. Posidonius [445]. Posin [916]. Position [49]. Positivism [282]. positivisme [127]. possibilité [562]. Possibilities [933]. Powder [310]. Powell [862]. Power [777, 151, 376, 411]. Practice [492]. Pradas [663]. Pratt [99]. Praxis [142]. Pre [846]. Pre-Huygenian [846]. Precursors [482]. Preface [48, 2, 108, 155, 221, 298, 321, 262, 385, 423, 440, 353, 468, 496, 525, 580, 627, 703, 787, 742, 843, 890, 947]. Prehistoric [920]. Prelude [905, 721]. premier [908]. Prentice [101, 571]. Present [209, 829]. Pressure [479]. Prester [566]. Preventive [943]. Priestley [209]. prima [122]. Primer [868]. primitifs [823]. Primitive [216, 745]. Princeps [249]. principe [200, 225]. Principle [135, 340]. Principles [209, 345, 745]. Printed [239, 637, 208]. Prinzipien [209]. privilégiés [35]. Problem [576]. Problems [390, 58, 767, 707]. procédés [200]. Proceedings [878]. Processo [760]. Proclus [908]. Professor [399]. Profile [191]. Prognostications [266, 712]. Progres [180, 98]. Progress [796, 855, 878, 23, 75, 257]. progresso [516]. projection [323]. Proklus [908]. Prolegomenon [891]. Promise [373]. Promotion [204]. propos [356]. Pseudo [750]. Pseudo-Science [750]. Psychiatric [498, 673]. Psychiatry [673, 886]. Psychologische [776]. Psychology [99, 350]. psychophysique [62]. Ptolemaic [156]. Ptolémée. [919]. Ptolemy [289, 705, 509, 509]. Public [327, 41]. Publications [639, 750]. publié [762]. publiques [93]. Published [948, 97]. Pursuit [64, 689].

Q [916]. Qarāfī [50]. Quadrans [692]. Quadratic [58]. Quadripartite [509]. Quadrivium [363]. quae [183]. Qualifications [627, 947]. Quantitative [748]. Quarrel [416]. quartam [728]. Quelpaert [755]. Quelques [303, 66]. quem [110]. Queries [226, 641, 836, 403, 181, 527, 506, 543, 564, 304, 162, 446, 715, 3, 359, 475]. Quest [816]. Quest. [230]. quickens [388]. Quiggin [480]. química [651]. Quine [206]. Qusta [538].

R [911, 168, 408, 17, 24, 517, 647, 869, 610, 918, 618, 667]. R. [848]. Rabbi [809, 13]. Rabelais [706]. Rabin [721]. Raccolta [341, 729]. Race [433, 215]. Rackham [12]. Radio [92]. Raghallaigh [771]. Raghensis [183]. Rainbow [50]. Raising [520]. Ralph [738]. Ramazzini [189]. Raphael [63]. Raven [824]. Ray [521, 393, 340, 824]. Raymond [32, 364, 929]. Rays [138]. Razi [184]. Read [364, 649]. Reader [650]. Reason [855, 887, 938, 522]. Réaumur [652]. récents [98]. recherche [67]. Recipe [756]. Reckoner [264]. Reconstruction [108]. Record [204, 113]. recueillies [652]. Reed [311]. réels [34]. Reese [346]. Reference [737, 43, 44]. References [581, 601]. referred [268]. Reflected [687]. Reflection [581]. Reflections [686]. Réflexion [709]. Regimen [3]. Reiser [373]. relatifs [198]. Relating [209, 41]. Relations [231, 18]. relativiste [35]. Religieux [812]. Religion [454, 604, 221, 925, 200, 66, 82, 197]. Religious [620]. Remarks [6, 8, 355]. Remedies [890]. Renaissance [416, 290, 532, 120, 616]. Renaud [309]. René [885]. répétitions [562]. Report [613, 116, 43]. Represented [112]. Republic [155]. Republican [365]. Republication [843]. rerum [448]. Research [10, 829, 415, 576, 88, 483]. Resources [878]. Respiratory [325]. Rethe [791]. Retina [379]. Retreat [887]. Reucker [431]. Revealed [95]. Review [934, 15, 27, 29, 40, 380, 783, 435, 676, 523, 617, 875, 769, 928, 821, 827, 855, 924, 191, 375, 763, 121, 335, 911, 21, 31, 32, 65, 91, 92, 132, 136, 137, 138, 168, 186, 207, 209, 232, 250, 272, 291, 292, 342, 364, 417, 533, 591, 613, 650, 651, 664, 725, 726, 639, 765, 768, 825, 856, 944, 775, 549, 430, 271, 72, 88, 139, 140, 449, 771, 659, 39, 76, 133, 273, 514, 660, 661, 662, 908, 941, 942, 943, 936, 164, 334, 455, 510, 663, 142, 170, 310, 429, 666, 883, 491, 206, 730]. Review [926, 546, 731, 817, 534, 949, 63, 84, 201, 202, 234, 248, 314, 369, 372, 412, 434, 454, 484, 485, 486, 550, 809, 819, 826, 867, 34, 35, 89, 99, 308, 95, 313, 371, 381, 408, 436, 482, 566, 674, 677, 681, 736, 772, 777, 870, 914, 916, 927, 69, 365, 406, 593, 339, 77, 811, 96, 19, 149, 880, 938, 94, 338, 416, 646, 810, 16, 64, 249, 858, 307, 431, 433, 141, 12, 17, 20, 23, 36, 60, 74, 98, 120, 144, 146, 147, 167, 213, 228, 259, 379, 410, 432, 448, 456, 459]. Review [481, 738, 165, 547, 569, 864, 872, 463, 575, 737, 886, 887, 923, 878, 37, 152, 219, 257, 275, 315, 377, 781, 30, 90, 169, 343, 665, 733, 734, 823, 860, 873, 876, 937, 881, 512, 28, 38, 192, 215, 413, 479, 530, 531, 567, 595, 678, 682, 857, 492, 852, 102, 127, 529, 24, 81, 82, 83, 75, 78, 100, 103, 128, 129, 130, 131, 143, 145, 150, 151, 172, 187, 190, 194, 196, 197, 203, 205, 210, 214, 216, 217, 229, 230, 231, 251, 253, 255, 256, 274, 288, 345, 349, 350, 366, 367, 373]. Review [409, 411, 450, 460, 480, 483, 487, 490, 517, 594, 647, 655, 672, 675, 778, 782, 879, 863, 13, 122, 135, 252, 939, 70, 104, 622, 761, 854, 922, 874, 882, 86, 341, 764, 869, 574, 620, 648, 657, 779, 828, 97, 193, 376, 616, 909, 770, 861, 610, 117, 188, 200, 290, 347, 370, 414, 568, 612, 615, 722, 80, 14, 18, 26, 33, 41, 42, 43, 61, 62, 68, 79, 85, 87, 93, 123, 124, 125, 134, 153, 166, 183, 184, 189, 195, 198, 199, 204, 208, 247, 258, 287, 289, 293, 306, 309, 312]. Review [316, 340, 346, 362, 363, 368, 418, 420, 452, 453, 508, 509, 511, 515, 516, 532, 570, 573, 589, 592, 611, 614, 621, 652, 728, 729, 732, 735, 762, 766, 773, 774, 812, 814, 815, 816, 818, 820, 830, 865, 866, 885, 910, 913, 918, 919, 920, 921, 929, 71, 218, 348, 361, 419, 428, 462, 464, 513, 545, 551, 654, 656, 673, 679, 680, 721, 780, 829, 853, 862, 884, 917, 119, 336, 458, 596, 461, 619, 760, 915, 478, 813, 912, 724, 254, 374, 767, 25, 925, 477, 776, 935, 723, 940, 822]. Review [22, 415, 118, 337, 451, 576, 727, 148, 211, 378, 73, 101, 171, 185, 212, 233, 311, 344, 407, 457, 488, 489, 518, 519, 520, 521, 522, 548, 571, 572, 590, 597, 618, 649, 653, 658, 667, 668, 669, 670, 671, 824, 859, 868, 871, 877, 930, 931, 932, 933, 66, 67, 126, 405]. révolution [816]. Rewald [531]. Rey [76, 864]. Rheticus [122]. Rhyme [522]. Ricci [290]. Ricerche [621]. Richard [191, 533, 594, 680, 829, 211, 233, 847, 735]. Richens [667]. Richer [324]. Richter [511]. Rider [576]. Ring [846]. Rise [146, 458]. Ritter [36]. riunione [516]. riyadiyyat [592]. Road [371, 781]. Roback [512]. Robbins [533, 289]. Robert [783, 875, 944, 546, 314, 149, 16, 17, 913, 671, 640, 325, 158, 744, 719]. Robertus [441]. Robinson [129]. Rock [866]. Rodgers [407, 590, 653]. Roemer [5, 792]. Roman [920]. Romance [663, 932]. Romanorum [728]. Romantic [930]. Romanticism [338]. Rome [611, 369]. Ronchi [209]. Rosen [765, 122]. Rosenblatt [867]. Rosenfield [255]. Rossiter [168]. Rostovtzeff [339]. Rothenburg [809]. roues [754]. Rousseau [951]. Roxbrough [573]. Royal [875, 204, 687, 302]. Rozbroj [70]. Rozelle [118]. Rt [132]. Rubber [489]. Rudbeck [165]. Rudolf [291, 374]. Rudolphinae [952]. Rufinus [612]. Rukeyser [429]. Runner [129]. Russell [291, 367, 657, 779]. Ruth [870, 215]. Rutherford [132, 132]. Rynin [764].

S [15, 435, 821, 132, 209, 664, 768, 72, 202, 772, 738, 131, 210, 869, 612, 921, 171, 311, 667]. Saadia [867]. Sachs [659]. Sacrobosco [909]. Saggio [894]. Salerni [3]. Salo [248, 485]. Salt [835]. Salvemini [172]. Samuel [726, 825, 867, 938, 781, 14, 337, 405]. Sanceau [566]. Sand [264, 885]. Sand-Reckoner [264]. Sandreuter [431]. Sanitatis [3]. santé [68]. Sarton [54, 923, 852]. Satellite [802]. Saturn [846]. Saudiana [341]. Sayili [926]. Schechter [83]. Scheiner [55]. Schillinger [934]. Schillp [574, 657]. Schilpp [491]. Schmidt [144]. Schneider [816]. Schoen [939]. Scholar [576]. Scholars [123]. Scholastic [249]. Scholem [485]. School [336, 464]. Schools [43, 43]. Schrecker [879]. Schuchert [192]. Schullian [939, 853].

Schutz [663]. Schweizer [87]. schweizerischen [307]. Science [958, 959, 686, 500, 21, 168, 342, 726, 704, 750, 285, 730, 926, 731, 914, 898, 338, 569, 923, 561, 215, 857, 852, 345, 647, 716, 154, 794, 44, 173, 235, 317, 437, 440, 382, 494, 535, 552, 563, 608, 732, 627, 698, 798, 839, 957, 899, 947, 721, 606, 471, 925, 758, 670, 864, 85, 10, 168, 27, 40, 291, 417, 726, 872, 169, 876, 413, 203, 231, 33, 918, 767, 590, 868, 29, 768, 372, 16, 734, 75, 196, 615, 195, 868]. Sciences [821, 891, 688, 180, 864, 570, 26, 96, 515, 254, 311, 198]. Scientific [114, 505, 803, 223, 772, 54, 705, 678, 345, 373, 517, 497, 13, 324, 592, 742, 654, 387, 649, 777, 738]. scientifico [516]. scientifique [76, 874, 924]. scientifiques [529, 134]. Scientist [228, 916, 172, 230]. Scientists [607, 615, 820, 28]. scienze [516]. scienziati [516]. Scotland [41]. Scott [193]. Scripta [525]. Scritti [729, 341]. Sea [337, 376]. Seals [78]. Sebastian [530]. Secolo [760, 516]. Second [14, 2, 108, 221, 321, 496, 742, 839, 890, 947, 415, 317]. Secondary [43]. Section [10]. Sector [95]. Sedgwick [168]. Segal [147]. Seidel [944]. sekaichizu [121]. sekreterare [125]. Selbstbiographie [861]. Selection [631]. Selections [783]. Semantics [374]. Semicentennial [32]. Séquard. [817]. seres [272]. Servant [868]. Serveto [706]. Sesmat [34, 35]. Seventeenth [914, 16, 843, 393]. Seventeenth-Century [393]. Seventh [552]. Seventieth [698]. Seventy [798, 839, 957, 899, 878]. Seventy-First [798]. Seventy-Five [878]. Seventy-Fourth [957]. Seventy-Second [839]. Seventy-Third [899]. Sewall [862]. Sex [145, 721, 456]. Seymour [290]. Sha [756]. Shadid [420]. Shadows [856]. Sheep [20]. Sherborn [573]. Sherrington [253]. Sherwood [65, 168]. shi [80]. Shinkichi [80]. Ship [530]. Shorey [39]. Short [168, 342, 462, 311]. Shryock [829]. Shuichi [765]. Sibylla [815]. Sickness [622]. Side [201]. Sidney [585, 559, 406]. siècle [68, 816, 93]. Siècles [180]. Sigerist [830, 534, 258]. Sigfried [274]. Sigmund [949]. Significance [240]. Silliman [726]. Silphium [179]. Silverman [259]. Simon [419, 912]. Simpson [669]. since [168]. Singer [523, 721, 342]. Sir [302, 392, 917]. Six [549]. Sixteenth [121, 266, 600, 843]. Sixteenth-Century [266]. Sixth [440, 535, 959]. Sixtieth [173, 732]. Sixty [235, 317, 437, 382, 494, 535, 552, 563, 608]. Sixty-eighth [563]. Sixty-Fifth [494]. Sixty-First [235]. Sixty-fourth [437]. Sixty-ninth [608]. Sixty-second [317]. Sixty-Seventh [552]. Sixty-Sixth [535]. Sixty-third [382]. Slavery [401]. Smallpox [681, 80]. Smith [449, 448, 456]. Sobre [848, 272]. Social [435, 676, 436, 339, 231, 631, 869, 40, 345, 594]. sociale [885]. Societa [516]. société [936]. Societies [738]. Society [958, 32, 16, 204, 921, 245, 649, 687, 302, 675, 593]. Sociological [593, 81]. Sociology [130]. Soil [401]. Solomon [83]. Solution [576]. Some [789, 750, 72, 948, 6, 3, 303, 193]. Sommario [760]. Son [531, 323]. Song [394]. Song-Birds [394]. Sorcery [537]. Sorokin [490]. Source [918, 597]. Sources [712, 721]. Soutar [167]. Southard [22]. Southwest [866]. Souvenirs [430]. Soviet [830, 667]. sozialen [776]. Space [126, 274]. Spain [484]. Spallanzani [326]. Spanish [238, 848]. Sparks [138]. Speak [820].

Special [737, 43, 44]. Species [488, 344]. Spectroscopy [396]. Speculative [281]. Speter [398]. Sphere [804, 909, 833]. Spheres [243, 694]. Spiller [944]. Spinach [805]. Spinka [478]. Spinoza [424]. Spirit [361, 387, 388]. Spiritus [388]. Spring [357]. Spurious [158]. Squire. [573]. Srole [676]. St [602]. Stace [381]. stades [919]. Stand [810]. Stanley [449, 448, 214]. Stapleton [675]. Star [416, 541, 49]. Star-Crossed [416]. Stars [283]. State [209]. States [140, 290, 312, 119, 596, 457, 520, 944, 738, 41]. Station [194]. Statistical [690, 415]. Statistics [803]. Statistique [180]. Status [435]. Stearn [681, 681]. Steele [719]. Stefansson [418, 94, 782]. Stephanus [728]. Stephen [379, 316]. Steps [503]. Stereometry [892]. Stevenson [917]. Stevin [912]. Stiles [159]. Still [810]. Stillman [328]. Stone [234, 487]. Stones [434]. Storehouse [150]. Storia [209]. Story [941, 916, 316, 407, 489]. Streptomycin [781]. Structure [31, 879]. Struik [928, 857]. studie [42, 666]. Studies [827, 751, 9, 58, 753, 560, 634, 224, 703, 843, 471, 279, 907, 372, 732]. Study [851, 911, 109, 102, 82, 197, 904, 869, 80, 42, 8, 742, 118]. Stuldreher [815]. Stuldreher-Nienhuis [815]. Subjects [884]. Successors [735]. suivis [273]. Sulfa [463]. sulla [621]. sull’Eresia [760]. Sullivan [187]. Suma [722]. Summary [949]. Sun [810]. Superstition [82]. supersunt [183]. Surgeon [293]. Surgical [679]. Survey [474, 933, 219]. Surviving [51]. Susanne [461]. Svenska [153]. Swann [772]. Sweden [561]. Swedenborg [190]. Sylvanus [86]. Sylvester [410]. Symmes [694]. Symposia [459, 432, 456]. synekdèmos [117]. Synthesis [518]. Syphilis [80]. System [435]. Systematics [488]. système [34]. Systèmes [35]. Systems [676].

T [168, 207, 139, 772, 17, 100]. T. [647, 448]. tô [80]. Table [45, 105]. Tables [300]. Tabulae [721, 952]. Takahashi [80]. talen [42]. Talmud [485, 412]. Tan [756]. t’ang [952]. Tanksukname [52]. Tanksukname-I-Ilhani [52]. Tannehill [521]. Tannery [134, 812, 363, 708]. taoiste [200]. Tarih [925]. Tarski [254]. Tartarotti [832]. Taylor [168, 65, 417]. Teach [449]. Teacher [682]. Teachers [627, 947]. Teaching [898, 716, 794]. Technical [43]. technique [864]. Technology [21, 16, 606]. Telescope [390]. Telescopes [904]. Temkin [622]. Temperaments [355]. Tempo [669]. Tempore [267, 268]. temps [166]. Tenney [310]. Tennyson [501]. tenu [762]. teoría [651]. Terence [381]. Terminology [605]. terminus [110]. Terra [483]. terre [570]. Territorial [401]. Test [489]. Testament [314]. Testimonies [660]. Tests [240]. Tetrabiblos [509, 289]. Texas [780]. Text [726, 884]. Textbooks [742]. Textes [198]. Texts [659]. Th [432]. thébaine [166]. Their [607, 170, 81]. Theodosius [344, 668]. Theognis [631]. Theon [611, 611]. théorie [209, 142]. Theories [326, 209]. Theory [223, 801, 222, 243, 950, 355, 504, 629, 694, 873]. These [434]. Thesis [726]. Things [856]. Third [262, 353, 525, 899, 947, 382]. Thirty [703]. Thirty-eight [703]. Thomas [944, 436, 205, 188, 419, 679, 548, 649, 391, 177, 582, 689, 833, 649].

Thompson [79, 195, 766, 378, 520]. Thomson [726]. Thorndike [612, 285, 909]. Thorp [944]. Those [745]. Thou [810]. Though [65]. Thought [40, 498, 594, 911, 661, 678, 103]. Thousand [783, 932]. three [353, 783, 771, 190, 122]. Thunberg [499]. Time [274, 267, 48, 649, 145]. Times [811, 20, 387, 765]. Tissue [504]. Tjerneld [663]. Tobias [33]. Today [886]. Togan [306]. Toledo [452]. Tolerance [217]. Tome [134, 198, 611, 67, 722]. Tomorrow [315]. Tools [170, 745]. toponymie [347]. Tops [129]. Torch [406]. Torrey [407]. Tortoise [325]. Tower [787]. Townson [325]. Trachtenberg [82]. Tract [540]. Tractatus [335]. Tracts [806, 539]. Trades [513]. Tradition [25, 547, 249]. Traditore [356]. traducciones [452]. Traduttore [356]. Tragic [64]. Traité [822]. Traités [663]. Transactions [190, 687]. Transfusion [788]. Transition [653]. Translation [3]. Translations [278, 539]. Translators [693]. Transmission [851]. travaux [430]. Travel [251]. Travels [613]. Travers [180]. Treatise [63, 764]. Treatises [450, 122]. Treatment [209, 6]. Trees [489]. Trembley [652, 652]. Trenchard [546]. Trends [485, 649]. tresor [910]. Trial [537]. Tribute [703]. Tridosa [340]. Trofim [668]. Troubled [886]. Truest [926]. Truth [734, 620]. Tswett [584]. Tuberculosis [6]. Tubes [489]. Tukan [592]. Turath [592]. Turcs [85]. Türklerinde [730]. Turks [730]. Turn [756]. Turnbull [813]. Turnips [20]. Turquet [558]. Twenty [23]. Twenty-Five [23]. Two [772, 671]. Tycho [811]. Tyler [168]. Tyndall [2]. Tyson [458]. Tysoniana [583]. Tzu [835, 835]. Tzu-liu-ching [835, 835].

U [52, 772]. Ugaritic [819]. Ukers [932]. Ulich [783]. Ulum [52]. Ulum-U-Funun-I-Khatai [52]. Unanue [272]. Understanding [731, 30, 620]. Unicorn [930]. Unified [291]. Union [269, 830, 667]. unique [773]. United [944, 140, 738, 290, 41, 312, 596, 520, 457]. Universal [747, 834]. Universe [149]. universelle [66]. University [316, 368]. Unnoted [222]. Unpublished [726]. unserer [104]. Unsolved [94]. Untersuchung [858]. Upham [135]. Upson [451]. Uranium [749]. Urdang [219]. Use [244, 331, 749, 804, 695, 400, 135, 538]. Uzuncarsilioglu [119].

V [206, 858, 17, 657, 125, 770]. Vagbhata [340]. Validi [306]. Valla [222]. Vallentin [64]. Vallicrosa [13, 452]. Vannoccio [449]. Vao [487]. Variability [668]. Variable [283]. variae [118]. Vasco [209]. Vaticana [621, 621]. Vaud [68]. Vázquez [451]. ve [119, 925]. Velocity [160, 5]. Ver [273]. Verborgen [815]. Verdoorn [615, 670]. veritate [19]. vérité [67]. Vernacular [140]. Verne [409]. Vers [885]. Version [389]. versus [302, 940]. Vesalius [545, 721]. Vetenskaps [153]. Vetenskapsakademiens [125]. Vetus [692]. Victor [291, 213]. Vida [621]. Vienna [464]. Viewpoint [474, 488]. Views [790]. VII [560]. VIII [634, 44]. Vilhjalmur [94, 782, 418]. Vinci [64, 511, 511, 511]. Vindication [496]. Virdung [386]. Virginia [477]. Visierkunst [892]. Vision [209]. Visorandi [892]. Visual [459]. vital [200]. Vivificat [388]. Vogelstein [369]. Vol [45, 105, 456, 459, 411, 647, 341, 940, 572]. Volume [106, 295, 319, 466, 554, 578, 624, 684, 700, 740, 785, 840, 960, 310, 852, 102, 657, 48, 2, 108, 155, 221, 298, 321, 262, 453, 385, 423, 440, 353, 468, 496, 525, 580, 627, 703, 787, 742, 843, 890, 947]. Vore [875]. Vorzeit [877]. Voyage [14]. Voyages [927]. vs [777].

W [676, 136, 168, 291, 726, 206, 234, 248, 454, 485, 64, 144, 152, 30, 38, 256, 290, 61, 289, 453, 877]. Waard [762, 812]. Wachsmuth [104]. Waddell [289]. Wade [315]. Waele [761]. Wagner [681]. Wahshiya [238]. wal [592]. wal-falak [592]. Walcott [185]. Walkington [582]. Walter [27, 381, 614, 776]. Walzer [610]. Want [933]. War [886, 401, 571, 596, 436]. Wargentin [125]. Warner [435, 676]. Warren [820]. Was [73]. Washington [10, 940, 88]. Water [634, 468, 754]. Waters [726]. Watson [30, 28]. Way [349, 199, 614]. Wayman [410]. Weather [113]. Weaver [820]. Webb [862]. Webber [572]. Weeks [141]. Wehrot [77]. Weinberg [911]. Weinberger [940, 940]. Weinstock [728]. Welch [419]. Welfare [258]. Weltordnung [863]. Wentworth [195, 766, 378]. Werken [815, 814]. Werkwoth [833]. Werner [133, 514, 662, 122]. West [873, 451, 671, 303]. Western [381, 779, 346, 402, 688]. Westfall [79, 520]. Wheats [658]. Wheeler [726]. wheels [754]. Wherry [74]. Which [607]. White [102, 251, 251]. Whitehead [103, 574]. Whittaker [399]. Who [315, 692]. Wholiness [205]. Whyte [619]. Wilbraham [212]. Wilbur [677]. Wilde [209]. Wilhelm [125, 287]. Wilkins [747]. Wilkinson [31]. Willard [726, 944, 429]. William [435, 250, 417, 613, 726, 139, 140, 661, 95, 36, 120, 146, 448, 886, 37, 131, 187, 193, 361, 596, 932, 396, 74, 512, 419]. Williams [781, 733]. Willibald [340]. Willy [135, 930]. Wilmer [189]. Wilson [140, 194, 460, 290, 489]. Winslow [551]. Winter [368]. Winthrop [802]. Wintner [375]. Wisdom [783, 402, 63]. Wissenschaften [307]. Wissenschaftliche [861]. Witchcraft [41, 380]. within [95]. Wittfogel [921]. Wives [549]. Wolf [940, 21]. Wolfson [903]. Wood [510]. Woodger [291]. Woodpecker [36]. Woods [618]. Woodward [364]. Word [332]. Work [607, 796, 726, 810, 17, 13, 848, 72, 879]. Workers [189]. Works [546, 60, 41, 208, 511, 278, 636, 188]. World [121, 886, 878, 675, 73, 933, 339, 646, 24, 196, 306]. Wright [591, 177, 315, 189, 591]. Writings [380, 178, 650]. Wrongly [906]. Wrote [692]. Wyrouboff [708].

XIV [647, 44]. XIX [654]. XV [134]. XVI [293, 760]. XVIIe [93]. XVIIIe [68, 816]. XXXI [45, 2]. XXXII [105, 48, 108]. XXXIII [155, 221, 262]. XXXIV [298, 321, 385, 423, 440, 353]. XXXV [468, 496, 525]. XXXVI [580].

Yale [726]. Yankee [857]. Year [300]. Yearbook [415, 457]. Years [783, 726, 712, 23, 737, 878, 353, 673, 932]. Yesterday [886]. York [41].

Yoshitomo [121]. Young [209, 825, 20]. Youssef [184].

Zachariae [183]. Zeit [616, 104]. Zeki [306]. Zellforschung [144]. Zeus. [197]. Zeydel [530]. Zilboorg [350]. Zimara [157]. Zinner [616]. Zoologist [488]. Zoology [930]. Zoroastre [25]. zum [908]. Zur [883, 763, 616, 723].

References

Anonymous:1940:FMa

[1] Anonymous. Front matter. Isis, 31(2):299–302, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225753.

Sarton:1940:SPVa

[2] George Sarton. Second preface to volume XXXI: Faraday to Tyndall. Isis, 31(2):303–304, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225754.

Silvette:1940:SBN

[3] Herbert Silvette. Some bibliographical notes and queries on Philemon Holland’s translation of the Regimen Sanitatis Salerni. Isis, 31(2):305– 308, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225755.

Stimson:1940:HHO

[4] Dorothy Stimson. Hartlib, Haak and Oldenburg: Intelligencers. Isis, 31 (2):309–326, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225756.

Cohen:1940:RFD

[5] I. Bernard Cohen. Roemer and the first determination of the velocity of light (1676). Isis, 31(2):327–379, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225757.

Pease:1940:SRD

[6] Arthur Stanley Pease. Some remarks on the diagnosis and treatment of tuberculosis in antiquity. Isis, 31(2):380–393, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225758.

Chatley:1940:LME

[7] Herbert Chatley. The lunar mansions in Egypt. Isis, 31(2):394–397, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225759.

Sarton:1940:RSB

[8] George Sarton. Remarks on the study of Babylonian mathematics. Isis, 31(2):398–404, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225760.

Gandz:1940:SBMa

[9] Solomon Gandz. Studies in Babylonian mathematics II. Conflicting interpretations of Babylonian mathematics. Isis, 31(2):405–425, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225761.

Anonymous:1940:CIW

[10] Anonymous. Carnegie Institution of Washington. Division of Historical Research. Section of the History of Science. Isis, 31(2):426–428, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225762.

Adams:1940:NC

[11] C. W. Adams, W. Burke-Gaffney, and Dirk J. Struik. Notes and correspondence. Isis, 31(2):429–432, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225763.

Kofoid:1940:BRBa

[12] Charles A. Kofoid. Book review: Natural History by Pliny; H. Rackham. Isis, 31(2):433–434, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225764.

Norman:1940:BRBa

[13] Daniel Norman. Book review: Rabbi Abraham Ibn Ezra’s Work in Scientific Astronomy by Joseph M. Millas Vallicrosa. Isis, 31(2):435–436, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225765.

Sarton:1940:BRBa

[14] George Sarton. Book review: The Second Voyage of Christopher Columbus from Cadiz to Hispaniola and the Discovery of the Lesser Antilles by Samuel Eliot Morison. Isis, 31(2):436–437, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225766.

Ashley-Montagu:1940:BRBa

[15] M. F. Ashley-Montagu. Book review: Abarbanel by Jacob S. Minkin. Isis, 31(2):438, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225767.

Jones:1940:BRB

[16] Richard F. Jones. Book review: Science, Technology, and Society in Seventeenth Century England by Robert K. Merton. Isis, 31(2):438–441, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225768.

Kofoid:1940:BRBb

[17] Charles A. Kofoid. Book review: The Life and Work of Robert Hooke. (Part V) Micrographia, 1665 by R. T. Gunther. Isis, 31(2):441–442, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225769.

Sarton:1940:BRBb

[18] George Sarton. Book review: Bayle’s Relations with England and the English by Léo Pierre Courtines. Isis, 31(2):442–444, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225770.

Hocking:1940:BRBa

[19] Richard Hocking. Book review: De veritate by Herbert of Cherbury; Meyrick H. Carre. Isis, 31(2):444–449, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225771.

Kofoid:1940:BRBc

[20] Charles A. Kofoid. Book review: Sheep and Turnips, Being the Life and Times of Arthur Young by Amelia Defries. Isis, 31(2):449–450, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225772.

Cohen:1940:BRBa

[21] I. Bernard Cohen. Book review: A History of Science, Technology, and Philosophy in the Eighteenth Century by A. Wolf. Isis, 31(2):450–451, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225773.

Viets:1940:BRB

[22] Henry R. Viets. Book review: The Open Mind: Elmer Ernest Southard, 1876–1920 by Frederick P. Gay. Isis, 31(2):451–453, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225774.

Kofoid:1940:BRBd

[23] Charles A. Kofoid. Book review: Twenty-Five Years of Health Progress by Louis I. Dublin; Alfred J. Lotka. Isis, 31(2):453–454, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225775.

Minkin:1940:BRBa

[24] Jacob S. Minkin. Book review: The Jew in the Medieval World by Jacob R. Marcus. Isis, 31(2):454–457, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225776.

Temkin:1940:BRB

[25] Owsei Temkin. Book review: Les Mages Hellénises. Zoroastre, Ostanès et Hystaspe d’après la Tradition Grecque by Joseph Bidez; Franz Cumont. Isis, 31(2):458–462, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225777.

Sarton:1940:BRBc

[26] Georges Sarton. Book review: Notices et Discours by Institut de France, Académie des sciences. Isis, 31(2):462–464, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225778.

Ashley-Montagu:1940:BRBb

[27] M. F. Ashley-Montagu. Book review: Background to Modern Science by Joseph Needham; Walter Pagel. Isis, 31(2):465–466, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225779.

M:1940:BRBa

[28] R. K. M. Book review: Scientists Are Human by David Lindsay Watson. Isis, 31(2):466–467, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225780.

Ashley-Montagu:1940:BRBc

[29] M. F. Ashley-Montagu. Book review: Science for the Citizen by Lancelot Hogben. Isis, 31(2):467–469, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225781.

Lenzen:1940:BRBa

[30] V. F. Lenzen. Book review: On Understanding Physics by W. H. Watson. Isis, 31(2):469–471, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225782.

Cohen:1940:BRBb

[31] I. Bernard Cohen. Book review: The Structure of Aristotelian Logic by James Wilkinson Miller. Isis, 31(2):471–473, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225783.

Cohen:1940:BRBc

[32] I. Bernard Cohen. Book review: A Semicentennial History of the American Mathematical Society, 1888–1938 by Raymond Clare Archibald. Isis, 31(2):473–475, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225784.

Sarton:1940:BRBd

[33] George Sarton. Book review: Number, the Language of Science by Tobias Dantzig. Isis, 31(2):475–476, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225785.

Geheniau:1940:BRBa

[34] Jules Géhéniau. Book review: Le système absolu classique et les mouvements réels by Augustin Sesmat. Isis, 31(2):476–478, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225786.

Geheniau:1940:BRBb

[35] Jules Géhéniau. Book review: Les Systèmes privilégiés de la Physique relativiste by Augustin Sesmat. Isis, 31(2):478–479, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225787.

Kofoid:1940:BRBe

[36] Charles A. Kofoid. Book review: The California Woodpecker and I by William Emerson Ritter. Isis, 31(2):479–480, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225788.

Leikind:1940:BRBa

[37] Morris C. Leikind. Book review: The History of Bacteriology by William Bulloch. Isis, 31(2):480–482, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225789.

M:1940:BRBb

[38] R. K. M. Book review: The Philosophy of the Act by George Herbert Mead; Charles W. Morris. Isis, 31(2):482–483, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225790.

Diller:1940:BRBa

[39] Aubrey Diller. Book review: Platonism Ancient and Modern by Paul Shorey. Isis, 31(2):483–485, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225791.

Ashley-Montagu:1940:BRBd

[40] M. F. Ashley-Montagu. Book review: Social Thought from Lore to Science by Harry Elmer Barnes; Howard Becker. Isis, 31(2):485–486, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225792.

Sarton:1940:BRBe

[41] George Sarton. Book review: List of Works in the New York Public Library Relating to Witchcraft in the United States by George F. Black; List of Works Relating to Witchcraft in Europe; A Calendar of Cases of Witchcraft in Scotland, 1510–1727. Isis, 31(2):486–487, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225793.

Sarton:1940:BRBf

[42] George Sarton. Book review: De studie van de oostersche talen in Nederland in de 18e en de 19e eeuw (The Study of Oriental Languages in the Netherlands in the Eighteenth and Nineteenth Centuries) by Jan Nat.

Isis, 31(2):487–488, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225794.

Sarton:1940:BRBg

[43] George Sarton. Book review: Report of the Consultative Committee on Secondary Education with Special Reference to Grammar Schools and Technical High Schools by Secondary Education. Isis, 31(2):488–490, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225795.

Sarton:1940:FEC

[44] George Sarton and Frances Siegel. Fifty-eighth critical bibliography of the history and philosophy of science and of the history of civilization (to October 1939 —with special reference to cent. VIII to XIV inclusive). Isis, 31(2):491–608, April 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225796.

Anonymous:1940:BTCa

[45] Anonymous. Brief table of contents of vol. XXXI (1939). Isis, 31(2): 609–611, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225797.

Anonymous:1940:BM

[46] Anonymous. Back matter. Isis, 31(2):??, April 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225798.

Anonymous:1940:FMb

[47] Anonymous. Front matter. Isis, 32(1):3–8, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226045.

Sarton:1940:PVX

[48] George Sarton. Preface to volume XXXII: Time is money. Isis, 32(1): 9–13, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226047.

Holmes:1940:PNS

[49] Urban T. Holmes, Jr. The position of the North Star ca. 1250. Isis, 32(1): 14–15, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226048.

Sayili:1940:AQH

[50] Aydin M. Sayili. Al Qarāfī and his explanation of the rainbow. Isis, 32 (1):16–26, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226049.

Johnson:1940:GMC

[51] M. C. Johnson. Greek, Moslem and Chinese instrument design in the surviving Mongol equatorials of 1279 A.D. Isis, 32(1):27–43, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226050.

Adnan:1940:TIU

[52] Abdulhak Adnan. Sur le Tanksukname-I-Ilhani der Ulum-U-Funun-I-Khatai. (French). Isis, 32(1):44–47, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226051.

Mendelsohn:1940:BNB

[53] Charles J. Mendelsohn. Bibliographical note on the “De Cifris” of Leone Battista Alberti. Isis, 32(1):48–51, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226052.

Haraszti:1940:DSS

[54] Zoltán Haraszti. Dr. Sarton on scientific incunabula. Isis, 32(1):52–62, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226053.

McColley:1940:CSD

[55] Grant McColley. Christopher Scheiner and the decline of neo-Aristotelianism. Isis, 32(1):63–69, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226054.

Ockenden:1940:GHL

[56] R. E. Ockenden. George Henry Lewes (1817–1878). Isis, 32(1):70–86, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226055.

Montagu:1940:KAA

[57] M. F. Ashley Montagu. Knowledge of the ape in antiquity. Isis, 32(1): 87–102, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226056.

Gandz:1940:SBMb

[58] Solomon Gandz. Studies in Babylonian mathematics iii: Isoperimetric problems and the origin of the quadratic equations. Isis, 32(1):103–115, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226057.

Sarton:1940:NCa

[59] George Sarton, H. W. Davies, W. F. Durand, W. Pagel, Bernard Drummond, Dirk J. Struik, C. D. Leake, Paul Schrecker, W. Ganzenmüller, Gudmund Björck, Jean Pelseneer, and Dietrich Mahnke. Notes and correspondence. Isis, 32(1):116–134, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226058.

Kofoid:1940:BRBf

[60] Charles A. Kofoid. Book review: The Genuine Works of Hippocrates by Francis Adams. Isis, 32(1):135–136, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226059.

Sarton:1940:BRBh

[61] George Sarton. Book review: Aristotle.–On the Heavens by W. K. C. Guthrie. Isis, 32(1):136, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226060.

Sarton:1940:BRBi

[62] George Sarton. Book review: Antécédents gréco-arabes de la psychophysique by Léon Gauthier. Isis, 32(1):136–138, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226061.

Gandz:1940:BRBa

[63] Solomon Gandz. Book review: The Beginning of Wisdom. An Astrological Treatise by Abraham Ibn Ezra by Raphael Levy; Francisco Cantera; Abraham Ibn Ezra. Isis, 32(1):138–141, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226062.

K:1940:BRB

[64] C. A. K. Book review: Leonardo da Vinci; The Tragic Pursuit of Perfection by Antonina Vallentin; E. W. Dickes. Isis, 32(1):141–143, July 1940.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226063.

Cohen:1940:BRBd

[65] I. Bernard Cohen. Book review: Galileo and the Freedom of Though by F. Sherwood Taylor. Isis, 32(1):143–145, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226064.

deSantiliana:1940:BRB

[66] George de Santiliana. Book review: Attraction universelle et religion naturelle chez quelques commentateurs anglais de Newton by Hélène Metzger. Isis, 32(1):145–148, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226065.

deSantillana:1940:BRBa

[67] George de Santillana. Book review: OEuvres complètes. Tome I. De la recherche de la vérité, livres I et II by Malebranche. Isis, 32(1):148– 150, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226066.

Sarton:1940:BRBj

[68] George Sarton. Book review: Médecine et santé dans le Pays de Vaud au XVIIIe siècle, 1675–1798 by Eugène Olivier. Isis, 32(1):150–153, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226067.

Guerlac:1940:BRB

[69] Henry Guerlac. Book review: Guyton-Morveau, chimiste et conventionnel (1737–1816) by Georges Bouchard. Isis, 32(1):153–156, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226068.

Pagel:1940:BRBa

[70] W. Pagel. Book review: Jean Paul Marat (1743–1793) by H. Rozbroj. Isis, 32(1):156–157, July 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226069.

Saunders:1940:BRB

[71] J. B. de C. M. Saunders. Book review: Memorialia Herman Boerhaave, Optimi Medici by Herman Boerhaave. Isis, 32(1):158, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226070.

Davis:1940:BRBa

[72] H. T. Davis. Book review: Karl Pearson. An Appreciation of Some Aspects of His Life and Work by E. S. Pearson. Isis, 32(1):158–164, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226071.

Zirkle:1940:BRBa

[73] Conway Zirkle. Book review: The World Was My Garden by David Fairchild. Isis, 32(1):164–165, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226072.

Kofoid:1940:BRBg

[74] Charles A. Kofoid. Book review: William B. Wherry, Bacteriologist by Martin Fischer. Isis, 32(1):165–166, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226073.

Montagu:1940:BRBa

[75] M. F. Ashley Montagu. Book review: Science in Progress by George A. Baitsell. Isis, 32(1):166–167, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226074.

Diller:1940:BRBb

[76] Aubrey Diller. Book review: La maturité de la pensée scientifique en Grèce by Abel Rey. Isis, 32(1):167–168, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226075.

Hartner:1940:BRB

[77] Willy Hartner. Book review: Wehrot und Arang by Josef Markwart. Isis, 32(1):168–174, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226076.

Montagu:1940:BRBb

[78] M. F. Ashley Montagu. Book review: Cylinder Seals by H. Frankfort. Isis, 32(1):174–175, July 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226077.

Sarton:1940:BRBk

[79] George Sarton. Book review: The Medieval Library by James Westfall Thompson. Isis, 32(1):175–177, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226078.

Sakanishi:1940:BRB

[80] Shio Sakanishi. Book review: Ezo tô bi shi kô (A Historical Study of Smallpox and Syphilis in the Island of Ezo) by Shinkichi Takahashi. Isis, 32(1):177–178, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226079.

Minkin:1940:BRBb

[81] Jacob S. Minkin. Book review: The Pharisees, the Sociological Background of Their Faith by Louis Finkelstein. Isis, 32(1):179–181, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226080.

Minkin:1940:BRBc

[82] Jocob S. Minkin. Book review: Jewish Magic and Superstition, a Study in Folk Religion by Joshua Trachtenberg. Isis, 32(1):182–183, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226081.

Minkin:1940:BRBd

[83] Jacob S. Minkin. Book review: Solomon Schechter (1847–1915). A Biography by Norman Bentwich. Isis, 32(1):184–185, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226082.

Gandz:1940:BRBb

[84] Solomon Gandz. Book review: Islam by Henry Massé; Halide Edib. Isis, 32(1):185–186, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226083.

Sarton:1940:BRBl

[85] George Sarton. Book review: La science chez les Turcs Ottomans by Abdulhak Adnan. Isis, 32(1):186–189, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226084.

Pogo:1940:BRB

[86] A. Pogo. Book review: The Inscriptions of Petén by Sylvanus Griswold Morley. Isis, 32(1):189–193, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226085.

Sarton:1940:BRBm

[87] George Sarton. Book review: Grosse Schweizer Forscher by Eduard Fueter. Isis, 32(1):193–197, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226086.

Davis:1940:BRBb

[88] H. T. Davis. Book review: Cooperation in Research by Carnegie Institution of Washington. Isis, 32(1):197–200, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226087.

Gerth:1940:BRB

[89] Hans H. Gerth. Book review: Freedom of Inquiry and Expression by Edward P. Cheney. Isis, 32(1):200–203, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226088.

Lenzen:1940:BRBb

[90] V. F. Lenzen. Book review: Die Methode der Physik by Hugo Dingler. Isis, 32(1):203–205, July 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226089.

Cohen:1940:BRBe

[91] I. Bernard Cohen. Book review: The Concepts of the Calculus by Carl B. Boyer. Isis, 32(1):205–210, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226090.

Cohen:1940:BRBf

[92] I. Bernard Cohen. Book review: History of Radio to 1926 by Gleason L. Archer. Isis, 32(1):210–211, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226091.

Sarton:1940:BRBn

[93] George Sarton. Book review: Catalogue des manuscrits alchimiques latins. I. Manuscrits des bibliothèques publiques de Paris antérieurs au XVIIe siècle by James Corbett. Isis, 32(1):211–212, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226092.

Johnson:1940:BRB

[94] Francis R. Johnson. Book review: Unsolved Mysteries of the Arctic by Vilhjalmur Stefansson. Isis, 32(1):212–214, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226093.

Gould:1940:BRB

[95] L. McK. Gould. Book review: The Discoveries of Antarctica within the American Sector as Revealed by Maps and Documents by William Herbert Hobbs. Isis, 32(1):214–218, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226094.

Hobbs:1940:BRB

[96] Wm. H. Hobbs. Book review: The Birth and Development of the Geological Sciences by Frank Dawson Adams. Isis, 32(1):218–220, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226095.

Romer:1940:BRB

[97] Alfred Sherwood Romer. Book review: Published Papers and Addresses of John Campbell Merriam. Isis, 32(1):220–222, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226096.

Kofoid:1940:BRBh

[98] Charles A. Kofoid. Book review: Les progrès récents de l’embryologie expérimentale by Maurice Caullery. Isis, 32(1):222–223, July 1940. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/226097.

Ginzburg:1940:BRB

[99] Benjamin Ginzburg. Book review: The Logic of Modern Psychology by Carroll C. Pratt. Isis, 32(1):223–225, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226098.

Montagu:1940:BRBc

[100] M. F. Ashley Montagu. Book review: The Meaning of the Humanities by T. M. Greene. Isis, 32(1):225–226, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226099.

Zirkle:1940:BRBb

[101] Conway Zirkle. Book review: Hunger and History: The Influence of Hunger on Human History by E. Parmalee Prentice. Isis, 32(1):227–229, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226100.

Merton:1940:BRB

[102] Robert K. Merton. Book review: Crime and the Man by Earnest Albert Hooton; The American Criminal: An Anthropological Study. Volume I, The Native White Criminal of Native Parentage. Isis, 32(1):229–238, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226101.

Montagu:1940:BRBd

[103] M. F. Ashley Montagu. Book review: Modes of Thought by Alfred North Whitehead. Isis, 32(1):239, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226102.

Pagel:1940:BRBb

[104] W. Pagel. Book review: Der Arzt in der Dichtung unserer Zeit by Bruno Wachsmuth. Isis, 32(1):239–240, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226103.

Anonymous:1940:BTCb

[105] Anonymous. Brief table of contents of vol. XXXII (1940). Isis, 32(1): ??, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226046.

Anonymous:1940:VI

[106] Anonymous. Volume information. Isis, 32(1):??, July 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226044.

Anonymous:1940:FMc

[107] Anonymous. Front matter. Isis, 32(2):241–244, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226240.

Sarton:1940:SPVb

[108] George Sarton. Second preface to volume XXXII: Reconstruction. Isis, 32(2):245–247, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226241.

DeSantillana:1940:EPS

[109] George De Santillana. Eudoxus and Plato. a study in chronology. Isis, 32(2):248–262, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226242.

Gandz:1940:HDN

[110] Solomon Gandz. Heron’s date. a new terminus ante quem (+150). Isis, 32(2):263–266, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226243.

Kibre:1940:AMA

[111] Pearl Kibre. The Alkimia Minor ascribed to Albertus Magnus. Isis, 32(2): 267–300, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226244.

Thorndike:1940:ANR

[112] Lynn Thorndike. Arabic numerals as represented in a Basel manuscript. Isis, 32(2):301–303, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226245.

Thorndike:1940:WRD

[113] Lynn Thorndike. A weather record for 1399–1406 A.D. Isis, 32(2):304– 323, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226246.

Arthos:1940:PDS

[114] John Arthos. Poetic diction and scientific language. Isis, 32(2):324–338, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226247.

Sarton:1940:NCb

[115] George Sarton, M. B. Singer, and A. Kaplan. Notes and correspondence. Isis, 32(2):339–345, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226248.

Sarton:1940:BR

[116] George Sarton. Isis report for 1939. Isis, 32(2):346–348, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226249.

S:1940:BRB

[117] G. S. Book review: Le synekdèmos d’Hiéroklès et l’Opuscule géographique de Georges de Chypre by Ernest Honigmann. Isis, 32(2):349–350, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226250.

White:1940:BRB

[118] Lynn White, Jr. Book review: Compositiones variae from Codex 490, Biblioteca Capitolare, Lucca, Italy: An Introductory Study by Rozelle Parker Johnson. Isis, 32(2):350–352, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226251.

Sayili:1940:BRB

[119] Aydin M. Sayili. Book review: Anadolu Beylikleri ve Akkoyunlu ve Karakoyunlu Devletleri (Anatolian States and the Ak Koyunlu and Kara Koyunlu Governments) by I. Hakki Uzuncars.ilioğlu. Isis, 32(2):352–354, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226252.

Kofoid:1940:BRBi

[120] Charles A. Kofoid. Book review: Engineers and Engineering in the Renaissance by William Barclay Parsons. Isis, 32(2):354–356, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226253.

Clack:1940:BRB

[121] Robert Wood Clack. Book review: Jurokuseiki sekaichizu jo no Nihon (Japan on World Maps of the Sixteenth Century) by Yoshitomo Okamoto. Isis, 32(2):356–358, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226254.

Norman:1940:BRBb

[122] Daniel Norman. Book review: Three Copernican Treatises (The Commentariolus of Copernicus, the Letter against Werner, the Narratio prima of Rheticus) by Edward Rosen. Isis, 32(2):358–359, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226255.

Sarton:1940:BRBo

[123] George Sarton. Book review: English Scholars by David C. Douglas. Isis, 32(2):359–360, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226256.

Sarton:1940:BRBp

[124] George Sarton. Book review: The Collected Letters of Antoni van Leeuwenhoek by Antoni van Leeuwenhoek. Isis, 32(2):360–361, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226257.

Sarton:1940:BRBq

[125] George Sarton. Book review: Pehr Wilhelm Wargentin, Kungl. Vetenskapsakademiens sekreterare och astronom, 1749–1783 by N. V. E. Nordenmark. Isis, 32(2):361–363, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226258.

deSantillana:1940:BRBb

[126] George de Santillana. Book review: The Kantian Philosophy of Space by Christopher Browne Garnett. Isis, 32(2):363–365, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226259.

Metzger:1940:BRB

[127] Hélène Metzger. Book review: Essai sur les origines intuitives du positivisme by Pierre Ducassé; Méthode et intuition chez Auguste Comte. Isis, 32(2):365–367, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226260.

Montagu:1940:BRBe

[128] M. F. Ashley Montagu. Book review: My Life by Havelock Ellis. Isis, 32(2):367–368, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226261.

Montagu:1940:BRBf

[129] M. F. Ashley Montagu. Book review: Runner of the Mountain Tops. The Life of Louis Agassiz by Mabel L. Robinson. Isis, 32(2):368–369, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226262.

Montagu:1940:BRBg

[130] M. F. Ashley Montagu. Book review: Emile Durkheim and His Sociology by Harry Alpert. Isis, 32(2):369–370, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226263.

Montagu:1940:BRBh

[131] M. F. Ashley Montagu. Book review: Life and Letters of Dr. William Beaumont by Jesse S. Myer. Isis, 32(2):371–372, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226264.

Cohen:1940:BRBg

[132] I. Bernard Cohen. Book review: Rutherford (1871–1937), Being the Life and Letters of the Rt. Hon. Lord Rutherford, O. M. by A. S. Eve. Isis, 32(2):372–375, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226265.

Diller:1940:BRBc

[133] Aubrey Diller. Book review: Paideia: The Ideals of Greek Culture by Werner Jaeger; Gilbert Highet. Isis, 32(2):375–376, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226266.

Sarton:1940:BRBr

[134] G. Sarton. Book review: Mémoires scientifiques. Tome XV. Correspondance by Paul Tannery; A. Dies; J. Pasquier; Pierre Louis. Isis, 32(2): 376–377, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226267.

Norman:1940:BRBc

[135] Daniel Norman. Book review: The Principle and Use of the Astrolabe by Willy Hartner; Arthur Upham Pope. Isis, 32(2):378, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226268.

Cohen:1940:BRBh

[136] I. Bernard Cohen. Book review: Matter and Light. The New Physics by Louis de Broglie; W. H. Johnston. Isis, 32(2):378–380, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226269.

Cohen:1940:BRBi

[137] I. Bernard Cohen. Book review: The Decline of Mechanism in Modern Physics by A. D’Abro. Isis, 32(2):380–382, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226270.

Cohen:1940:BRBj

[138] I. Bernard Cohen. Book review: Sparks, Lightning, Cosmic Rays. An Anecdotal History of Electricity by Dayton C. Miller. Isis, 32(2):382– 383, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226271.

Davis:1940:BRBc

[139] Tenney L. Davis. Book review: A History of Chemistry by Forris Jewett Moore; William T. Hall. Isis, 32(2):384, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226272.

Davis:1940:BRBd

[140] Tenney L. Davis. Book review: Catalogue of Latin and Vernacular Alchemical Manuscripts in the United States and Canada by William Jerome Wilson. Isis, 32(2):385–386, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226273.

Koenig:1940:BRB

[141] F. O. Koenig. Book review: Discovery of the Elements by Mary Elvira Weeks. Isis, 32(2):386–389, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226274.

Farber:1940:BRB

[142] Eduard Farber. Book review: Kurze Geschichte der Katalyse in Praxis und Theorie by Alwin Mittasch. Isis, 32(2):389–390, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226275.

Montagu:1940:BRBi

[143] M. F. Ashley Montagu. Book review: The Organism by Kurt Goldstein. Isis, 32(2):390–393, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226276.

Kofoid:1940:BRBj

[144] Charles A. Kofoid. Book review: Hundert Jahre Zellforschung by A. Aschoff; E. Kuster; W. J. Schmidt. Isis, 32(2):393–394, ???? 1940. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/226277.

Montagu:1940:BRBj

[145] M. F. Ashley Montagu. Book review: Pain, Sex and Time by Gerald Heard. Isis, 32(2):394–396, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226278.

Kofoid:1940:BRBk

[146] Charles A. Kofoid. Book review: An Analysis of the De generatione animalium of William Harvey by Arthur William Meyer; The Rise of Embryology by Arthur William Meyer. Isis, 32(2):396–398, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226279.

Kofoid:1940:BRBl

[147] Charles A. Kofoid. Book review: The Conquest of the Arctic by Louis Segal. Isis, 32(2):398–399, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226280.

Wohlstetter:1940:BRB

[148] Albert Wohlstetter. Book review: Charles Peirce’s Empiricism by Justus Buchler. Isis, 32(2):399–403, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226281.

Hocking:1940:BRBb

[149] Richard Hocking. Book review: The Universe and God by Robert Edward David Clark. Isis, 32(2):403–405, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226282.

Montagu:1940:BRBk

[150] M. F. Ashley Montagu. Book review: The Storehouse of Civilization by C. C. Furnas. Isis, 32(2):405–406, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226283.

Montagu:1940:BRBl

[151] M. F. Ashley Montagu. Book review: The Power of the Charlatan by Grete de Francesco. Isis, 32(2):406–408, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226284.

Leikind:1940:BRBb

[152] Morris C. Leikind. Book review: Man against Microbe by Joseph W. Bigger. Isis, 32(2):409, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226285.

Sarton:1940:BRBs

[153] George Sarton. Book review: Kungl. Svenska Vetenskaps Akademien by Bengt Hildebrand. Isis, 32(2):409–410, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226286.

S:1940:FNC

[154] G. S. and Frances Siegel. Fifty-ninth critical bibliography of the history and philosophy of science and of the history of civilization. Isis, 32(2): 411–502, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226287.

Sarton:1941:PVX

[155] George Sarton. Preface to volume XXXIII: To the Republic of Letters. Isis, 33(1):1–3, March 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330644.

Diller:1941:PPM

[156] Aubrey Diller. The parallels on the Ptolemaic maps. Isis, 33(1):4–7, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330645.

Tallmadge:1941:PMM

[157] G. Kasten Tallmadge. The perpetual motion machine of Mark Antony Zimara. Isis, 33(1):8–14, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330646.

Montagu:1941:SPR

[158] M. F. Ashley Montagu. A spurious portrait of Robert Hooke (1635–1703). Isis, 33(1):15–17, March 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330647.

Cohen:1941:ADB

[159] I. Bernard Cohen. Anquetil-Duperron, Benjamin Franklin, and Ezra Stiles. Isis, 33(1):17–23, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330648.

Boyer:1941:EEV

[160] Carl B. Boyer. Early estimates of the velocity of light. Isis, 33(1):24– 40, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330649.

Sarton:1941:EBI

[161] George Sarton. The end of the Belgian Isis. Isis, 33(1):41–54, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330650.

Sarton:1941:QA

[162] George Sarton, Loren C. MacKinney, and I. Bernard Cohen. Queries and answers. Isis, 33(1):55–63, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330651.

Armstrong:1941:NC

[163] Eva V. Armstrong. Notes and correspondence. Isis, 33(1):63–67, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330652.

Durand:1941:BRB

[164] Dana B. Durand. Book review: Le Livre de Ethiques d’Aristote by Maistre Nicole Oresme. Isis, 33(1):68–71, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330653.

Larsell:1941:BRB

[165] O. Larsell. Book review: Atlantica by Olaus Rudbeck. Isis, 33(1):71, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330654.

Sarton:1941:BRBa

[166] George Sarton. Book review: Les dessins ébauchés de la nécropole thébaine (au temps du nouvel empire) by Marcelle Baud. Isis, 33(1):71–73, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330655.

Kofoid:1941:BRBa

[167] C. A. Kofoid. Book review: Nature in Greek Poetry by George Soutar. Isis, 33(1):73–74, March 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330656.

Cohen:1941:BRBa

[168] I. Bernard Cohen. Book review: The Growth of Science. An Outline History by A. P. Rossiter; The March of Mind. A Short History of Science by F. Sherwood Taylor; A Short History of Science by W. T. Sedgwick; H. W. Tyler; R. P. Bigelow; Science since 1500. A Short History of Mathematics, Physics, Chemistry, Biology by H. T. Pledge. Isis, 33(1):74–79, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330657.

Lenzen:1941:BRB

[169] V. F. Lenzen. Book review: The Philosophy of Physical Science by Arthur Eddington. Isis, 33(1):79–80, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330658.

Farber:1941:BRB

[170] E. Farber. Book review: The Tools of the Chemist, Their Ancestry and American Evolution by Ernest Child. Isis, 33(1):80–81, March 1941. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/330659.

Zirkle:1941:BRBa

[171] Conway Zirkle. Book review: A History of Agriculture in Europe and America by N. S. B. Gras. Isis, 33(1):81–82, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330660.

Montagu:1941:BRBa

[172] M. F. Ashley Montagu. Book review: Historian and Scientist by Gaetano Salvemini. Isis, 33(1):82–83, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330661.

Sarton:1941:SCB

[173] George Sarton and Frances Siegel. Sixtieth critical bibliography of the history and philosophy of science and of the history of civilization (to October 1940). Isis, 33(1):84–180, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330662.

Viets:1941:ADa

[174] Henry R. Viets. Administrative documents. Isis, 33(1):181–186, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330663.

Anonymous:1941:FMa

[175] Anonymous. Front matter. Isis, 33(1):??, March 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330643.

Singer:1941:CGB

[176] Dorothea Waley Singer. The cosmology of Giordano Bruno (1548–1600). Isis, 33(2):187–196, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330740.

Gushee:1941:TWD

[177] Vera Gushee. Thomas Wright of Durham, astronomer. Isis, 33(2):197– 218, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330741.

Krakeur:1941:MWD

[178] Lester Gilbert Krakeur, Raymond Leslie Krueger, and Diderot. The mathematical writings of Diderot. Isis, 33(2):219–232, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330742.

Andrews:1941:SAL

[179] Alfred C. Andrews. The Silphium of the Ancients: A lesson in crop control. Isis, 33(2):232–236, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330743.

Pelseneer:1941:ASD

[180] Jean Pelseneer. Aspect statistique du progres des sciences en Belgique, À travers les siècles. Isis, 33(2):237–242, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330744.

Elwes:1941:QA

[181] Hugh G. Elwes and G. S. Queries and answers. Isis, 33(2):243–245, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330745.

Anonymous:1941:NC

[182] Anonymous. Notes and correspondence. Isis, 33(2):245, June 1941. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/330746.

Sarton:1941:BRBb

[183] George Sarton. Book review: Opera philosophica fragmentaque quae supersunt by Abi Bakr Mohammadi Filii Zachariae Raghensis; Paulus Kraus. Isis, 33(2):246–248, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330747.

Sarton:1941:BRBc

[184] George Sarton. Book review: La physiognomonie arabe et le Kitab al-firasa de Fakhr al-din al-Razi by Youssef Mourad. Isis, 33(2):248–249, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330748.

Zirkle:1941:BRBb

[185] Conway Zirkle. Book review: The Badianus Manuscript by Emily Walcott Emmart; Joannes Badianus. Isis, 33(2):249–251, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330749.

Cohen:1941:BRBb

[186] I. Bernard Cohen. Book review: The Photismi de lumine of Maurolycus. A Chapter in Late Medieval Optics by Henry Crew. Isis, 33(2):251–253, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330750.

Montagu:1941:BRBb

[187] M. F. Ashley Montagu. Book review: Isaac Newton by John William Navin Sullivan. Isis, 33(2):253–254, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330751.

S:1941:BRBa

[188] G. S. Book review: Cotton Mather, a Bibliography of His Works by Thomas James Holmes; Cotton Mather. Isis, 33(2):254–260, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330752.

Sarton:1941:BRBd

[189] George Sarton. Book review: De morbis artificum Bernardini Ramazzini diatriba. Diseases of Workers by Bernardino Ramazzini; Wilmer Cave Wright. Isis, 33(2):260–261, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330753.

Montagu:1941:BRBc

[190] M. F. Ashley Montagu. Book review: Three Transactions on the Cerebrum by Emmanuel Swedenborg. Isis, 33(2):262–263, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330754.

C:1941:BRB

[191] I B. C. Book review: Profile of Genius. “Poor Richard Pamphlets” by Franklin; Nathan G. Goodman. Isis, 33(2):263–264, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330755.

M:1941:BRBa

[192] M. F. A. M. Book review: O. C. Marsh Pioneer in Paleontology by Charles Schuchert; Clara M. LeVene. Isis, 33(2):264, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330756.

Romer:1941:BRB

[193] Alfred S. Romer. Book review: Some Memories of a Palaeontologist by William Berryman Scott. Isis, 33(2):264–267, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330757.

Montagu:1941:BRBd

[194] M. F. Ashley Montagu. Book review: To the Finland Station by Edmund Wilson. Isis, 33(2):267–269, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330758.

Sarton:1941:BRBe

[195] George Sarton. Book review: Science and the Classics by D’Arcy Wentworth Thompson. Isis, 33(2):269–270, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330759.

Montagu:1941:BRBe

[196] M. F. Ashley Montagu. Book review: Science and Politics in the Ancient World by Benjamin Farrington. Isis, 33(2):270–273, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330760.

Montagu:1941:BRBf

[197] M. F. Ashley Montagu. Book review: Zeus. A Study in Ancient Religion by Arthur Bernard Cook. Isis, 33(2):273–274, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330761.

Sarton:1941:BRBf

[198] George Sarton and A. Diller. Book review: Anecdota atheniensia et alia. Tome II: Textes grecs relatifs a l’histoire des sciences by Armand Delatte. Isis, 33(2):274–277, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330762.

Sarton:1941:BRBg

[199] George Sarton. Book review: The Chinese Way in Medicine by Edward Hicks Hume. Isis, 33(2):277–278, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330763.

S:1941:BRBb

[200] G. S. Book review: Les procédés de “nourrir le principe vital” dans la religion taoiste ancienne by Henri Maspero. Isis, 33(2):278–279, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330764.

Gandz:1941:BRBa

[201] Solomon Gandz. Book review: The Other Side of the Jordan by Nelson Glueck. Isis, 33(2):279–281, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330765.

Gandz:1941:BRBb

[202] Solomon Gandz. Book review: The Kosher Code of the Orthodox Jew by S. I. Levin; Edward A. Boyden. Isis, 33(2):282–285, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330766.

Montagu:1941:BRBg

[203] M. F. Ashley Montagu. Book review: Literature and Science by Grant Mc-Colley. Isis, 33(2):285–287, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330767.

Sarton:1941:BRBh

[204] George Sarton. Book review: The Record of the Royal Society of London for the Promotion of Natural Knowledge. Isis, 33(2):287–288, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330768.

Montagu:1941:BRBh

[205] M. F. Ashley Montagu. Book review: Hunger for Wholiness by Thomas H. Howells. Isis, 33(2):288–289, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330769.

Friedman:1941:BRB

[206] Bernard Friedman. Book review: Mathematical Logic by W. V. Quine. Isis, 33(2):289–291, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330770.

Cohen:1941:BRBc

[207] I. Bernard Cohen. Book review: The Development of Mathematics by E. T. Bell. Isis, 33(2):291–293, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330771.

Sarton:1941:BRBi

[208] George Sarton. Book review: Bibliography of Mathematical Works Printed in America through 1850 by Louis C. Karpinski. Isis, 33(2):293–294, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330772.

Cohen:1941:BRBd

[209] I. Bernard Cohen. Book review: Storia della luce by Vasco Ronchi; The History and Present State of Discoveries Relating to Vision, Light, and Colours by Joseph Priestley; Geschichte der Optik by Emil Wilde; Die Prinzipien der physikalischen Optik, historisch und erkenntnispsychologisch entwickelt (The Principles of Physical Optics, an Historical and Philosophical Treatment) by Ernst Mach; J. S. Anderson; A. F. A. Young; Geschichte der Optik by Edmund Hoppe; Les théories sur la nature de la lumière de Descartes à nos jours et l’évolution de la théorie physique by Ch. E. Papanastassiou. Isis, 33(2):294–296, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330773.

Montagu:1941:BRBi

[210] M. F. Ashley Montagu. Book review: Adventures of a Biologist by J. B. S. Haldane. Isis, 33(2):297–298, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330774.

Wright:1941:BRB

[211] John K. Wright. Book review: The Nature of Geography by Richard Hartshorne. Isis, 33(2):298–300, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330775.

Zirkle:1941:BRBc

[212] Conway Zirkle. Book review: The Englishman’s Food by J. C. Drummond; Ann Wilbraham. Isis, 33(2):300–301, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330776.

Kofoid:1941:BRBb

[213] C. A. Kofoid. Book review: Annual Review of Physiology by James Murray Luck; Victor E. Hall. Isis, 33(2):301–302, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330777.

Montagu:1941:BRBj

[214] M. F. Ashley Montagu. Book review: The Discovery of Man by Stanley Casson. Isis, 33(2):302–303, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330778.

M:1941:BRBb

[215] M. F. A. M. Book review: Race: Science and Politics by Ruth Benedict. Isis, 33(2):303–304, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330779.

Montagu:1941:BRBk

[216] M. F. Ashley Montagu. Book review: The Economic Life of Primitive Peoples by M. J. Herskovits. Isis, 33(2):304–305, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330780.

Montagu:1941:BRBl

[217] M. F. Ashley Montagu. Book review: Tolerance by Hendrik van Loon. Isis, 33(2):305–306, June 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330781.

Saunders:1941:BRB

[218] J. B. de C. M. Saunders. Book review: American Doctors of Destiny by Frank J. Jirka. Isis, 33(2):306–307, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330782.

Leikind:1941:BRBa

[219] Morris C. Leikind. Book review: History of Pharmacy, a Guide and a Survey by Edward Kremers; George Urdang. Isis, 33(2):307–308, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330783.

Anonymous:1941:FMb

[220] Anonymous. Front matter. Isis, 33(2):??, June 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330739.

Sarton:1941:SPV

[221] George Sarton. Second preface to volume XXXIII: Laplace’s religion. Isis, 33(3):309–312, September 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330785.

McColley:1941:GVU

[222] Grant McColley. George valla: An unnoted advocate of the geoheliocentric theory. Isis, 33(3):312–314, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330786.

Emery:1941:STE

[223] Clark Emery. Scientific theory in Erasmus Darwin’s “The Botanic Garden” (1789–91). Isis, 33(3):315–325, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330787.

Sarton:1941:MGS

[224] George Sarton. Mlle. Goichon’s studies on Avicennian metaphysics. Isis, 33(3):326–329, September 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330788.

Schrecker:1941:NED

[225] Paul Schrecker. Notes sur l’evolution du principe de la moindre action. Isis, 33(3):329–334, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330789.

Adams:1941:QA

[226] Frank D. Adams. Queries and answers. Isis, 33(3):335, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330790.

S:1941:NCa

[227] G. S., I. Bernard Cohen, Courtney Robert Hall, and A. Pogo. Notes and correspondence. Isis, 33(3):335–342, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330791.

Kofoid:1941:BRBc

[228] Charles A. Kofoid. Book review: Amos Eaton, Scientist and Educator, 1776–1842 by Ethel M. McAllister. Isis, 33(3):343, September 1941. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/330792.

Montagu:1941:BRBm

[229] M. F. Ashley Montagu. Book review: Darwinism in the English Novel by Leo J. Henkin. Isis, 33(3):343–344, September 1941. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330793.

Montagu:1941:BRBn

[230] M. F. Ashley Montagu. Book review: Quest. The Evolution of a Scientist by Leopold Infeld. Isis, 33(3):344–345, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330794.

Montagu:1941:BRBo

[231] M. F. Ashley Montagu. Book review: The Social Relations of Science by J. G. Crowther. Isis, 33(3):345–347, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330795.

Cohen:1941:BRBe

[232] I. Bernard Cohen. Book review: A History of Geometrical Methods by J. L. Coolidge. Isis, 33(3):347–350, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330796.

Zirkle:1941:BRBd

[233] Conway Zirkle. Book review: The American and His Food by Richard Osborn Cummings. Isis, 33(3):350–351, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330797.

Gandz:1941:BRBc

[234] Solomon Gandz. Book review: From the Stone Age to Christianity by W. F. Albright. Isis, 33(3):351–352, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330798.

Sarton:1941:SFC

[235] George Sarton and Frances Siegel. Sixty-first critical bibliography of the history and philosophy of science and of the history of civilization (to may 1941). Isis, 33(3):353–430, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330799.

Anonymous:1941:AD

[236] Anonymous. Administrative documents. Isis, 33(3):431–432, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330800.

Anonymous:1941:FMc

[237] Anonymous. Front matter. Isis, 33(3):??, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330784.

Darby:1941:IWM

[238] George O. S. Darby. Ibn Wahshiya in mediaeval Spanish literature. Isis, 33(4):433–438, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330620.

Hummel:1941:PHD

[239] A. W. Hummel. The printed herbal of 1249 A.D. Isis, 33(4):439–442, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330621.

Bayon:1941:SDH

[240] H. P. Bayon. The significance of the demonstration of the Harveyan circulation by experimental tests. Isis, 33(4):443–453, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330622.

Gordon:1941:MAA

[241] Maurice Bear Gordon. Medicine among the Ancient Hebrews. Isis, 33 (4):454–485, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330623.

Zirkle:1941:JCB

[242] Conway Zirkle. The jumar or cross between the horse and the cow. Isis, 33(4):486–506, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330624.

Miller:1941:TCS

[243] Wm. Marion Miller. The theory of concentric spheres. Isis, 33(4):507– 514, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330625.

Andrews:1941:AUL

[244] Alfred C. Andrews. Alimentary use of lovage in the classical period. Isis, 33(4):514–518, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330626.

Sheehan:1941:MLP

[245] Donal Sheehan. The Manchester literary and philosophical society. Isis, 33(4):519–523, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330627.

S:1941:NCb

[246] G. S. Notes and correspondence. Isis, 33(4):524–526, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330628.

Sarton:1941:BRBj

[247] George Sarton. Book review: Un glossaire de matiere medicale de Maimonide by Max Meyerhof. Isis, 33(4):527–529, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330629.

Gandz:1941:BRBd

[248] Solomon Gandz. Book review: Essays on Maimonides by Moses ben Maimon; Salo W. Baron. Isis, 33(4):529–532, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330630.

Kibre:1941:BRB

[249] Pearl Kibre. Book review: Princeps Concordiae: Pico della Mirandola and the Scholastic Tradition by Avery Dulles. Isis, 33(4):532–534, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330631.

Cohen:1941:BRBf

[250] I. Bernard Cohen. Book review: John and William Bartram, Botanists and Explorers by Ernest Earnest. Isis, 33(4):534–538, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330632.

Montagu:1941:BRBp

[251] M. F. Ashley Montagu. Book review: Extracts from the European Travel Journal of Lewis H. Morgan by Leslie A. White; Pioneers in American Anthropology: The Bandelier-Morgan Letters by Leslie A. White. Isis, 33 (4):538–539, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330633.

Nykl:1941:BRB

[252] A. R. Nykl. Book review: Huellas del Islam by Miguel Asin Palacios. Isis, 33(4):539–544, December 1941. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330634.

Montagu:1941:BRBq

[253] M. F. Ashley Montagu. Book review: Man on His Nature by Charles Sherrington. Isis, 33(4):544–545, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330635.

Symon:1941:BRB

[254] Keith R. Symon. Book review: Introduction to Logic and to the Methodology of Deductive Sciences by Alfred Tarski. Isis, 33(4):546–548, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330636.

Montagu:1941:BRBr

[255] M. F. Ashley Montagu. Book review: From Beast-Machine to Man-Machine by Leonora Cohen Rosenfield. Isis, 33(4):548–550, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330637.

Montagu:1941:BRBs

[256] M. F. Ashley Montagu. Book review: The Cambridge Bibliography of English Literature by F. W. Bateson. Isis, 33(4):550–551, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330638.

Leikind:1941:BRBb

[257] Morris C. Leikind. Book review: Progress in Medicine by Iago Galdston. Isis, 33(4):551–553, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330639.

Sarton:1941:BRBk

[258] George Sarton. Book review: Medicine and Human Welfare by Henry E. Sigerist. Isis, 33(4):553, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330640.

Kofoid:1941:BRBd

[259] Charles A. Kofoid. Book review: Magic in a Bottle by Milton Silverman. Isis, 33(4):553–555, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330641.

Viets:1941:ADb

[260] Henry R. Viets. Administrative documents. Isis, 33(4):555–556, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330642.

Anonymous:1941:FMd

[261] Anonymous. Front matter. Isis, 33(4):??, December 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/330619.

Sarton:1942:TPV

[262] George Sarton. Third preface to volume XXXIII: Brave Busbecq (1522– 1592). Isis, 33(5):557–575, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330665.

Neugebauer:1942:CAP

[263] O. Neugebauer. The chronology of the Aramaic papyri from Elephantine. Isis, 33(5):575–578, March 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330666.

vonErhardt:1942:ASR

[264] Rudolf von Erhardt and Erika von Erhardt-Siebold. Archimedes’ sandreckoner: Aristarchos and Copernicus. Isis, 33(5):578–602, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330667.

Sagui:1942:BED

[265] C. L. Sagui. Les bâtisseurs d’églises du moyen âge (I maestri comacini). (French/Italian) [The church builders of the Middle Ages (the Masters of Como)]. Isis, 33(5):602–609, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330668.

Buhler:1942:SCP

[266] Curt F. Buhler. Sixteenth-Century prognostications: Libri Impressi cum Notis Manuscriptis — Part II. Isis, 33(5):609–620, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330669.

Pagel:1942:JBV

[267] Walter Pagel. John Baptist van Helmont: De Tempore and the history of the biological concept of time. Isis, 33(5):621–623, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330670.

Weiss:1942:NGI

[268] Helene Weiss. Notes on the Greek ideas referred to in van Helmont: De Tempore. Isis, 33(5):624, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330671.

Wilson:1942:UCL

[269] William Jerome Wilson. The union catalog of the library of Congress. Isis, 33(5):625–629, March 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330672.

S:1942:NCa

[270] G. S. Notes and correspondence. Isis, 33(5):629–633, March 1942. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/330673.

Crew:1942:BRB

[271] Henry Crew. Book review: Benjamin Franklin’s Experiments by I. Bernard Cohen; Benjamin Franklin. Isis, 33(5):634–636, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330674.

Cohen:1942:BRBa

[272] I. Bernard Cohen. Book review: Observaciones sobre el clima de Lima, y su influencin ea los seres organizados, en especial el hombre by Jose Hipolito Unanue. Isis, 33(5):636–638, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330675.

Diller:1942:BRB

[273] Aubrey Diller. Book review: Les opuscules mathématiques de Didyme, Diophane, et Anthemius, suivis du fragment mathématique de Bobbio by Paul Ver Eecke. Isis, 33(5):639–640, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330676.

Montagu:1942:BRBa

[274] M. F. Ashley Montagu. Book review: Space, Time and Architecture by Sigfried Giedion. Isis, 33(5):640–641, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330677.

Leikind:1942:BRBa

[275] Morris C. Leikind. Book review: Germs and the Man by Justina Hill. Isis, 33(5):641–642, March 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330678.

Viets:1942:AD

[276] Henry R. Viets. Administrative documents. Isis, 33(5):643–648, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330679.

Anonymous:1942:FMa

[277] Anonymous. Front matter. Isis, 33(5):??, March 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330664.

Thorndike:1942:TWG

[278] Lynn Thorndike and Peter of Abano. Translations of works of Galen from the Greek by Peter of Abano. Isis, 33(6):649–653, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330710.

deSantillana:1942:NGS

[279] Giorgio Diaz de Santillana. New Galilean studies. Isis, 33(6):654–656, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330711.

Cohen:1942:BPC

[280] I. Bernard Cohen. The Compendium Physicae of Charles Morton (1627– 1698). Isis, 33(6):657–671, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330712.

Romer:1942:SHA

[281] Alfred Romer. The speculative history of atomic charges, 1873–1895. Isis, 33(6):671–683, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/330713.

Frank:1942:CIP

[282] Philipp Frank. Concerning an interpretation of positivism. Isis, 33(6): 683–687, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330714.

Bobrovnikoff:1942:DVS

[283] N. T. Bobrovnikoff. The discovery of variable stars. Isis, 33(6):687– 689, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330715.

Pease:1942:FFA

[284] Arthur Stanley Pease. Fossil fishes again. Isis, 33(6):689–690, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330716.

Durand:1942:MES

[285] Dana B. Durand. Magic and experimental science. The achievement of Lynn Thorndike. Isis, 33(6):691–712, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330717.

S:1942:NCb

[286] G. S. Notes and correspondence. Isis, 33(6):713–714, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330718.

Sarton:1942:BRBa

[287] George Sarton. Book review: Allgemeiner politischer und historischer Briefwechsel by Gottfried Wilhelm Leibniz. Isis, 33(6):715–716, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330719.

Montagu:1942:BRBb

[288] Montague Francis Ashley Montagu. Book review: Nietzsche by Crane Brinton. Isis, 33(6):716–718, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330720.

Sarton:1942:BRBb

[289] George Sarton. Book review: Manetho by Manetho; W. G. Waddell; Tetrabiblos by Ptolemy; Frank Egleston Robbins. Isis, 33(6):718–719, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330721.

S:1942:BRBa

[290] G. S. Book review: Census of Medieval and Renaissance Manuscripts in the United States and Canada by Seymour de Ricci; W. J. Wilson. Isis, 33(6):719–720, June 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330722.

Cohen:1942:BRBb

[291] I. Bernard Cohen. Book review: International Encyclopedia of Unified Science by Otto Neurath; Rudolf Carnap; Charles W. Morris; Niels Bohr; John Dewey; Bertrand Russell; Leonard Bloomfield; Victor F. Lenzen; Ernest Nagel; J. H. Woodger. Isis, 33(6):721–723, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330723.

Cohen:1942:BRBc

[292] I. Bernard Cohen. Book review: A Mathematician’s Apology by G. H. Hardy; Mathematics and the Imagination by Edward Kasner; James Newman. Isis, 33(6):723–725, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330724.

Sarton:1942:BRBc

[293] George Sarton. Book review: Index-Catalogue of the Library of the Surgeon General’s Office by Claudius F. Mayer; Bio-Bibliography of XVI. Century Medical Authors by Claudius F. Mayer. Isis, 33(6):726–727, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330725.

Durand:1942:AD

[294] Dana B. Durand. Administrative documents. Isis, 33(6):728–732, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330726.

Anonymous:1942:VIa

[295] Anonymous. Volume information. Isis, 33(6):733–739, June 1942. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/330708.

Anonymous:1942:FMb

[296] Anonymous. Front matter. Isis, 33(6):??, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330709.

Anonymous:1942:FMc

[297] Anonymous. Front matter. Isis, 34(1):1–2, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225988.

Sarton:1942:PVX

[298] George Sarton. Preface to volume XXXIV: O columns! Isis, 34(1):3– 4, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225989.

Neugebauer:1942:AA

[299] O. Neugebauer. Archimedes and Aristarchus. Isis, 34(1):4–6, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225990.

Thorndike:1942:OAT

[300] Lynn Thorndike. Other astronomical tables beginning in the year 1361. Isis, 34(1):6–7, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225991.

Arber:1942:NGM

[301] Agnes Arber. Nehemiah Grew (1641–1712) and Marcello Malpighi (1628– 1694): An essay in comparison. Isis, 34(1):7–16, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225992.

Emery:1942:SJH

[302] Clark Emery. “Sir” John Hill versus the Royal Society. Isis, 34(1):16– 20, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225993.

Renaud:1942:QCA

[303] H. P. J. Renaud. Quelques constructeurs d’astrolabes en occident musulman. (French) [Some constructors of astrolabes in the Muslim West]. Isis, 34(1):20–23, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225994.

Renaud:1942:QA

[304] H. P. J. Renaud, Eugene V. Prostov, L. Carrington Goodrich, and George Sarton. Queries and answers. Isis, 34(1):24–27, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225995.

Thorndike:1942:NC

[305] Lynn Thorndike, George Sarton, and Daniel Norman. Notes and correspondence. Isis, 34(1):27–30, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225996.

Sarton:1942:BRBd

[306] George Sarton. Book review: Biruni’s Picture of the World by A. Zeki Validi Togan. Isis, 34(1):31–32, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225997.

Klebs:1942:BRB

[307] A. C. Klebs. Book review: Geschichte der exakten Wissenschaften in der schweizerischen Aufklärung, 1680–1780 by Eduard Fueter. Isis, 34(1):32– 33, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225998.

Ginzburg:1942:BRB

[308] Benjamin Ginzburg. Book review: The Methodology of Pierre Duhem by Armand Lowinger. Isis, 34(1):33–34, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225999.

Sarton:1942:BRBe

[309] George Sarton. Book review: Les manuscrits arabes de l’Escurial by H. P. J. Renaud. Isis, 34(1):34–35, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226000.

Farber:1942:BRB

[310] Eduard Färber. Book review: The Chemistry of Powder and Explosives. Volume 1 by Tenney L. Davis. Isis, 34(1):35–36, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226001.

Zirkle:1942:BRBa

[311] Conway Zirkle. Book review: A Short History of the Plant Sciences by Howard S. Reed. Isis, 34(1):36, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226002.

Sarton:1942:BRBf

[312] George Sarton. Book review: Ornithologists of the United States Army Medical Corps by Edgar Erskine Hume. Isis, 34(1):36–38, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226003.

Graubard:1942:BRB

[313] Mark Graubard. Book review: How to Find Happiness and Keep It by M. F. Ashley Montagu. Isis, 34(1):38, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226004.

Gandz:1942:BRB

[314] Solomon Gandz. Book review: Introduction to the Old Testament by Robert H. Pfeiffer. Isis, 34(1):38–39, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226005.

Leikind:1942:BRBb

[315] Morris C. Leikind. Book review: The Man Who Lived for Tomorrow by Wade Wright Oliver. Isis, 34(1):39–40, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226006.

Sarton:1942:BRBg

[316] George Sarton. Book review: That They May Have Life. The Story of the American University of Beirut 1866–1941 by Stephen B. L. Penrose, Jr. Isis, 34(1):40–41, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226007.

Sarton:1942:SSC

[317] George Sarton and Frances Siegel. Sixty-second critical bibliography of the history and philosophy of science and of the history of civilization (to April 1942). Isis, 34(1):42–94, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226008.

Anonymous:1942:BMa

[318] Anonymous. Back matter. Isis, 34(1):??, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226009.

Anonymous:1942:VIb

[319] Anonymous. Volume information. Isis, 34(1):??, Summer 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225987.

Anonymous:1942:FMd

[320] Anonymous. Front matter. Isis, 34(2):95–96, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226208.

Sarton:1942:SPV

[321] George Sarton and Léo Lesquereux. Second preface to volume XXXIV: Lesquereux (1806–89). Isis, 34(2):97–108, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226209.

vonErhardt:1942:HPA

[322] Rudolf von Erhardt and Erika von Erhardt-Siebold. The helix in Plato’s astronomy. Isis, 34(2):108–110, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226210.

Delevsky:1942:LPM

[323] Jacques Delevsky. L’invention de la projection de Mercator et les enseignements de son histoire. (French) [The invention of the Mercator projection and the lessons of its history]. Isis, 34(2):110–117, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226211.

Olmsted:1942:SEJ

[324] John W. Olmsted. The scientific expedition of Jean Richer to Cayenne (1672–1673). Isis, 34(2):117–128, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226212.

Maluf:1942:RTR

[325] N. S. Rustum Maluf. Robert Townson and the respiratory movements of the tortoise. Isis, 34(2):128–132, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226213.

Galambos:1942:AOF

[326] Robert Galambos. The avoidance of obstacles by flying bats: Spallanzani’s ideas (1794) and later theories. Isis, 34(2):132–140, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226214.

Munsterberg:1942:BCB

[327] Margaret Munsterberg. The Bowditch Collection in the Boston Public Library. Isis, 34(2):140–142, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226215.

Koenig:1942:JMS

[328] F. O. Koenig. John Maxson Stillman (1852–1923). Isis, 34(2):142–146, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226216.

Montagu:1942:BM

[329] M. F. Ashley Montagu. Bronislaw Malinowski (1884–1942). Isis, 34(2): 146–150, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226217.

Romer:1942:EHA

[330] Alfred Romer. The experimental history of atomic charges, 1895–1903. Isis, 34(2):150–161, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226218.

Andrews:1942:AUH

[331] Alfred C. Andrews. Alimentary use of hoary mustard in the classical period. Isis, 34(2):161–162, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226219.

Spitzer:1942:OWJ

[332] Leo Spitzer. Origin of the word jumar. Isis, 34(2):163–164, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226220.

Anonymous:1942:NC

[333] Anonymous. Notes and correspondence. Isis, 34(2):165–166, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226221.

Durand:1942:BRB

[334] Dana B. Durand. Book review: Giovanni Marliani and Late Medieval Physics by Marshall Clagett. Isis, 34(2):166–168, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226222.

Clagett:1942:BRB

[335] Marshall Clagett. Book review: Note on the Tractatus Physici Falsely Attributed to Giovanni Marliani. Isis, 34(2):168, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226223.

Sayili:1942:BRB

[336] Aydin M. Sayili. Book review: The Palace School of Muhammad the Conqueror by Barnette Miller. Isis, 34(2):168–169, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226224.

Wilson:1942:BRB

[337] William Jerome Wilson. Book review: Admiral of the Ocean Sea: A Life of Christopher Columbus by Samuel Eliot Morison. Isis, 34(2):169– 172, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226225.

Johnson:1942:BRB

[338] Francis R. Johnson. Book review: Natural Science in German Romanticism by Alexander Gode-von Aesch. Isis, 34(2):172–173, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226226.

Hammond:1942:BRB

[339] Mason Hammond. Book review: The Social and Economic History of the Hellenistic World by M. Rostovtzeff. Isis, 34(2):173–174, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226227.

Sarton:1942:BRBh

[340] George Sarton. Book review: Vagbhat.a’s As.t.āṅgahr.dayasam .hita by Luise Hilgenberg; Willibald Kirfel; The Principle of Tridosa in Ayurveda by Dhirendra Nath Ray. Isis, 34(2):174–177, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226228.

Rentz:1942:BRB

[341] George Rentz. Book review: Raccolta di scritti editi e inediti. Vol. I: L’Arabia Saudiana (1938) by Carlo Alfonso Nallino; Maria Nallino. Isis, 34(2):177, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226229.

Cohen:1942:BRBd

[342] I. Bernard Cohen and Charles A. Kofoid. Book review: A Short History of Science to the Nineteenth Century by Charles Singer. Isis, 34(2):177– 180, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226230.

Lenzen:1942:BRB

[343] V. F. Lenzen. Book review: Between Physics and Philosophy by Philipp Frank. Isis, 34(2):180, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226231.

Zirkle:1942:BRBb

[344] Conway Zirkle. Book review: Genetics and the Origin of Species by Theodosius Dobzhansky. Isis, 34(2):181, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226232.

Montagu:1942:BRBc

[345] M. F. A. Montagu. Book review: Social Science Principles in the Light of Scientific Method by Joseph Mayer. Isis, 34(2):181–182, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226233.

Sarton:1942:BRBi

[346] George Sarton. Book review: Music in Western Civilization by Paul Henry Láng; Music in the Middle Ages by Gustave Reese. Isis, 34(2):182– 186, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226234.

S:1942:BRBb

[347] G. S. Book review: La toponymie française by Albert Dauzat. Isis, 34 (2):186–187, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226235.

Saunders:1942:BRB

[348] J. B. de C. M. Saunders. Book review: The March of Medicine. Isis, 34(2): 187, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226236.

Montagu:1942:BRBd

[349] M. F. Ashley Montagu. Book review: Native African Medicine by George Way Harley. Isis, 34(2):187–189, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226237.

Montagu:1942:BRBe

[350] M. F. Ashley Montagu. Book review: A History of Medical Psychology by Gregory Zilboorg. Isis, 34(2):189–190, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226238.

Anonymous:1942:BMb

[351] Anonymous. Back matter. Isis, 34(2):??, Autumn 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226239.

Anonymous:1943:FMa

[352] Anonymous. Front matter. Isis, 34(3):191–192, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225836.

Sarton:1943:TPV

[353] George Sarton. Third preface to volume XXXIV: The years “forty-three”. Isis, 34(3):193–195, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225837.

Mundy:1943:JG

[354] John Mundy. John of Gmunden. Isis, 34(3):196–205, Winter 1943. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/225838.

Sarton:1943:RTT

[355] George Sarton and Erika von Erhardt-Siebold. Remarks on the theory of temperaments. Isis, 34(3):205–208, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225839.

Koyre:1943:TTP

[356] Alexandre Koyre. Traduttore–traditore. A propos de Copernic et de Galilée. Isis, 34(3):209–210, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225840.

Neave:1943:ES

[357] E. W. J. Neave. The Epsom spring. Isis, 34(3):210–211, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225841.

Montagu:1943:KPH

[358] M. F. Ashley Montagu. Karl Pearson and the historical method in ethnology. Isis, 34(3):211–214, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225842.

Thorndike:1943:QA

[359] Lynn Thorndike. Queries and answers. Isis, 34(3):214, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225843.

vonErhardt:1943:NC

[360] Rudolf von Erhardt and Erika von Erhardt. Notes and correspondence. Isis, 34(3):214–215, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225844.

Saunders:1943:BRBa

[361] J. B. de C. M. Saunders. Book review: Hippocratic Medicine, Its Spirit and Method by William Arthur Heidel. Isis, 34(3):216, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225845.

Sarton:1943:BRBa

[362] George Sarton. Book review: Al-H .asan ibn al-Haitham by Mus.t.afa Naz.if Bey. Isis, 34(3):217–218, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225846.

Sarton:1943:BRBb

[363] George Sarton. Book review: Quadrivium de Georges Pachymère by Paul Tannery. Isis, 34(3):218–219, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225847.

Cohen:1943:BRBa

[364] I. Bernard Cohen. Book review: Ploughs and Politicks. Charles Read of New Jersey and His Notes on Agriculture. 1715–44 by Carl Raymond Woodward. Isis, 34(3):219–220, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225848.

Guerlac:1943:BRBa

[365] Henry Guerlac. Book review: Lazare Carnot, Republican Patriot by Huntley Dupre. Isis, 34(3):220–221, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225849.

Montagu:1943:BRBa

[366] M. F. Ashley Montagu. Book review: What Nietzsche Means by George Allen Morgan. Isis, 34(3):221, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225850.

Montagu:1943:BRBb

[367] M. F. Ashley Montagu. Book review: The Bertrand Russell Case by John Dewey; Horace M. Kallen. Isis, 34(3):221–222, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225851.

Sarton:1943:BRBc

[368] George Sarton. Book review: Papyri in the University of Michigan Collection by John Garrett Winter. Isis, 34(3):222–223, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225852.

Gandz:1943:BRBa

[369] Solomon Gandz. Book review: Rome by Hermann Vogelstein. Isis, 34(3): 223–224, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225853.

S:1943:BRBa

[370] G. S. Book review: Early Muslim Architecture by Keppel Archibald Cameron Creswell. Isis, 34(3):224–225, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225854.

Graubard:1943:BRBa

[371] Mark Graubard. Book review: The Road to Disappearance by Angie Debo. Isis, 34(3):225–226, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225855.

Gandz:1943:BRBb

[372] Solomon Gandz. Book review: Studies in the History of Science. Isis, 34 (3):226, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225856.

Montagu:1943:BRBc

[373] M. F. Ashley Montagu. Book review: The Promise of Scientific Humanism by Oliver L. Reiser. Isis, 34(3):226–229, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225857.

Symon:1943:BRB

[374] Keith R. Symon. Book review: Introduction to Semantics by Rudolf Carnap. Isis, 34(3):229, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225858.

C:1943:BRB

[375] I. B. C. Book review: The Analytical Foundations of Celestial Mechanics by Aurel Wintner. Isis, 34(3):230, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225859.

Romer:1943:BRB

[376] Alfred Romer. Book review: Sea Power in the Machine Age by Bernard Brodie. Isis, 34(3):230–231, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225860.

Leikind:1943:BRB

[377] Morris C. Leikind. Book review: The Advance of the Fungi by E. C. Large. Isis, 34(3):231–232, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225861.

Wrinch:1943:BRB

[378] Dorothy Wrinch. Book review: Growth and Form by D’Arcy Wentworth Thompson. Isis, 34(3):232–234, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225862.

Kofoid:1943:BRBa

[379] Charles A. Kofoid. Book review: The Retina by Stephen L. Polyak. Isis, 34(3):234–235, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225863.

Bainton:1943:BRB

[380] Roland H. Bainton. Book review: Materials toward a History of Witchcraft by Henry Charles Lea; Arthur C. Howland; Minor Historical Writings and Other Essays by Arthur C. Howland. Isis, 34(3):235–236, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225864.

Graubard:1943:BRBb

[381] Mark Graubard. Book review: The Destiny of Western Man by Walter Terence Stace. Isis, 34(3):236–237, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225865.

Sarton:1943:STC

[382] George Sarton and Frances Siegel. Sixty-third critical bibliography of the history and philosophy of science and of the history of civilization (to October 1942). Isis, 34(3):238–286, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225866.

Anonymous:1943:BMa

[383] Anonymous. Back matter. Isis, 34(3):??, Winter 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225867.

Anonymous:1943:FMb

[384] Anonymous. Front matter. Isis, 34(4):287–288, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225628.

Sarton:1943:FPVa

[385] George Sarton. Fourth preface to volume XXXIV: James Henry Breasted (1865–1935). The father of American Egyptology. Isis, 34(4):289–291, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225629.

Thorndike:1943:AVM

[386] Lynn Thorndike. Another Virdung manuscript. Isis, 34(4):291–293, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225630.

Stearns:1943:SSE

[387] Raymond Phineas Stearns. The scientific spirit in England in early modern times (c. 1600). Isis, 34(4):293–300, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225631.

Crew:1943:LOS

[388] Henry Crew. Littera Occidit — Spiritus Vivificat. (Latin) [The letter kills — the spirit quickens]. Isis, 34(4):300–301, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225632.

Wiener:1943:CNK

[389] Philip P. Wiener. A critical note on Koyré’s version of Galileo. Isis, 34(4):301–302, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225633.

Allen:1943:PCD

[390] Phyllis Allen. Problems connected with the development of the telescope (1609–1687). Isis, 34(4):302–311, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225634.

Brown:1943:MCL

[391] Harcourt Brown. The Mersenne correspondence: A lost letter by Thomas Hobbes. Isis, 34(4):311–312, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225635.

Hodgen:1943:SMH

[392] Margaret T. Hodgen. Sir Matthew Hale and the ’method’ of invention. Isis, 34(4):313–318, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225636.

Arber:1943:SCN

[393] Agnes Arber. A Seventeenth-Century naturalist: John Ray. Isis, 34(4): 319–324, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225637.

Moore:1943:GDS

[394] John Robert Moore. Goldsmith’s degenerate song-birds: An Eighteenth-Century fallacy in ornithology. Isis, 34(4):324–327, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225638.

Norman:1943:LMA

[395] Hilda L. Norman. Leopardi and the Machine Age. Isis, 34(4):327–337, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225639.

Adams:1943:WAM

[396] C. W. Adams. William Allen Miller and William Hallowes Miller (a note to the early history of spectroscopy). Isis, 34(4):337–339, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225640.

Thorndike:1943:FB

[397] Lynn Thorndike. Frederick Barry (1876–1943). Isis, 34(4):339–340, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225641.

Weeks:1943:MS

[398] Mary Elvira Weeks. Max Speter (1883–1942). Isis, 34(4):340–344, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225642.

Hardie:1943:PWH

[399] C. D. Hardie. Professor Whittaker and the history of physics. Isis, 34(4): 344–346, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225643.

Mason:1943:HUG

[400] Howard S. Mason. History of the use of graphic formulas in organic chemistry. Isis, 34(4):346–354, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225644.

Zirkle:1943:SET

[401] Conway Zirkle. Soil exhaustion, the territorial limitation of slavery, and the civil war. Isis, 34(4):355–359, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225645.

Coomaraswamy:1943:EWW

[402] Ananda K. Coomaraswamy. Eastern wisdom and Western knowledge. Isis, 34(4):359–363, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225646.

Eisler:1943:QA

[403] Robert Eisler. Queries and answers. Isis, 34(4):363–364, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225647.

Montagu:1943:NC

[404] M. F. Ashley Montagu. Notes and correspondence. Isis, 34(4):364–366, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225648.

vonErhardt-Siebold:1943:BRB

[405] Erika von Erhardt-Siebold. Book review: An English 13th Century Bestiary by Samuel A. Ives; Hellmut Lehmann-Haupt. Isis, 34(4):366–367, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225649.

Guerlac:1943:BRBb

[406] Henry Guerlac. Book review: Torch & Crucible. The Life and Death of Antoine Lavoisier by Sidney J. French. Isis, 34(4):367–368, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225650.

Zirkle:1943:BRBa

[407] Conway Zirkle. Book review: John Torrey, A Story of North American Botany by Andrew Denny Rodgers. Isis, 34(4):368–369, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225651.

Graubard:1943:BRBc

[408] Mark Graubard. Book review: Liebig and after Liebig by F. R. Moulton. Isis, 34(4):369–370, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225652.

Montagu:1943:BRBd

[409] M. F. Ashley Montagu. Book review: Jules Verne [1828–1905] by Kenneth Allott. Isis, 34(4):370–371, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225653.

Kofoid:1943:BRBb

[410] Charles A. Kofoid and G. Sarton. Book review: Edward Sylvester Morse by Dorothy G. Wayman. Isis, 34(4):371–373, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225654.

Montagu:1943:BRBe

[411] Ashley Montagu. Book review: The Cambridge Economic History of Europe. Vol 1 by J. H. Clapham; Eileen Power. Isis, 34(4):373, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225655.

Gandz:1943:BRBc

[412] Solomon Gandz. Book review: A Commentary on the Palestinian Talmud by Louis Ginzberg. Isis, 34(4):373–374, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225656.

M:1943:BRB

[413] M. F. A. M. Book review: The Diffusion of Science by Jesse Lee Bennett. Isis, 34(4):374–375, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225657.

S:1943:BRBb

[414] G. S. Book review: The Garment of God by John C. Merriam. Isis, 34 (4):375, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225658.

Walker:1943:BRB

[415] Helen M. Walker. Book review: The Second Yearbook of Research and Statistical Methodology by Oscar Krisen Buros. Isis, 34(4):375–377, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225659.

Johnson:1943:BRB

[416] Francis R. Johnson. Book review: The Star-Crossed Renaissance: The Quarrel about Astrology and Its Influence in England by Don Cameron Allen. Isis, 34(4):377–378, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225660.

Cohen:1943:BRBb

[417] I. Bernard Cohen. Book review: Physics, the Pioneer Science by Lloyd William Taylor. Isis, 34(4):378–379, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225661.

Sarton:1943:BRBd

[418] George Sarton. Book review: Iceland by Vilhjalmur Stefansson; Greenland. Isis, 34(4):379–380, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225662.

Saunders:1943:BRBb

[419] J. B. de C. M. Saunders. Book review: William Henry Welch and the Heroic Age of American Medicine by Simon Flexner; James Thomas Flexner. Isis, 34(4):381–382, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225663.

Sarton:1943:BRBe

[420] George Sarton. Book review: A Doctor for the People by Michael A. Shadid. Isis, 34(4):382, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225664.

Anonymous:1943:BMb

[421] Anonymous. Back matter. Isis, 34(4):??, Spring 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225665.

Anonymous:1943:FMc

[422] Anonymous. Front matter. Isis, 34(5):383–384, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225736.

Sarton:1943:FPVb

[423] George Sarton. Fifth preface to volume XXXIV: Aimé Bonpland (1773– 1858). Isis, 34(5):385–399, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225737.

Arber:1943:SB

[424] Agnes Arber. Spinoza and Boethius. Isis, 34(5):399–403, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225738.

Cohen:1943:FEH

[425] I. Bernard Cohen. Franklin’s experiments on heat absorption as a function of color. Isis, 34(5):404–407, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225739.

Goodrich:1943:CC

[426] L. Carrington Goodrich. Cotton in China. Isis, 34(5):408–410, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225740.

Vallicrosa:1943:NC

[427] J. Ma. Millás Vallicrosa, Lynn Thorndike, and M. F. Ashley Montagu. Notes and correspondence. Isis, 34(5):410–412, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225741.

Saunders:1943:BRBc

[428] J. B. de C. M. Saunders. Book review: The Doctors Mayo by Helen Clapesattle. Isis, 34(5):412–414, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225742.

Farber:1943:BRB

[429] Eduard Färber. Book review: Willard Gibbs by Muriel Rukeyser. Isis, 34(5):414–415, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225743.

Cramer:1943:BRB

[430] Marc Cramer. Book review: Souvenirs et travaux d’un chimiste by Amé Pictet. Isis, 34(5):415–417, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225744.

Klebs:1943:BRB

[431] A. C. Klebs. Book review: Festschrift für Jacques Brodbeck-Sandreuter by Karl Reucker. Isis, 34(5):417–418, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225745.

Kofoid:1943:BRBc

[432] Charles A. Kofoid. Book review: Biological Symposia by Jaques Cattell; Th. Dobzhansky. Isis, 34(5):418, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225746.

Kluckhohn:1943:BRB

[433] Clyde Kluckhohn. Book review: Man’s Most Dangerous Myth: The Fallacy of Race by M. F. Ashley Montagu. Isis, 34(5):419–420, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225747.

Gandz:1943:BRBd

[434] Solomon Gandz. Book review: What Mean These Stones? by Millar Burrows. Isis, 34(5):420–421, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225748.

Barber:1943:BRB

[435] Bernard Barber. Book review: The Social Life of a Modern Community by William Lloyd Warner; Paul S. Lunt; The Status System of a Modern Community. Isis, 34(5):421, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225749.

Graubard:1943:BRBd

[436] Mark Graubard. Book review: War as a Social Institution by Jesse D. Clarkson; Thomas C. Cochran. Isis, 34(5):422, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225750.

Sarton:1943:SFC

[437] George Sarton and Frances Siegel. Sixty-fourth critical bibliography of the history and philosophy of science and of the history of civilization (to June 1943). Isis, 34(5):423–462, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225751.

Anonymous:1943:BMc

[438] Anonymous. Back matter. Isis, 34(5):??, Summer 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225752.

Anonymous:1943:FMd

[439] Anonymous. Front matter. Isis, 34(6):463–464, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225889.

Sarton:1943:SPV

[440] George Sarton. Sixth preface to volume XXXIV: Defence of the History of Science (Philadelphia 1941). Isis, 34(6):465–466, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225890.

Thorndike:1943:RA

[441] Lynn Thorndike. Robertus Anglicus. Isis, 34(6):467–469, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225891.

Sarton:1943:FMC

[442] George Sarton. The feminine monarchie of Charles Butler, 1609. Isis, 34(6):469–472, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225892.

Sampson:1943:DFH

[443] Harriet Sampson. Dr. Faber and his celebrated cordial. Isis, 34(6):472– 496, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225893.

Epstein:1943:OEH

[444] Hans J. Epstein. The origin and earliest history of falconry. Isis, 34(6): 497–509, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225894.

Drabkin:1943:PCE

[445] I. E. Drabkin. Posidonius and the circumference of the Earth. Isis, 34 (6):509–512, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225895.

Sarton:1943:QA

[446] George Sarton. Queries and answers. Isis, 34(6):512–513, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225896.

S:1943:NC

[447] G. S. Notes and correspondence. Isis, 34(6):513, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225897.

Kofoid:1943:BRBd

[448] Charles A. Kofoid. Book review: T. Lucretii Cari De rerum natura by William Ellery Leonard; Stanley Barney Smith. Isis, 34(6):514, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225898.

Davis:1943:BRB

[449] Tenney L. Davis. Book review: The Pirotechnia of Vannoccio Biringuccio by Biringuccio; Cyril Stanley Smith; Martha Teach Gnudi. Isis, 34(6): 514–516, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225899.

Montagu:1943:BRBf

[450] M. F. Ashley Montagu. Book review: The Embryological Treatises of Hieronymous Fabricius of Aquapendente by Howard B. Adelmann. Isis, 34(6):516–517, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225900.

Wilson:1943:BRB

[451] William Jerome Wilson. Book review: Compendium and Description of the West Indies by Antonio Vázquez de Espinosa; Charles Upson Clark. Isis, 34(6):517–518, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225901.

Sarton:1943:BRBf

[452] George Sarton. Book review: Las traducciones orientales en los manuscritos de la Biblioteca Catedral de Toledo by José Maria Millás Vallicrosa. Isis, 34(6):518–519, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225902.

Sarton:1943:BRBg

[453] George Sarton. Book review: Eminent Chinese of the Ch’ing Period (1644–1912). Volume I by Arthur W. Hummel. Isis, 34(6):519–522, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225903.

Gandz:1943:BRBe

[454] Solomon Gandz. Book review: Archaeology and the Religion of Israel by W. F. Albright. Isis, 34(6):522–523, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225904.

Edsall:1943:BRB

[455] John T. Edsall. Book review: Biochemistry and Morphogenesis by Joseph Needham. Isis, 34(6):523–525, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225905.

Kofoid:1943:BRBe

[456] Charles A. Kofoid. Book review: Sex Hormones. Vol. IX, Biological Symposia by F. C. Koch; Philip E. Smith. Isis, 34(6):525, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225906.

Zirkle:1943:BRBb

[457] Conway Zirkle. Book review: United States Department of Agriculture Yearbook. Isis, 34(6):525–526, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225907.

Schultz:1943:BRB

[458] Adolph H. Schultz. Book review: Edward Tyson, M.D., F.R.S. 1650–1708 and the Rise of Human and Comparative Anatomy in England by M. F.

Ashley Montagu. Isis, 34(6):526–527, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225908.

Kofoid:1943:BRBf

[459] Charles A. Kofoid. Book review: Visual Mechanisms. Biological Symposia, Vol. 7 by Heinrich Klüver. Isis, 34(6):527–528, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225909.

Montagu:1943:BRBg

[460] M. F. Ashley Montagu. Book review: The Academic Man by Logan Wilson. Isis, 34(6):528–529, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225910.

Singer:1943:BRB

[461] Milton B. Singer. Book review: Philosophy in a New Key by Susanne K. Langer. Isis, 34(6):529–530, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225911.

Saunders:1943:BRBd

[462] J. B. de C. M. Saunders. Book review: A Short History of Cardiology by James B. Herrick. Isis, 34(6):530–531, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225912.

Leake:1943:BRB

[463] Chauncey D. Leake. Book review: Behind the Sulfa Drugs by Iago Galdston. Isis, 34(6):531, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225913.

Saunders:1943:BRBe

[464] J. B. de C. M. Saunders. Book review: British Medicine and the Vienna School by Max Neuburger. Isis, 34(6):531–532, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225914.

Pogo:1943:AD

[465] Alexander Pogo and Dana B. Durand. Administrative documents. Isis, 34(6):532–534, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225915.

Anonymous:1943:VI

[466] Anonymous. Volume information. Isis, 34(6):535–541, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225888.

Anonymous:1943:BMd

[467] Anonymous. Back matter. Isis, 34(6):??, Autumn 1943. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225916.

Sarton:1944:PVX

[468] George Sarton. Preface to volume XXXV: Faraday drinks a glass of water. Isis, 35(1):3, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/330681.

Muntner:1944:MBA

[469] Suessman Muntner. Maimonides’ book for Al-Fadil. Isis, 35(1):3–5, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330682.

Green:1944:JLF

[470] H. Gwynedd Green and H. J. J. Winter. John Landen, F.R.S. (1719–1790) — mathematician. Isis, 35(1):6–10, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330683.

Shryock:1944:NSH

[471] Richard H. Shryock. The need for studies in the history of American science. Isis, 35(1):10–13, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330684.

Means:1944:AAP

[472] Philip Ainsworth Means. Ancient American papermaking. Isis, 35(1): 13–15, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330685.

Coomaraswamy:1944:GE

[473] Ananda K. Coomaraswamy. Gradation and evolution. Isis, 35(1):15– 16, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330686.

Neuburger:1944:HSC

[474] Max Neuburger. An historical survey of the concept of nature from a medical viewpoint. Isis, 35(1):16–28, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330687.

Zeitlin:1944:QA

[475] Jacob Zeitlin and G. S. Queries and answers. Isis, 35(1):28–29, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330688.

Renaud:1944:NC

[476] H. P. J. Renaud, Lynn Thorndike, G. S., Alexandre Koyre, Earle O. Whittier, and R. Bennett Hovey. Notes and correspondence. Isis, 35(1):29–33, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330689.

Thorndike:1944:BRB

[477] Lynn Thorndike. Book review: The Philosophy of Marsilio Ficino by Paul Oskar Kristeller; Virginia Conant. Isis, 35(1):33–34, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330690.

Stimson:1944:BRB

[478] Dorothy Stimson. Book review: John Amos Comenius, That Incomparable Moravian by Matthew Spinka. Isis, 35(1):35, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330691.

M:1944:BRBa

[479] M. F. A. M. Book review: Barometric Pressure by Paul Bert. Isis, 35(1): 35–36, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330692.

Montagu:1944:BRBa

[480] M. F. Ashley Montagu. Book review: Haddon the Head-Hunter by A. Hingston Quiggin. Isis, 35(1):36–37, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330693.

Kofoid:1944:BRB

[481] Charles A. Kofoid. Book review: Joseph Grinnell’s Philosophy of Nature by Joseph Grinnell. Isis, 35(1):37, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330694.

Graubard:1944:BRB

[482] Mark Graubard. Book review: Marco Polo’s Precursors by Leonardo Olschki; The Land of the Great Image by Maurice Collis. Isis, 35(1):37– 39, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330695.

Montagu:1944:BRBb

[483] M. F. Ashley Montagu. Book review: Research on Early Man in Burma by Hellmut De Terra; Hallam L. Movius,; Edwin H. Colbert; J. Bequaert. Isis, 35(1):39–40, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330696.

Gandz:1944:BRBa

[484] Solomon Gandz. Book review: The Jews in Spain by Abraham A. Neuman. Isis, 35(1):40–41, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330697.

Gandz:1944:BRBb

[485] Solomon Gandz. Book review: The Jewish Community by Salo W. Baron; Marriage Laws in the Bible and in the Talmud by Louis M. Epstein; Major Trends in Jewish Mysticism by Gershom G. Scholem. Isis, 35(1):41–42, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330698.

Gandz:1944:BRBc

[486] Solomon Gandz. Book review: The Arabs by Philip K. Hitti. Isis, 35(1): 42–43, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330699.

Montagu:1944:BRBc

[487] M. F. Ashley Montagu. Book review: Stone Men of Malekula. Vao by John Layard. Isis, 35(1):43–44, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330700.

Zirkle:1944:BRBa

[488] Conway Zirkle. Book review: Systematics and the Origin of Species from the Viewpoint of a Zoologist by Ernst Mayr. Isis, 35(1):44–45, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330701.

Zirkle:1944:BRBb

[489] Conway Zirkle. Book review: Trees and Test Tubes: The Story of Rubber by Charles Morrow Wilson. Isis, 35(1):45–46, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330702.

Montagu:1944:BRBd

[490] M. F. Ashley Montagu. Book review: The Crisis of Our Age by P. A. Sorokin. Isis, 35(1):46–47, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330703.

Frances:1944:BRB

[491] Madeleine Frances. Book review: The Philosophy of G. E. Moore by Paul Arthur Schilpp. Isis, 35(1):47–48, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330704.

Mayer:1944:BRB

[492] Claudius F. Mayer. Book review: A Handbook of Medical Library Practice by Janet Doe. Isis, 35(1):48–51, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330705.

Anonymous:1944:AD

[493] Anonymous. Administrative documents. Isis, 35(1):51–52, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330706.

Sarton:1944:SFC

[494] George Sarton and Frances Siegel. Sixty-fifth critical bibliography of the history and philosophy of science and of the history of civilization (to December 1943). Isis, 35(1):53–94, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330707.

Anonymous:1944:FMa

[495] Anonymous. Front matter. Isis, 35(1):??, Winter 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330680.

Sarton:1944:SPV

[496] George Sarton. Second preface to volume XXXV vindication of Father Hell. Isis, 35(2):97–105, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330591.

Mullett:1944:GMS

[497] Charles F. Mullett. Gervase Markham: Scientific amateur. Isis, 35(2): 106–118, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/330592.

Galdston:1944:DMP

[498] Iago Galdston. Descartes and modern psychiatric thought. Isis, 35(2): 118–128, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/330593.

Svedelius:1944:CPT

[499] Nils Svedelius. Carl Peter Thunberg (1743–1828). Isis, 35(2):128–134, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330594.

Browne:1944:AHH

[500] C. A. Browne. Alexander von Humboldt as historian of science in Latin America. Isis, 35(2):134–139, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330595.

Emery:1944:BTA

[501] Clark Emery. The background of Tennyson’s “Airy Navies”. Isis, 35(2): 139–147, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/330596.

Guerlac:1944:GLB

[502] Henry Guerlac. George Lincoln Burr. Isis, 35(2):147–152, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330597.

Boyer:1944:FSD

[503] Carl B. Boyer. Fundamental steps in the development of numeration. Isis, 35(2):153–168, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330598.

Wilson:1944:CTD

[504] J. Walter Wilson. Cellular tissue and the dawn of the Cell theory. Isis, 35(2):168–173, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330599.

Buhler:1944:SMI

[505] Curt F. Buhler. Scientific and medical incunabula in American libraries. Isis, 35(2):173–175, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330600.

Merrill:1944:QA

[506] E. D. Merrill and George Sarton. Queries and answers. Isis, 35(2):176– 178, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330601.

Thomson:1944:NC

[507] William Thomson. Notes and correspondence. Isis, 35(2):179–180, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330602.

Sarton:1944:BRBa

[508] George Sarton. Book review: Generation of Animals by Aristotle; A. L. Peck. Isis, 35(2):181, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330603.

Sarton:1944:BRBb

[509] George Sarton. Book review: Ptolemy’s Tetrabiblos or Quadripartite by Ptolemy. Isis, 35(2):181–182, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330604.

Epstein:1944:BRB

[510] Hans J. Epstein. Book review: The Art of Falconry by Casey A. Wood; F. Marjorie Fyfe. Isis, 35(2):182–184, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330605.

Sarton:1944:BRBc

[511] George Sarton. Book review: The Literary Works of Leonardo da Vinci by Leonardo da Vinci; Jean Paul Richter; Irma A. Richter; The Notebooks of Leonardo da Vinci by Edward MacCurdy. Isis, 35(2):184–187, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330606.

Levine:1944:BRB

[512] Albert J. Levine. Book review: William James [1842–1910] by A. A. Roback. Isis, 35(2):187–188, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330607.

Saunders:1944:BRBa

[513] J. B. de C. M. Saunders. Book review: Exploring the Dangerous Trades by Alice Hamilton. Isis, 35(2):188, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330608.

Diller:1944:BRB

[514] Aubrey Diller. Book review: Paideia: The Ideals of Greek Culture by Werner Jaeger; Gilbert Highet. Isis, 35(2):188–189, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330609.

Sarton:1944:BRBd

[515] George Sarton. Book review: The Contribution of Holland to the Sciences by A. J. Barnouw; B. Landheer. Isis, 35(2):189–190, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330610.

Sarton:1944:BRBe

[516] George Sarton. Book review: Societa italiana per il progresso delle scienze. Anno 100 degrees dalla 1a riunione degli scienziati italiani. Un secolo di progresso scientifico italiano, 1839–1939. Isis, 35(2):190–191, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330611.

Montagu:1944:BRBe

[517] M. F. Ashley Montagu. Book review: The Scientific Life by John R. Baker. Isis, 35(2):191–192, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330612.

Zirkle:1944:BRBc

[518] Conway Zirkle. Book review: Evolution: The Modern Synthesis by Julian Huxley. Isis, 35(2):192–194, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330613.

Zirkle:1944:BRBd

[519] Conway Zirkle. Book review: Garden Islands of the Great East by David Fairchild. Isis, 35(2):194–195, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330614.

Zirkle:1944:BRBe

[520] Conway Zirkle. Book review: A History of Livestock Raising in the United States, 1607–1860 by James Westfall Thompson. Isis, 35(2):195, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330615.

Zirkle:1944:BRBf

[521] Conway Zirkle. Book review: Hurricanes by Ivan Ray Tannehill. Isis, 35(2):195–196, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330616.

Zirkle:1944:BRBg

[522] Conway Zirkle. Book review: Man’s Food, Its Rhyme or Reason by Mark Graubard. Isis, 35(2):196–197, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330617.

Berry:1944:BRB

[523] George D. W. Berry. Book review: Philosophical Essays in Honor of Edgar Arthur Singer, Jr by F. P. Clarke; M. C. Nahm. Isis, 35(2):197– 198, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330618.

Anonymous:1944:FMb

[524] Anonymous. Front matter. Isis, 35(2):??, Spring 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330590.

Sarton:1944:TPV

[525] George Sarton. Third preface to volume XXXV: Scripta Manent. Isis, 35(3):201–205, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330728.

Schmidt:1944:CLD

[526] Olaf H. Schmidt. The computation of the length of daylight in Hindu astronomy. Isis, 35(3):205–211, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330729.

Goodrich:1944:QA

[527] L. Carrington Goodrich and C. W. Adams. Queries and answers. Isis, 35(3):211–212, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330730.

Montagu:1944:NC

[528] M. F. Ashley Montagu. Notes and correspondence. Isis, 35(3):212–213, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330731.

Meyerhof:1944:BRB

[529] Max Meyerhof. Book review: Jabir ibn Hayyan. Contribution à l’histoire des idées scientifiques dans l’Islam by Paul Kraus. Isis, 35(3):213–217, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/330732.

M:1944:BRBb

[530] M. F. A. M. Book review: The Ship of Fools by Sebastian Brant; Edwin H. Zeydel. Isis, 35(3):217, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330733.

M:1944:BRBc

[531] M. F. A. M. Book review: Letters to His Son Lucien by Camille Pissarro; Lucien Pissarro; John Rewald; Lionel Abel. Isis, 35(3):217–218, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330734.

Sarton:1944:BRBf

[532] George Sarton. Book review: A Catalogue of the Medieval and Renaissance Manuscripts and Incunabula in the Boston Medical Library by James F. Ballard. Isis, 35(3):218–219, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330735.

Cohen:1944:BRB

[533] I. Bernard Cohen. Book review: What Is Mathematics? by Richard Courant; Herbert Robbins. Isis, 35(3):219–220, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330736.

Galdston:1944:BRB

[534] Iago Galdston. Book review: Civilization and Disease by Henry E. Sigerist. Isis, 35(3):220, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330737.

Sarton:1944:SSCa

[535] George Sarton and Frances Siegel. Sixty-sixth critical bibliography of the history and philosophy of science and of the history of civilization (to July 1944). Isis, 35(3):221–278, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330738.

Anonymous:1944:FMc

[536] Anonymous. Front matter. Isis, 35(3):??, Summer 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330727.

Honigmann:1944:TSA

[537] Ernest Honigmann. A trial for sorcery on August 22, A.D. 449. Isis, 35 (4):281–284, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330839.

Worrell:1944:QIL

[538] W. H. Worrell. Qusta Ibn Luqa on the use of the celestial globe. Isis, 35(4):285–293, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330840.

Thorndike:1944:LTA

[539] Lynn Thorndike. The Latin translations of the astrological tracts of Abraham Avenezra. Isis, 35(4):293–302, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330841.

Kibre:1944:ATA

[540] Pearl Kibre. An alchemical tract attributed to Albertus Magnus. Isis, 35(4):303–316, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330842.

Rufus:1944:KSM

[541] W. Carl Rufus and Celia Chao. A Korean star map. Isis, 35(4):316– 326, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330843.

Sarton:1944:G

[542] George Sarton. Grassmann — 1844. Isis, 35(4):326–330, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330844.

Miller:1944:QA

[543] Genevieve Miller, H. Bruce Collier, Erika von Erhardt-Siebold, and George Sarton. Queries and answers. Isis, 35(4):331–333, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330845.

Archibald:1944:NC

[544] R. C. Archibald, I. Bernard Cohen, and George Sarton. Notes and correspondence. Isis, 35(4):333–337, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330846.

Saunders:1944:BRBb

[545] J. B. de C. M. Saunders. Book review: A Bio-Bibliography of Andreas Vesalius by Harvey Cushing; The Harvey Cushing Collection of Books and Manuscripts by Harvey Cushing. Isis, 35(4):338–341, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330847.

Fulton:1944:BRB

[546] John F. Fulton. Book review: The Life and Works of The Honourable Robert Boyle by Louis Trenchard More. Isis, 35(4):341–342, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330848.

LeCorbeiller:1944:BRB

[547] P. Le Corbeiller. Book review: La tradition litteraire des Ideologues by Emile Caillet. Isis, 35(4):342–343, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330849.

Zirkle:1944:BRBh

[548] Conway Zirkle. Book review: Naturalist at Large by Thomas Barbour. Isis, 35(4):343–344, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330850.

Collier:1944:BRB

[549] H. Bruce Collier. Book review: Djin Ping Meh by Artur Kibat; Otto Kibat; King Ping Meh by Franz Kuhn; Chin P’ing Mei, the Adventurous History of Hsi Men and His Six Wives by Bernard Miall; Franz Kuhn; The Golden Lotus by Clement Egerton. Isis, 35(4):344–346, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330851.

Gandz:1944:BRBd

[550] Solomon Gandz. Book review: The Jews and Medicine by Harry Friedenwald. Isis, 35(4):346–347, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330852.

Saunders:1944:BRBc

[551] J. B. de C. M. Saunders. Book review: The Conquest of Epidemic Disease by Charles-Edward A. Winslow. Isis, 35(4):347–348, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330853.

Sarton:1944:SSCb

[552] George Sarton and Frances Siegel. Sixty-seventh critical bibliography of the history and philosophy of science and of the history of civilization. Isis, 35(4):349–362, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330854.

Pogo:1944:AD

[553] Alexander Pogo. Administrative documents. Isis, 35(4):363–368, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330855.

Anonymous:1944:VI

[554] Anonymous. Volume information. Isis, 35(4):369–373, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330837.

Anonymous:1944:FMd

[555] Anonymous. Front matter. Isis, 35(4):??, Autumn 1944. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330838.

Anonymous:1945:FM

[556] Anonymous. Front matter. Isis, 36(1):1–2, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225667.

Thorndike:1945:PLC

[557] Lynn Thorndike. Peter of Limoges on the Comet of 1299. Isis, 36(1): 3–6, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225668.

Thorndike:1945:FPT

[558] Lynn Thorndike. Franco de Polonia and the Turquet. Isis, 36(1):6–7, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/225669.

Desch:1945:FSM

[559] Cecil H. Desch. Francis Sidney Marvin, 1863–1943. Isis, 36(1):7–9, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225670.

Neugebauer:1945:SAA

[560] O. Neugebauer. Studies in ancient astronomy. VII. Magnitudes of lunar eclipses in Babylonian mathematical astronomy. Isis, 36(1):10–15, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225671.

Lindroth:1945:HSS

[561] Sten Lindroth. History of science in Sweden. Isis, 36(1):16–19, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225672.

Delevsky:1945:NPR

[562] Jacques Delevsky. Note sur la possibilité des répétitions cosmologiques. Isis, 36(1):19–21, October 1945. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225673.

Sarton:1945:SEC

[563] George Sarton and Frances Siegel. Sixty-eighth critical bibliography of the history and philosophy of science and of the history of civilization (to January 1945). Isis, 36(1):22–69, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225674.

Pettengill:1945:QA

[564] George E. Pettengill. Queries and answers. Isis, 36(1):69, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225675.

Renaud:1945:NC

[565] H. P. J. Renaud. Notes and correspondence. Isis, 36(1):69–70, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225676.

Graubard:1945:BRB

[566] Mark Graubard. Book review: The Land of Prester John by Elaine Sanceau. Isis, 36(1):70–71, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225677.

M:1945:BRB

[567] M. F. A. M. Book review: The Gobi Desert by Mildred Cable; Francesca French. Isis, 36(1):71, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225678.

S:1945:BRB

[568] G. S. Book review: The Bhagavad Gita by Franklin Edgerton. Isis, 36(1): 71–73, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225679.

LeCorbeiller:1945:BRB

[569] P. Le Corbeiller. Book review: Men of Science in America by Bernard Jaffe. Isis, 36(1):73–74, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225680.

Sarton:1945:BRBa

[570] George Sarton. Book review: Critique et géologie. Contribution à l’histoire des sciences de la terre (1882–1942) by Emmanuel de Margerie. Isis, 36 (1):74–75, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225681.

Zirkle:1945:BRBa

[571] Conway Zirkle. Book review: Food, War and the Future by E. Parmalee Prentice. Isis, 36(1):75–76, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225682.

Zirkle:1945:BRBb

[572] Conway Zirkle. Book review: The Citrus Industry. Vol. I, History, Botany and Breeding by Herbert John Webber; Leon Dexter Batchelor. Isis, 36(1): 76–77, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225683.

Sarton:1945:BRBb

[573] George Sarton. Book review: Squire. Memories of Charles Davies Sherborn [1861–1942] by John Roxbrough Norman. Isis, 36(1):77–78, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225684.

Roberts:1945:BRB

[574] Leo Roberts. Book review: The Philosophy of Alfred North Whitehead by Paul Arthur Schillp. Isis, 36(1):78–81, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225685.

Leake:1945:BRB

[575] Chauncey D. Leake and Philip Ainsworth Means. Book review: Aesculapius in Latin America by Aristides A. Moll. Isis, 36(1):81–83, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225686.

Wilson:1945:BRB

[576] William Jerome Wilson. Book review: The Scholar and the Future of the Research Library: A Problem and Its Solution by Fremont Rider. Isis, 36(1):83–86, October 1945. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225687.

Anonymous:1945:BM

[577] Anonymous. Back matter. Isis, 36(1):??, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225688.

Anonymous:1945:VI

[578] Anonymous. Volume information. Isis, 36(1):??, October 1945. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225666.

Anonymous:1946:FMa

[579] Anonymous. Front matter. Isis, 36(2):87–88, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225868.

Sarton:1946:PVX

[580] George Sarton. Preface to volume XXXVI: Confiteor. Isis, 36(2):89– 91, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225869.

Boyer:1946:ARL

[581] Carl B. Boyer. Aristotelian references to the Law of Reflection. Isis, 36(2): 92–95, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225870.

Mullett:1946:TWH

[582] Charles F. Mullett. Thomas Walkington and his “optick glasse”. Isis, 36(2):96–105, January 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225871.

Montagu:1946:T

[583] M. F. Ashley Montagu. Tysoniana. Isis, 36(2):105–108, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225872.

Zechmeister:1946:MTI

[584] L. Zechmeister. Mikhail Tswett — the inventor of chromatography. Isis, 36(2):108–109, January 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225873.

Brown:1946:GSB

[585] Harcourt Brown. George Sidney Brett. Isis, 36(2):110–114, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225874.

Goodrich:1946:EDF

[586] L. Carrington Goodrich and Fêng Chia-shêng. The early development of firearms in China. Isis, 36(2):114–123, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225875.

Farmer:1946:GEA

[587] Henry George Farmer. “Ghosts”. An excursus on Arabic musical bibliographies. Isis, 36(2):123–130, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225876.

Sarton:1946:NC

[588] George Sarton and Aubrey Diller. Notes and correspondence. Isis, 36 (2):130–134, January 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225877.

Sarton:1946:BRBa

[589] George Sarton. Book review: Jac. Berzelius Brev by Berzelius. Isis, 36 (2):134–135, January 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225878.

Zirkle:1946:BRBa

[590] Conway Zirkle. Book review: John Merle Coulter. Missionary in Science by Andrew Denny Rodgers. Isis, 36(2):135, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225879.

Cohen:1946:BRBa

[591] I. Bernard Cohen. Book review: The Wright Brothers. A Biography Authorized by Orville Wright by Fred C. Kelly. Isis, 36(2):136–140, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225880.

Sarton:1946:BRBb

[592] George Sarton. Book review: Turath al-‘arab al-‘ilmi fil-riyad.iyyat walfalak [The Scientific Heritage of the Arabs in Mathematics and Astronomy] by K. H. Tukan. Isis, 36(2):140–142, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225881.

Gurwitsch:1946:BRB

[593] Aron Gurwitsch. Book review: Society and Nature. A Sociological Inquiry by Hans Kelsen. Isis, 36(2):142–146, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225882.

Montagu:1946:BRB

[594] M. F. Ashley Montagu. Book review: Social Darwinism in American Thought, 1860–1915 by Richard Hofstadter. Isis, 36(2):146–147, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225883.

M:1946:BRB

[595] M. F. A. M. Book review: Man the Measure by Erich Kahler. Isis, 36(2): 147, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225884.

Shryock:1946:BRB

[596] Richard Harrison Shryock. Book review: Medical Education in the United States before the Civil War by William Frederick Norwood. Isis, 36(2): 147–149, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225885.

Zirkle:1946:BRBb

[597] Conway Zirkle. Book review: A Source Book of Agricultural Chemistry by Charles A. Browne. Isis, 36(2):149–150, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225886.

Anonymous:1946:BMa

[598] Anonymous. Back matter. Isis, 36(2):??, January 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225887.

Anonymous:1946:FMb

[599] Anonymous. Front matter. Isis, 36(3–4):151–152, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225602.

Sarton:1946:FDS

[600] George Sarton. Floating docks in the Sixteenth Century. Isis, 36(3–4): 153–154, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225603.

Rudolph:1946:ECR

[601] Richard C. Rudolph. Early Chinese references to fossil fish. Isis, 36(3–4): 155, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225604.

Thorndike:1946:JSA

[602] Lynn Thorndike. John of St. Amand on the magnet. Isis, 36(3–4):156– 157, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225605.

Thorndike:1946:DA

[603] Lynn Thorndike. Dalmatius again. Isis, 36(3–4):158, October 1946. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/225606.

Pelseneer:1946:RL

[604] Jean Pelseneer. La religion de Laplace. Isis, 36(3–4):158–160, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225607.

Richeson:1946:FTE

[605] A. W. Richeson. On Faraday’s terminology in electrolysis. Isis, 36(3–4): 160–162, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/225608.

Schlesinger:1946:AHL

[606] Arthur M. Schlesinger. An American historian looks at science and technology. Isis, 36(3–4):162–166, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225609.

Adams:1946:AWS

[607] C. W. Adams. The age at which scientists do their best work. Isis, 36 (3–4):166–169, October 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225610.

Sarton:1946:SNC

[608] George Sarton and Frances Siegel. Sixty-ninth critical bibliography of the history and philosophy of science and of the history of civilization (to December 1945). Isis, 36(3–4):170–248, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225611.

Reymond:1946:NC

[609] Arnold Reymond, G. S., and I. B. Cohen. Notes and correspondence. Isis, 36(3–4):248–251, October 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225612.

Rosenthal:1946:BRB

[610] Franz Rosenthal. Book review: On Medical Experience by Galen; R. Walzer. Isis, 36(3–4):251–255, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225613.

Sarton:1946:BRBc

[611] G. Sarton. Book review: Commentaires de Pappus et de Théon d’Alexandrie sur l’Almageste. Tome III by Adolphe Rome; Theon d’Alexandrie. Isis, 36(3–4):255–256, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225614.

S:1946:BRBa

[612] G. S. Book review: The Herbal of Rufinus by Lynn Thorndike; Francis S. Benjamin, Jr. Isis, 36(3–4):256–257, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225615.

Cohen:1946:BRBb

[613] I. Bernard Cohen. Book review: Diary of a Journey through the Carolinas, Georgia, and Florida, from July 1, 1765, to April 10, 1766 by John Bartram; Travels in Georgia and Florida, 1773–74, a Report to Dr. John Fothergill by William Bartram. Isis, 36(3–4):257–259, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225616.

Sarton:1946:BRBd

[614] George Sarton. Book review: The Way of an Investigator by Walter Bradford Cannon. Isis, 36(3–4):259–260, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225617.

S:1946:BRBb

[615] G. S. Book review: Science and Scientists in the Netherlands Indies by Pieter Honig; Frans Verdoorn. Isis, 36(3–4):260–261, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225618.

Rosen:1946:BRB

[616] Edward Rosen. Book review: Geschichte und Bibliographie der astronomischen Literatur in Deutchland zur Zeit der Renaissance by Ernst Zinner; Entstehung und Ausbreitung der coppernicanischen Lehre. Isis, 36(3–4):261–266, October 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225619.

Bobrovnikoff:1946:BRB

[617] N. T. Bobrovnikoff. Book review: The Comet of 1577 by C. Doris Hellman. Isis, 36(3–4):266–270, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225620.

Zirkle:1946:BRBc

[618] Conway Zirkle. Book review: The Woods Hole Marine Biological Laboratory by Frank R. Lillie. Isis, 36(3–4):270–271, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225621.

Singer:1946:BRB

[619] Dorothea Waley Singer. Book review: The Next Development in Man by L. L. Whyte. Isis, 36(3–4):271–273, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225622.

Roberts:1946:BRB

[620] Leo Roberts. Book review: Philosophical Understanding and Religious Truth by Erich Frank. Isis, 36(3–4):273, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225623.

Sarton:1946:BRBe

[621] George Sarton. Book review: Elenco dei manoscritti arabi islamici della Biblioteca Vaticana by Giorgio Levi Della Vida; Ricerche sulla formazione del piu antico fondo dei manoscritti orientali della Biblioteca Vaticana. Isis, 36(3–4):273–275, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225624.

Pagel:1946:BRB

[622] Walter Pagel. Book review: The Falling Sickness by Owsei Temkin. Isis, 36(3–4):275–278, October 1946. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225625.

Pogo:1946:AD

[623] Alexander Pogo. Administrative documents. Isis, 36(3–4):279–281, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225626.

Anonymous:1946:VI

[624] Anonymous. Volume information. Isis, 36(3–4):282–285, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225601.

Anonymous:1946:BMb

[625] Anonymous. Back matter. Isis, 36(3–4):286, October 1946. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225627.

Anonymous:1947:FMa

[626] Anonymous. Front matter. Isis, 37(1–2):1–4, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226150.

Sarton:1947:PVQ

[627] George Sarton. Preface to volume 37: Qualifications of teachers of the history of science. Isis, 37(1–2):5–7, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226151.

Holmes:1947:KBP

[628] S. J. Holmes. K. E. Von Baer’s perplexities over evolution. Isis, 37(1–2): 7–14, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226152.

Wilson:1947:DCT

[629] J. Walter Wilson. Dutrochet and the Cell theory. Isis, 37(1–2):14–21, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226153.

Lownes:1947:CDC

[630] Albert E. Lownes. Charles Darwin to Charles Harrison Blackley (an early chapter on pollen allergy). Isis, 37(1–2):21–24, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226154.

Montagu:1947:TDS

[631] M. F. Ashley Montagu. Theognis, Darwin, and social selection. Isis, 37(1– 2):24–26, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226155.

Hedgpeth:1947:FMN

[632] Joel W. Hedgpeth. Fishers of the murex (notes for a bibliography of marine natural history). Isis, 37(1–2):26–32, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226156.

Meyerhof:1947:EMM

[633] Max Meyerhof. The earliest mention of a manniparous insect. Isis, 37(1– 2):32–36, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226157.

Neugebauer:1947:SAA

[634] O. Neugebauer. Studies in ancient astronomy. VIII. The water clock in Babylonian astronomy. Isis, 37(1–2):37–43, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226158.

Renaud:1947:LDM

[635] H. P. J. Renaud. L’origine du mot “almanach”. Isis, 37(1–2):44–46, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226159.

Thorndike:1947:BFH

[636] Lynn Thorndike. Blasius the Franciscan and his works on Computus. Isis, 37(1–2):46–47, May 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226160.

Richeson:1947:FAP

[637] A. W. Richeson. The first arithmetic printed in English. Isis, 37(1–2): 47–56, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226161.

Miller:1947:AHP

[638] Wm. Marion Miller. An attempt to harness perpetual motion in Ohio in the Nineteenth Century. Isis, 37(1–2):57, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226162.

Cohen:1947:DAA

[639] I. Bernard Cohen. The development of aeronautics in America: A review of recent publications. Isis, 37(1–2):58–64, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226163.

Mach:1947:LEM

[640] Ernst Mach and Robert H. Lowie. Letters from Ernst Mach to Robert H. Lowie. Isis, 37(1–2):65–68, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226164.

Adams:1947:QA

[641] C. W. Adams, George Sarton, and James R. Ware. Queries and answers. Isis, 37(1–2):68–73, May 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226165.

Thorndike:1947:NC

[642] Lynn Thorndike, G. S., C. W. Adams, Jean Pelseneer, and Semyon P. Rudnykh. Notes and correspondence. Isis, 37(1–2):73–79, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226166.

Anonymous:1947:EB

[643] Anonymous. Editing of Isis. Isis, 37(1–2):79–80, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226167.

Anonymous:1947:P

[644] Anonymous. Personalia. Isis, 37(1–2):80, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226168.

Anonymous:1947:Oa

[645] Anonymous. Obituaries. Isis, 37(1–2):81, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226169.

Johnson:1947:BRB

[646] Francis R. Johnson. Book review: Copernicus and His World by Hermann Kesten; E. B. Ashton; Norbert Guterman. Isis, 37(1–2):82, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226170.

Montagu:1947:BRBa

[647] M. F. Ashley Montagu. Book review: Early Science in Oxford. Vol. XIV, Life and Letters of Edward Lhwyd by Gunther, R. T. Isis, 37(1–2):82–83, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226171.

Roberts:1947:BRBa

[648] Leo Roberts. Book review: Entretiens sur Descartes by Alexandre Koyré. Isis, 37(1–2):83–84, May 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226172.

Zirkle:1947:BRBa

[649] Conway Zirkle. Book review: Thomas Jefferson’s Garden Book by Edwin Morris Betts; =Thomas Jefferson and the Scientific Trends of His Time by Charles A. Browne; Jefferson and Agriculture by Everett E. Edwards; Papers Read before the American Philosophical Society in Celebration of the Bicentennial of Thomas Jefferson by Thomas Jefferson. Isis, 37(1–2): 84–85, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226173.

Cohen:1947:BRBa

[650] I. Bernard Cohen. Book review: A Benjamin Franklin Reader by Nathan G. Goodman; Benjamin Franklin’s Autobiographical Writings by Carl Van Doren. Isis, 37(1–2):85–86, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226174.

Cohen:1947:BRBb

[651] I. Bernard Cohen. Book review: Lavoisier y la formación de la teoría química moderna by Aldo Mieli. Isis, 37(1–2):86–87, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226175.

Sarton:1947:BRBa

[652] George Sarton. Book review: Correspondance inédite entre Réaumur et Abraham Trembley comprenant 113 lettres recueillies et annotées by Maurice Trembley. Isis, 37(1–2):87–89, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226176.

Zirkle:1947:BRBb

[653] Conway Zirkle. Book review: American Botany, 1873–1892; Decades of Transition by Andrew Denny Rodgers. Isis, 37(1–2):89–90, May 1947.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226177.

Saunders:1947:BRBa

[654] J. B. deC. M. Saunders. Book review: François Magendie, Pioneer in Experimental Physiology and Scientific Medicine in XIX Century France by J. M. D. Olmsted. Isis, 37(1–2):90–91, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226178.

Montagu:1947:BRBb

[655] M. F. Ashley Montagu. Book review: Patrick Geddes: Maker of the Future by Philip Boardman. Isis, 37(1–2):91–92, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226179.

Saunders:1947:BRBb

[656] J. B. deC. M. Saunders. Book review: Harvey Cushing. A Biography by John F. Fulton. Isis, 37(1–2):92–93, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226180.

Roberts:1947:BRBb

[657] Leo Roberts. Book review: The Philosophy of Bertrand Russell. Volume V in “The Library of Living Philosophers” by Arthur Schillp. Isis, 37(1– 2):93–95, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226181.

Zirkle:1947:BRBc

[658] Conway Zirkle. Book review: The Wheats of Classical Antiquity by Naum Jasny. Isis, 37(1–2):96, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226182.

Dijksterhuis:1947:BRB

[659] E. J. Dijksterhuis. Book review: Mathematical Cuneiform Texts by O. Neugebauer; A. Sachs. Isis, 37(1–2):96–97, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226183.

Diller:1947:BRBa

[660] Aubrey Diller. Book review: Asclepius. A Collection and Interpretation of the Testimonies by Emma J. Edelstein; Ludwig Edelstein. Isis, 37(1– 2):98, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226184.

Diller:1947:BRBb

[661] Aubrey Diller. Book review: Moira. Fate, Good and Evil, in Greek Thought by William Chase Greene. Isis, 37(1–2):98–99, May 1947. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/226185.

Diller:1947:BRBc

[662] Aubrey Diller. Book review: Paideia: The Ideals of Greek Culture by Werner Jaeger; Gilbert Highet. Isis, 37(1–2):99–100, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226186.

Epstein:1947:BRB

[663] Hans J. Epstein. Book review: The Romance of Daude de Pradas called Dels Auzels Cassadors by Alexander H. Schutz; Moamin et Ghatrif, Traités de fauconnerie et des chiens de chasse by Hakan Tjerneld. Isis, 37(1–2):100–103, May 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226187.

Cohen:1947:BRBc

[664] I. Bernard Cohen. Book review: History of Photography by Josef Maria Eder; Edward Epstean; History of Color Photography by Joseph S. Friedman. Isis, 37(1–2):103–104, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226188.

Lenzen:1947:BRBa

[665] Victor F. Lenzen. Book review: Foundations of Physics by Philipp Frank. Isis, 37(1–2):104–105, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226189.

Farber:1947:BRB

[666] Eduard Farber. Book review: Die Entwicklungsgeschichte der Chemie. Eine Studie by H. E. Fierz-David. Isis, 37(1–2):105–106, May 1947. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/226190.

Zirkle:1947:BRBd

[667] Conway Zirkle. Book review: The New Genetics in the Soviet Union by P. S. Hudson; R. H. Richens. Isis, 37(1–2):106–108, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226191.

Zirkle:1947:BRBe

[668] Conway Zirkle. Book review: Heredity and Its Variability by Trofim Denisovich Lysenko; Theodosius Dobzhansky. Isis, 37(1–2):108, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226192.

Zirkle:1947:BRBf

[669] Conway Zirkle. Book review: Tempo and Mode in Evolution by George Gaylord Simpson. Isis, 37(1–2):109–110, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226193.

Zirkle:1947:BRBg

[670] Conway Zirkle. Book review: Plants and Plant Science in Latin America by Frans Verdoorn. Isis, 37(1–2):110–111, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226194.

Zirkle:1947:BRBh

[671] Conway Zirkle. Book review: Two Billion Acre Farm: An Informal History of American Agriculture by Robert West Howard. Isis, 37(1–2):111– 112, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226195.

Montagu:1947:BRBc

[672] M. F. Ashley Montagu. Book review: A History of Comparative Anatomy. From Aristotle to the Eighteenth Century by F. J. Cole. Isis, 37(1–2):112– 114, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226196.

Saunders:1947:BRBc

[673] J. B. deC. M. Saunders. Book review: One Hundred Years of American Psychiatry by American Psychiatric Association. Isis, 37(1–2):114–118, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226197.

Graubard:1947:BRBa

[674] Mark Graubard. Book review: Configuration of Culture Growth by A. L. Kroeber. Isis, 37(1–2):118–119, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226198.

Montagu:1947:BRBd

[675] M. F. Ashley Montagu. Book review: Justice and World Society by Laurence Stapleton. Isis, 37(1–2):119–120, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226199.

Barber:1947:BRB

[676] Bernard Barber. Book review: The Social Systems of American Ethnic Groups by W. Lloyd Warner; Leo Srole. Isis, 37(1–2):120–121, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226200.

Graubard:1947:BRBb

[677] Mark Graubard. Book review: The East India Company and the British Empire in the Far East by Marguerite Eyer Wilbur. Isis, 37(1–2):121– 122, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226201.

M:1947:BRBa

[678] M. F. A. M. Book review: Art and Scientific Thought by Martin Johnson. Isis, 37(1–2):122, May 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226202.

Saunders:1947:BRBd

[679] J. B. deC. M. Saunders and G. S. Book review: The History of Surgical Anesthesia by Thomas E. Keys. Isis, 37(1–2):122–123, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226203.

Saunders:1947:BRBe

[680] J. B. deC. M. Saunders. Book review: History of Gynecology by Richard A. Leonardo. Isis, 37(1–2):123–124, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226204.

Graubard:1947:BRBc

[681] Mark Graubard. Book review: The Effect of Smallpox on the Destiny of the Amerindian by E. Wagner Stearn; Allen E. Stearn. Isis, 37(1–2): 124–125, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226205.

M:1947:BRBb

[682] M. F. A. M. Book review: Teacher in America by Jacques Barzun. Isis, 37 (1–2):125, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226206.

Anonymous:1947:BMa

[683] Anonymous. Back matter. Isis, 37(1–2):??, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226207.

Anonymous:1947:VIa

[684] Anonymous. Volume information. Isis, 37(1–2):??, May 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226149.

Anonymous:1947:FMb

[685] Anonymous. Front matter. Isis, 37(3–4):127, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225559.

Bridgman:1947:SFR

[686] Percy W. Bridgman. Science and freedom reflections of a physicist. Isis, 37(3–4):128–131, July 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225560.

Allen:1947:RSL

[687] Phyllis Allen. The Royal Society and Latin America as reflected in the Philosophical Transactions 1665–1730. Isis, 37(3–4):132–138, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225561.

Hendrickson:1947:WAN

[688] Walter B. Hendrickson. The Western Academy of Natural Sciences of Cincinnati. Isis, 37(3–4):138–145, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225562.

Mullett:1947:TMP

[689] Charles F. Mullett. Thomas Morrison and the pursuit of knowledge. Isis, 37(3–4):146–147, July 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225563.

Boyer:1947:NEG

[690] Carl B. Boyer. Note on an early graph of statistical data (Huygens 1669). Isis, 37(3–4):148–149, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225564.

Worrell:1947:NMC

[691] W. H. Worrell. Note on modern coptic inks. Isis, 37(3–4):149–150, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225565.

Thorndike:1947:WWQ

[692] Lynn Thorndike. Who wrote quadrans vetus? Isis, 37(3–4):150–153, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225566.

Levy:1947:NLT

[693] Raphael Levy. A note on the Latin translators of ibn ezra. Isis, 37(3–4): 153–155, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225567.

Zirkle:1947:TCS

[694] Conway Zirkle. The theory of concentric spheres: Edmund Halley, Cotton Mather, & John Cleves Symmes. Isis, 37(3–4):155–159, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225568.

Ling:1947:IUG

[695] Wang Ling. On the invention and use of gunpowder and firearms in China. Isis, 37(3–4):160–178, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225569.

Sarton:1947:NC

[696] George Sarton, C. Doris Hellman, and Sidney M. Edelstein. Notes & correspondence. Isis, 37(3–4):178–182, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225570.

Anonymous:1947:Ob

[697] Anonymous. Obituaries. Isis, 37(3–4):183, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225571.

Sarton:1947:SCB

[698] George Sarton and Frances Siegel. Seventieth critical bibliography of the history and philosophy of science and of the history of civilization (to January 1947). Isis, 37(3–4):184–282, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225572.

Pogo:1947:AD

[699] Alexander Pogo. Administrative documents. Isis, 37(3–4):282–284, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225573.

Anonymous:1947:VIb

[700] Anonymous. Volume information. Isis, 37(3–4):285–288, July 1947. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/225558.

Anonymous:1947:BMb

[701] Anonymous. Back matter. Isis, 37(3–4):??, July 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225574.

Anonymous:1947:FMc

[702] Anonymous. Front matter. Isis, 38(1–2):1–2, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225441.

Sarton:1947:PVT

[703] George Sarton. Preface to volume thirty-eight: A tribute to Gilbert Murray and a plea for Greek studies. Isis, 38(1–2):3–11, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225442.

Cowan:1947:HPS

[704] Thomas A. Cowan. The historian and the philosophy of science. Isis, 38 (1–2):11–18, November 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225443.

Kattsoff:1947:PSM

[705] Louis O. Kattsoff. Ptolemy and scientific method: A note on the history of an idea. Isis, 38(1–2):18–22, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225444.

Bayon:1947:CSR

[706] H. P. Bayon. Calvin, Serveto and Rabelais. Isis, 38(1–2):22–28, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225445.

Zirkle:1947:PP

[707] Conway Zirkle. Population problems. Isis, 38(1–2):28–33, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225446.

Sarton:1947:PJM

[708] George Sarton. Paul, Jules, and Marie Tannery (with a note on Grégoire Wyrouboff). Isis, 38(1–2):33–51, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225447.

Lejeune:1947:ALR

[709] Albert Lejeune. Archimède et la loi de la Réflexion. Isis, 38(1–2):51– 53, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225448.

Boyer:1947:NEE

[710] Carl B. Boyer. Note on epicycles & the ellipse from Copernicus to Lahire. Isis, 38(1–2):54–56, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225449.

Kennedy:1947:AKB

[711] E. S. Kennedy. Al-Kāshī’s “Plate of Conjunctions”. Isis, 38(1–2):56–59, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225450.

Dodson:1947:HSP

[712] Sarah Dodson. Holinshed’s sources for the prognostications about the years 1583 and 1588. Isis, 38(1–2):60–62, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/225451.

Dubs:1947:BA

[713] Homer H. Dubs. The beginnings of alchemy. Isis, 38(1–2):62–86, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225452.

Coomaraswamy:1947:GEI

[714] Ananda K. Coomaraswamy. Gradation and evolution–ii. Isis, 38(1–2):87– 94, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225453.

Sarton:1947:QA

[715] George Sarton and J. Pelseneer. Queries and answers. Isis, 38(1–2):94– 98, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/225454.

Richmond:1947:THS

[716] Donald Richmond and R. J. Forbes. The teaching of the history of science. Isis, 38(1–2):98–99, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225455.

S:1947:NC

[717] G. S., Chapman Grant, John F. Fulton, J. E. Richey, and David Haldane. Notes and correspondence. Isis, 38(1–2):100–103, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225456.

Hofsten:1947:ONE

[718] Nils V. Hofsten. Obituary notice: Erik Nordenskiöld (1872–1933). Isis, 38 (1–2):103–106, November 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225457.

Singer:1947:ONR

[719] Dorothea Waley Singer. Obituary notice: Robert Steele (1860–1944). Isis, 38(1–2):107, November 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225458.

Anonymous:1947:Oc

[720] Anonymous. Obituaries. Isis, 38(1–2):107–109, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225459.

Saunders:1947:BRBf

[721] J. B. deC. M. Saunders. Book review: A Prelude to Modern Science; Being a Discussion of the History, Sources and Circumstances of the “Tabulae Anatomicae Sex” of Vesalius by Charles Singer; C. Rabin. Isis, 38(1– 2):109–111, November 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225460.

S:1947:BRB

[722] G. S. Book review: The Suma Oriental of Tomé Pires by Armando Cortesão. Isis, 38(1–2):111, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225461.

Urdang:1947:BRB

[723] George Urdang. Book review: Die Historischen Grundlagen der Intravenösen Injektion. Ein Beitrag zur Medizingeschichte des 17. Jahrhunderts by Heinrich Buess. Isis, 38(1–2):111–114, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225462.

Studley:1947:BRB

[724] Duane Studley. Book review: Newton at the Mint by John Craig. Isis, 38 (1–2):114–115, November 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225463.

Cohen:1947:BRBd

[725] I. Bernard Cohen. Book review: Newton Demands the Muse. Newton’s “Opticks” and the Eighteenth Century Poets by Marjorie Hope Nicolson. Isis, 38(1–2):115–116, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225464.

Cohen:1947:BRBe

[726] I. Bernard Cohen. Book review: Benjamin Silliman, 1779–1864, Pathfinder in American Science by John F. Fulton; Elizabeth H. Thomson; The Early Work of Willard Gibbs in Applied Mechanics, Comprising the Text of His Hitherto Unpublished Ph.D. Thesis and Accounts of His Mechanical Inventions by Willard Gibbs; Lynde Phelps Wheeler; Everett Oyler Waters; Samuel William Dudley; Yale Science. The First Hundred Years, 1701–1801 by Louis W. McKeehan. Isis, 38(1–2):117–119, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225465.

Winter:1947:BRB

[727] H. J. J. Winter. Book review: Las Leyes de Ohm. Ensayo de Historia Científico y Humana by Cortes Pla. Isis, 38(1–2):119–120, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225466.

Sarton:1947:BRBb

[728] George Sarton. Book review: Codicum Romanorum partem quartam descripsit by Stephanus Weinstock. Isis, 38(1–2):120, November 1947. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/225467.

Sarton:1947:BRBc

[729] George Sarton. Book review: Raccolta di Scritti Editi e Inedili by Carlo Alfonso Nallino. Isis, 38(1–2):120–121, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225468.

Frye:1947:BRB

[730] Richard N. Frye. Book review: Osmanli Türklerinde Ilim. (Science among the Ottoman Turks) by A. Adnan Adivar. Isis, 38(1–2):121–125, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225469.

Fulton:1947:BRB

[731] John F. Fulton. Book review: On Understanding Science. An Historical Approach by James B. Conant. Isis, 38(1–2):125–127, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225470.

Sarton:1947:BRBd

[732] George Sarton. Book review: Studies and Essays in the History of Science and Learning Offered in Homage to George Sarton on the Occasion of His Sixtieth Birthday, 31 August 1944 by M. F. Ashley Montagu. Isis, 38 (1–2):127–128, November 1947. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/225471.

Lenzen:1947:BRBb

[733] V. F. Lenzen. Book review: The Ground of Induction by Donald Williams. Isis, 38(1–2):128–129, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

225472.

Lenzen:1947:BRBc

[734] V. F. Lenzen. Book review: Science and the Meaning of Truth by Martin Christopher Johnson. Isis, 38(1–2):129–130, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225473.

Sarton:1947:BRBe

[735] George Sarton. Book review: Richard Hakluyt and His Successors by Edward Lynam. Isis, 38(1–2):130, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225474.

Graubard:1947:BRBd

[736] Mark Graubard. Book review: Adventures of the Mind by Arturo Castiglioni. Isis, 38(1–2):131, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225475.

Leake:1947:BRB

[737] Chauncey D. Leake. Book review: The Development of Inhalation Anesthesia with Special Reference to the Years 1846–1900 by Barbara M. Duncum. Isis, 38(1–2):131–133, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/225476.

Kofoid:1947:BRB

[738] Charles A. Kofoid. Book review: Scientific Societies in the United States by Ralph S. Bates. Isis, 38(1–2):133, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/225477.

Anonymous:1947:BMc

[739] Anonymous. Back matter. Isis, 38(1–2):134, November 1947. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/225478.

Anonymous:1947:VIc

[740] Anonymous. Volume information. Isis, 38(1–2):??, November 1947. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/225440.

Anonymous:1948:FMa

[741] Anonymous. Front matter. Isis, 38(3–4):135–136, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226105.

Sarton:1948:SPV

[742] George Sarton. Second preface to volume 38: The study of early scientific textbooks. Isis, 38(3–4):137–148, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226106.

Cohen:1948:C

[743] I. Bernard Cohen. Comments. Isis, 38(3–4):149–150, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226107.

Patterson:1948:RHC

[744] Louise Diehl Patterson. Robert Hooke and the conservation of energy. Isis, 38(3–4):151–156, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226108.

Praus:1948:MPI

[745] Alexis A. Praus. Mechanical principles involved in primitive tools and those of the machine age. Isis, 38(3–4):157–160, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226109.

Denny:1948:LHD

[746] Margaret Denny. Linnaeus and his disciple in Carolina: Alexander Garden. Isis, 38(3–4):161–174, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226110.

Emery:1948:JWU

[747] Clark Emery. John Wilkins’ universal language. Isis, 38(3–4):174–185, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226111.

Tasch:1948:QMG

[748] Paul Tasch. Quantitative measurements and the Greek atomists. Isis, 38 (3–4):185–189, February 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226112.

Caley:1948:EKU

[749] Earle R. Caley. The earliest known use of a material containing uranium. Isis, 38(3–4):190–193, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226113.

Cramer:1948:SRE

[750] Frederick Cramer. Some recent European publications on ancient pseudoscience and its adversaries. Isis, 38(3–4):194–197, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226114.

Downey:1948:PAA

[751] Glanville Downey. Pappus of Alexandria on architectural studies. Isis, 38 (3–4):197–200, February 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226115.

Thorndike:1948:AOP

[752] Lynn Thorndike. Astronomical observations at Paris from 1312 to 1315. Isis, 38(3–4):200–205, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226116.

Lattin:1948:ECM

[753] Harriet Pratt Lattin. The Eleventh Century MS Munich 14436: Its contribution to the history of co-ordinates, of logic, of German studies in France. Isis, 38(3–4):205–225, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226117.

Sagui:1948:MBF

[754] C. L. Sagui. La meunerie de Barbegal (France) et les roues hydrauliques chez les anciens et au moyen âge. (French) [The milling of Barbegal (France) and water wheels in Ancient and Middle Ages]. Isis, 38(3–4): 225–231, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226118.

Sokol:1948:NQI

[755] A. E. Sokol. The name of quelpaert Island. Isis, 38(3–4):231–235, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226119.

Spooner:1948:DNT

[756] Roy C. Spooner and C. H. Wang. The Divine Nine Turn Tan Sha Method, a Chinese alchemical recipe. Isis, 38(3–4):235–242, February 1948. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/226120.

Sergescu:1948:NC

[757] Pierre Sergescu, J. A. Poliakov, George Sarton, and E. N. da C. Andrade. Notes and correspondence. Isis, 38(3–4):242–244, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226121.

Walter:1948:HS

[758] Emil J. Walter and Jean Pelseneer. The history of science. Isis, 38(3– 4):244–245, February 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226122.

Anonymous:1948:Oa

[759] Anonymous. Obituaries. Isis, 38(3–4):246–247, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226123.

Singer:1948:BRB

[760] Dorothea Waley Singer. Book review: Il Sommario del Processo di Giordano Bruno con Appendice di documenti sull’Eresia e l’Inquisizione a Modena nel Secolo XVI by Angelo Mercati. Isis, 38(3–4):247–248, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226124.

Pagel:1948:BRB

[761] Walter Pagel. Book review: J. B. van Helmont by Henry De Waele. Isis, 38 (3–4):248–249, February 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226125.

Sarton:1948:BRBa

[762] George Sarton. Book review: Journal tenu par Isaac Beeckman de 1604 à 1634, publié avec une introduction et des notes by Cornelis De Waard. Isis, 38(3–4):249–250, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226126.

C:1948:BRB

[763] I. B. C. Book review: Die Historischen Grundlagen der Intravenösen Injektion. Ein Beitrag zur Medizingeschichte des 17. Jahrhunderts by Heinrich Buess. Isis, 38(3–4):250–251, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226127.

Richards:1948:BRB

[764] I. A. Richards. Book review: A Treatise on Language by Alexander Bryan Johnson; David Rynin. Isis, 38(3–4):251–252, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226128.

Cohen:1948:BRBa

[765] I. Bernard Cohen. Book review: Einstein, His Life and Times by Philipp Frank; George Rosen; Shuichi Kusaka. Isis, 38(3–4):252–253, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226129.

Sarton:1948:BRBb

[766] George Sarton. Book review: A Glossary of Greek Fishes by D’Arcy Wentworth Thompson. Isis, 38(3–4):254, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226130.

Tasch:1948:BRB

[767] Paul Tasch. Book review: The Problems of Change in Greek Science by Clifford Pierson Osborne. Isis, 38(3–4):254–255, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226131.

Cohen:1948:BRBb

[768] I. Bernard Cohen. Book review: Science Advances by J. B. S. Haldane. Isis, 38(3–4):255–256, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226132.

Boyer:1948:BRB

[769] Carl B. Boyer. Book review: Mathematics as a Culture Clue, and Other Essays by Cassius Jackson Keyser. Isis, 38(3–4):256–257, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226133.

Rosenthal-Schneider:1948:BRB

[770] Ilse Rosenthal-Schneider and I. B. C. Book review: Geschichte der Physik by Max v. Laue. Isis, 38(3–4):258–260, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226134.

Davis:1948:BRB

[771] T. L. Davis. Book review: Three Centuries of Irish Chemists by Deasmumhan Ó. Raghallaigh. Isis, 38(3–4):260–261, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226135.

Graubard:1948:BRBa

[772] Mark Graubard. Book review: Two Blades of Grass — a History of Scientific Developments in the U. S. Department of Agriculture by T. Swann Harding. Isis, 38(3–4):261–262, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226136.

Sarton:1948:BRBc

[773] George Sarton. Book review: Catalogue des cartes nautiques manuscrites sur parchemin, 1300–1700. Cartes hollandaises. La Cartographie de la Compagnie des Indes Orientales, 1593–1743; La Mappemonde de Petrus Plancius gravée par Josua van den Ende, 1604, d’après l’unique exemplaire de la Bibliothèque Nationale de Paris by Marcel Destombes. Isis, 38(3–4):262, February 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226137.

Sarton:1948:BRBd

[774] George Sarton. Book review: Critique et géologie by Emmanuel De Margerie. Isis, 38(3–4):263–264, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226138.

Cole:1948:BRB

[775] F. J. Cole, Herbert Friedmann, and M. F. Ashley Montagu. Book review: History of Comparative Anatomy. From Aristotle to the Eighteenth Century by Francis J. Cole. Isis, 38(3–4):264–266, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226139.

Ulich:1948:BRB

[776] Robert Ulich. Book review: Psychologische Grundlagen der geschichtlichen und sozialen Entwicklung by Emil J. Walter. Isis, 38(3–4):266–267, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226140.

Graubard:1948:BRBb

[777] Mark Graubard. Book review: Scientific Man vs. Power Politics by Hans J. Morgenthau. Isis, 38(3–4):267–268, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226141.

Montagu:1948:BRBa

[778] M. F. Ashley Montagu. Book review: Art and Freedom by Horace M. Kallen. Isis, 38(3–4):268, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226142.

Roberts:1948:BRBa

[779] Leo Roberts. Book review: A History of Western Philosophy by Bertrand Russell. Isis, 38(3–4):268–270, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226143.

Saunders:1948:BRBa

[780] J. B. deC. M. Saunders. Book review: The Medical History of Early Texas, 1528–1853 by Pat Ireland Nixon. Isis, 38(3–4):270, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226144.

Leikind:1948:BRB

[781] Morris C. Leikind. Book review: Miracles from Microbes; The Road to Streptomycin by Samuel Epstein; Beryl Williams. Isis, 38(3–4):270–271, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226145.

Montagu:1948:BRBb

[782] M. F. Ashley Montagu. Book review: Not by Bread Alone by Vilhjalmur Stefansson. Isis, 38(3–4):271–272, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226146.

Ballou:1948:BRB

[783] Richard Boyd Ballou. Book review: Three Thousand Years of Educational Wisdom: Selections from Great Documents by Robert Ulich. Isis, 38(3– 4):272–273, February 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226147.

Anonymous:1948:BMa

[784] Anonymous. Back matter. Isis, 38(3–4):274, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226148.

Anonymous:1948:VIa

[785] Anonymous. Volume information. Isis, 38(3–4):275–278, February 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226104.

Anonymous:1948:FMb

[786] Anonymous. Front matter. Isis, 39(1–2):1–2, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226763.

Sarton:1948:PVT

[787] George Sarton. Preface to volume 39: The Tower of Babel. Isis, 39(1–2): 3–15, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226764.

Brown:1948:JDT

[788] Harcourt Brown. Jean Denis and transfusion of blood, Paris, 1667–1668. Isis, 39(1–2):15–29, May 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226765.

Clagett:1948:SGA

[789] Marshall Clagett. Some general aspects of physics in the middle ages. Isis, 39(1–2):29–44, May 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226766.

Pagel:1948:JVA

[790] Walter Pagel. Jung’s views on alchemy. Isis, 39(1–2):44–48, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226767.

Hellman:1948:GR

[791] C. Doris Hellman. Georg Rethe. Isis, 39(1–2):49–55, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226768.

Cohen:1948:RF

[792] I. Bernard Cohen. Roemer and Fahrenheit. Isis, 39(1–2):56–58, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226769.

Thorndike:1948:NC

[793] Lynn Thorndike, George Sarton, Joseph de Somogyi, L. Carrington Goodrich, and Eduard Farber. Notes and correspondence. Isis, 39(1– 2):58–66, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226770.

S:1948:THS

[794] G. S. and Harcourt Brown. Teaching of the history of science. Isis, 39 (1–2):67, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226771.

Anonymous:1948:Pa

[795] Anonymous. Personalia. Isis, 39(1–2):68, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226772.

Anonymous:1948:WP

[796] Anonymous. Work in progress. Isis, 39(1–2):68, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226773.

Anonymous:1948:Ob

[797] Anonymous. Obituaries. Isis, 39(1–2):68–69, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226774.

Sarton:1948:SFC

[798] George Sarton and Frances Siegel. Seventy-first critical bibliography of the history and philosophy of science and of the history of civilization (to October 1947). Isis, 39(1–2):70–139, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226775.

Anonymous:1948:BMb

[799] Anonymous. Back matter. Isis, 39(1–2):140–141, May 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226776.

Anonymous:1948:FMc

[800] Anonymous. Front matter. Isis, 39(3):143–144, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226308.

Holmes:1948:MBT

[801] S. J. Holmes. Micromerism in biological theory. Isis, 39(3):145–158, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226309.

Streeter:1948:JWJ

[802] John W. Streeter. John Winthrop, Junior, and the fifth satellite of Jupiter. Isis, 39(3):159–163, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226310.

Buhler:1948:SSI

[803] Curt F. Bühler. The statistics of scientific incunabula. Isis, 39(3):163– 168, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226311.

Lattin:1948:USM

[804] Harriet Pratt Lattin. Use of a sphere by Macrobiusxo. Isis, 39(3):168– 169, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226312.

Andrews:1948:OSC

[805] Alfred C. Andrews. Orach as the spinach of the classical period. Isis, 39(3):169–172, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226313.

Hellman:1948:ATC

[806] C. Doris Hellman. Additional tracts on the comet of 1577. Isis, 39(3):172– 174, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226314.

Anonymous:1948:Oc

[807] Anonymous. Obituaries. Isis, 39(3):174, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226315.

Anonymous:1948:Pb

[808] Anonymous. Personalia. Isis, 39(3):174–175, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226316.

Gandz:1948:BRBa

[809] Solomon Gandz. Book review: Rabbi Meir of Rothenburg by Irving A. Agus. Isis, 39(3):175, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226317.

Johnson:1948:BRB

[810] Francis R. Johnson. Book review: Sun, Stand Thou Still: The Life and Work of Copernicus the Astronomer by Angus Armitage. Isis, 39(3):175– 176, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226318.

Hellman:1948:BRB

[811] C. Doris Hellman. Book review: The Life and Times of Tycho Brahe by John Allyne Gade. Isis, 39(3):176–179, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226319.

Sarton:1948:BRBe

[812] George Sarton. Book review: Correspondance du P. Marin Mersenne, Religieux Minime by Marin Mersenne; Paul Tannery; Cornelis De Waard. Isis, 39(3):179–181, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226320.

Stimson:1948:BRB

[813] Dorothy Stimson. Book review: Hartlib, Dury and Comenius: Gleanings from Hartlib’s Papers by G. H. Turnbull. Isis, 39(3):181–182, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226321.

Sarton:1948:BRBf

[814] George Sarton. Book review: Hendrik Van Deventer, Medicinae Doctor, 1651–1724. Leven en Werken by A. J. M. Lamers. Isis, 39(3):182–184, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226322.

Sarton:1948:BRBg

[815] George Sarton. Book review: Verborgen Paradijzen. Het Leven en de Werken van Maria Sibylla Merian (1647–1717) by J. Stuldreher-Nienhuis. Isis, 39(3):184, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226323.

Sarton:1948:BRBh

[816] George Sarton. Book review: Quest for Mysteries by Heinrich Schneider; La franc-maçonnerie et le révolution intellectuelle du XVIIIe siècle by Bernard Faÿ. Isis, 39(3):184–187, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226324.

Fulton:1948:BRB

[817] John F. Fulton. Book review: Charles-Édouard Brown–Séquard. A Nineteenth Century Neurologist and Endocrinologist by J. M. D. Olmsted. Isis, 39(3):187–188, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226325.

Sarton:1948:BRBi

[818] George Sarton. Book review: Etudes d’esthétique médiévale by Edgar De Bruyne. Isis, 39(3):188–190, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226326.

Gandz:1948:BRBb

[819] Solomon Gandz. Book review: Ugaritic Mythology by Julian Obermann. Isis, 39(3):190–191, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226327.

Sarton:1948:BRBj

[820] George Sarton. Book review: The Scientists Speak by Warren Weaver. Isis, 39(3):191–192, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226328.

Bridgman:1948:BRB

[821] P. W. Bridgman. Book review: The Logic of the Sciences and the Humanities by F. S. C. Northrop. Isis, 39(3):192–194, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226329.

VanBiesbroeck:1948:BRB

[822] G. Van Biesbroeck. Book review: Traité de l’Astrolabe by Henri Michel. Isis, 39(3):194, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226330.

Lenzen:1948:BRB

[823] V. F. Lenzen. Book review: L’évolution de la notion de phénomène physique, des primitifs à Bohr et Louis de Broglie by Jean Pelseneer. Isis, 39(3):194–196, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226331.

Zirkle:1948:BRB

[824] Conway Zirkle. Book review: English Naturalists from Neckam to Ray by Charles E. Raven. Isis, 39(3):196–197, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/226332.

Cohen:1948:BRBc

[825] I. Bernard Cohen. Book review: History as a Literary Art: An Appeal to Young Historians by Samuel Eliot Morison. Isis, 39(3):197–198, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226333.

Gandz:1948:BRBc

[826] Solomon Gandz. Book review: The Poetry of History by Emery Neff. Isis, 39(3):198–199, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226334.

Brinton:1948:BRB

[827] Crane Brinton. Book review: The Dignity of Man. Studies in the Persistence of an Idea by Herschel Baker. Isis, 39(3):199–200, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226335.

Roberts:1948:BRBb

[828] Leo Roberts. Book review: The Foundation of Phenomenology by Martin Farber. Isis, 39(3):200–201, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226336.

Saunders:1948:BRBb

[829] John B. de C. M. Saunders. Book review: American Medical Research Past and Present by Richard H. Shryock. Isis, 39(3):201–202, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226337.

Sarton:1948:BRBk

[830] George Sarton. Book review: Medicine and Health in the Soviet Union by Henry E. Sigerist; Julia Older. Isis, 39(3):202–203, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226338.

Pogo:1948:AD

[831] Alexander Pogo. Administrative documents. Isis, 39(3):204–206, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226339.

Sarton:1948:GT

[832] George Sarton. Girolamo tartarotti, 1749. Isis, 39(4):207–212, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227709.

Thorndike:1948:TWM

[833] Lynn Thorndike. Thomas Werkwoth on the motion of the eighth sphere. Isis, 39(4):212–215, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227710.

McCracken:1948:AKU

[834] George E. McCracken. Athanasius Kircher’s universal polygraphy. Isis, 39(4):215–228, November 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227711.

Jung:1948:ASI

[835] Li Jung and Lien che Tu Fang. An account of the salt industry at Tzuliu-ching Tzu-liu-ching chi. Isis, 39(4):228–234, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227712.

Alston:1948:QA

[836] A. H. G. Alston, Laurence M. Klauber, A. P., and George Sarton. Queries and answers. Isis, 39(4):234–237, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227713.

Drabkin:1948:NC

[837] I. E. Drabkin, L. Carrington Goodrich, Lynn Thorndike, and G. S. Notes and correspondence. Isis, 39(4):237–240, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227714.

Anonymous:1948:Od

[838] Anonymous. Obituaries. Isis, 39(4):241, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227715.

Sarton:1948:SSC

[839] George Sarton and Frances Siegel. Seventy-second critical bibliography of the history and philosophy of science and of the history of civilization (to January 1948). Isis, 39(4):242–283, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227716.

Anonymous:1948:VIb

[840] Anonymous. Volume information. Isis, 39(4):284–286, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227707.

Anonymous:1948:FMd

[841] Anonymous. Front matter. Isis, 39(4):??, November 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227708.

Anonymous:1949:FMa

[842] Anonymous. Front matter. Isis, 40(1):1–2, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227412.

Sarton:1949:PVA

[843] George Sarton. Preface to volume 40: An appeal for the republication in book form of Father Bosmans’ studies on Belgian mathematics in the Sixteenth and Seventeenth Centuries. Isis, 40(1):3–6, February 1949. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/227413.

Diller:1949:AME

[844] Aubrey Diller. The ancient measurements of the Earth. Isis, 40(1):6– 9, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227414.

Tasch:1949:DAD

[845] Paul Tasch. Diogenes of Apollonia and Democritus. Isis, 40(1):10–12, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227415.

Shapley:1949:PHO

[846] Dora Shapley. Pre-Huygenian observations of Saturn’s ring. Isis, 40(1): 12–17, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227416.

Birkenmajer:1949:PLC

[847] Alexandre Birkenmajer. Pierre de Limoges commentateur de Richard de Fournival. (French) [Peter of Limoges commentator of Richard of Fournival]. Isis, 40(1):18–31, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227417.

Vallicrosa:1949:SAO

[848] J. M. Millás Vallicrosa. Sobre la autenticidad de una obra astronómica de R. Abraham ibn ’Ezra. (Spanish) [On the authenticity of a work of the astronomer R. Abraham ibn ’Ezra]. Isis, 40(1):32–33, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227418.

Thorndike:1949:MA

[849] Lynn Thorndike. More Abrahamismus. Isis, 40(1):34–35, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227419.

Rudolph:1949:JC

[850] Richard C. Rudolph. The jumar in China. Isis, 40(1):35–37, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227420.

Boas:1949:HPS

[851] Marie Boas. Hero’s pneumatica: A study of its transmission and influence. Isis, 40(1):38–48, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227421.

McKeon:1949:BRB

[852] Richard McKeon. Book review: Introduction to the History of Science. Volume III, Science and Learning in the Fourteenth Century by George Sarton. Isis, 40(1):49–56, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227422.

Saunders:1949:BRBa

[853] J. B. deC. M. Saunders. Book review: Nicolaus Pol Doctor 1494 by Max H. Fisch; Dorothy M. Schullian. Isis, 40(1):56–58, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227423.

Pagel:1949:BRB

[854] W. Pagel. Book review: Jheronimus Bosch by Jacques Combe; Ethel Duncan. Isis, 40(1):58–59, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227424.

Brinton:1949:BRB

[855] Crane Brinton. Book review: The Faith of Reason. The Idea of Progress in the French Enlightenment by Charles Frankel. Isis, 40(1):59–60, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227425.

Cohen:1949:BRBa

[856] I. Bernard Cohen. Book review: Images or Shadows of Divine Things by Jonathan Edwards; Perry Miller. Isis, 40(1):60–62, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227426.

Marx:1949:BRB

[857] Leo Marx. Book review: Yankee Science in the Making by Dirk J. Struik. Isis, 40(1):62–64, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227427.

King:1949:BRB

[858] Ronold W. P. King. Book review: Der Physiker L. V. Lorenz. Eine kritische Untersuchung by Mogens Pihl. Isis, 40(1):64–66, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227428.

Zirkle:1949:BRBa

[859] Conway Zirkle. Book review: Douglas of the Fir by Athelstan George Harvey. Isis, 40(1):66, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227429.

Lenzen:1949:BRBa

[860] V. F. Lenzen. Book review: Erinnerungen I by Max Planck. Isis, 40(1): 66–67, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227430.

Rosenthal-Schneider:1949:BRB

[861] Ilse Rosenthal-Schneider. Book review: Wissenschaftliche Selbstbiographie by Max Planck. Isis, 40(1):67–69, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227431.

Saunders:1949:BRBb

[862] J. B. deC. M. Saunders. Book review: Henry Sewall, Physiologist and Physician by Gerald B. Webb; Desmond Powell. Isis, 40(1):69, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227432.

Neugebauer:1949:BRB

[863] O. Neugebauer. Book review: Die Grosse Weltordnung by Anton Von Mörl. Isis, 40(1):69–70, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227433.

LeCorbeiller:1949:BRBa

[864] Philippe E. Le Corbeiller. Book review: L’apogée de la science technique grecque. Les sciences de la nature et de l’homme, les mathématiques d’Hippocrate à Platon by Abel Rey. Isis, 40(1):70–71, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227434.

Sarton:1949:BRBa

[865] George Sarton. Book review: Les portulans grecs by Armand Delatte. Isis, 40(1):71–72, February 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227435.

Sarton:1949:BRBb

[866] George Sarton. Book review: The Ancient Na-khi Kingdom of Southwest China by Joseph F. Rock. Isis, 40(1):72–73, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227436.

Gantz:1949:BRB

[867] Solomon Gantz. Book review: The Book of Beliefs and Opinions by Saadia Gaon; Samuel Rosenblatt. Isis, 40(1):73, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227437.

Zirkle:1949:BRBb

[868] Conway Zirkle. Book review: Science, Servant of Man: A Layman’s Primer for the Age of Science by I. Bernard Cohen. Isis, 40(1):73–75, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227438.

Ridenour:1949:BRB

[869] Louis N. Ridenour. Book review: The Control of Atomic Energy. A Study of Its Social, Economic, and Political Implications by James R. Newman; Byron S. Miller. Isis, 40(1):75–76, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227439.

Graubard:1949:BRBa

[870] Mark Graubard. Book review: Our Emergent Civilization by Ruth Nanda Anshen. Isis, 40(1):76, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227440.

Zirkle:1949:BRBc

[871] Conway Zirkle. Book review: Our Plundered Planet by Fairfield Osborn. Isis, 40(1):76–77, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227441.

LeCorbeiller:1949:BRBb

[872] P. Le Corbeiller. Book review: Philosophical Aspects of Modern Science by C. E. M. Joad. Isis, 40(1):77–78, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227442.

Lenzen:1949:BRBb

[873] V. F. Lenzen. Book review: Theory of Experimental Inference by C. West Churchman. Isis, 40(1):78, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227443.

Pelseneer:1949:BRB

[874] J. Pelseneer. Book review: L’humanisme scientifique de demain. Les grands courants de la pensée mathématique by F. Le Lionnais. Isis, 40(1): 78–79, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227444.

Bobrovnikoff:1949:BRB

[875] N. T. Bobrovnikoff. Book review: The Royal Art of Astrology by Robert Eisler; Encyclopedia of Astrology by Nicholas De Vore. Isis, 40(1):79– 81, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227445.

Lenzen:1949:BRBc

[876] V. F. Lenzen. Book review: The Growth of Physical Science by James Jeans. Isis, 40(1):81–82, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227446.

Zirkle:1949:BRBd

[877] Conway Zirkle. Book review: Botanik der Gegenwart und Vorzeit in Culturhistorischer Entwickelung by Karl F. W. Jessen. Isis, 40(1):82, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227447.

Leet:1949:BRB

[878] L. Don Leet. Book review: Seventy-Five Years of Progress in the Mineral Industry, 1871–1946, including Proceedings of the 75th Anniversary Celebration and World Conference on Mineral Resources, March, 1947 by A. B. Parsons. Isis, 40(1):82–83, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227448.

Nagel:1949:BRB

[879] Ernest Nagel. Book review: Work and History. An Essay on the Structure of Civilization by Paul Schrecker. Isis, 40(1):83–85, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227449.

Jackman:1949:BRBa

[880] Sidney Jackman. Book review: From Many One by Crane Brinton. Isis, 40(1):85, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227450.

Levin:1949:BRB

[881] Harry Levin. Book review: Essays in the History of Ideas by Arthur O. Lovejoy. Isis, 40(1):85–87, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227451.

Pfeiffer:1949:BRB

[882] Robert H. Pfeiffer. Book review: The Alphabet: A Key to the History of Mankind by David Diringer. Isis, 40(1):87–88, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227452.

Farber:1949:BRB

[883] Eduard Farber. Book review: Abendland und Kultur. Zur Kulturphilosophie der Gegenwart by Kurt Gihring. Isis, 40(1):88, February 1949. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/227453.

Saunders:1949:BRBc

[884] J. B. deC. M. Saunders. Book review: History of Medicine; A Correlative Text Arranged according to Subjects by Cecilia C. Mettler; Fred A. Mettler. Isis, 40(1):88–90, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227454.

Sarton:1949:BRBc

[885] George Sarton. Book review: Vers la médecine sociale by René Sand. Isis, 40(1):90–91, February 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227455.

Leake:1949:BRBa

[886] C. D. Leake. Book review: Psychiatry in a Troubled World: Yesterday’s War and Today’s Challenge by William C. Menninger. Isis, 40(1):91– 92, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227456.

Leake:1949:BRBb

[887] Chauncey D. Leake. Book review: No Retreat from Reason and Other Essays by Alfred E. Cohn. Isis, 40(1):92, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227457.

Anonymous:1949:BMa

[888] Anonymous. Back matter. Isis, 40(1):??, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227458.

Anonymous:1949:FMb

[889] Anonymous. Front matter. Isis, 40(2):93–94, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227033.

Sarton:1949:SPV

[890] George Sarton. Second preface to volume forty: In defense of Petrarca’s book on the remedies for good and evil fortune. Isis, 40(2):95–99, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227034.

Fulton:1949:BS

[891] John F. Fulton. A Prolegomenon to the sciences. Isis, 40(2):99–106, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227035.

Thorndike:1949:VAV

[892] Lynn Thorndike. Visierkunst, Ars Visorandi, or stereometry. Isis, 40(2): 106–107, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227036.

Eisler:1949:MAH

[893] Robert Eisler. Metallurgical anthropology in Hesiod and Plato and the date of a “phoenician lie”. Isis, 40(2):108–112, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227037.

Loria:1949:SBL

[894] Di Gino Loria. Saggio di una bibliografia Lagrangiana. (Italian) [essay on a bibliography of Lagrangiana]. Isis, 40(2):112–117, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227038.

Anonymous:1949:Oa

[895] Anonymous. Obituaries. Isis, 40(2):118, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227039.

Mayer:1949:NC

[896] Claudius F. Mayer, G. Sarton, and E. J. Dijksterhuis. Notes and correspondence. Isis, 40(2):119–122, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227040.

Anonymous:1949:Pa

[897] Anonymous. Personalia. Isis, 40(2):122, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227041.

Graubard:1949:THS

[898] Mark Graubard. Teaching of the history of science. Isis, 40(2):123, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227042.

Sarton:1949:STC

[899] George Sarton and Frances Siegel. Seventy-third critical bibliography of the history and philosophy of science and of the history of civilization (to November 1948). Isis, 40(2):124–193, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227043.

Pogo:1949:AD

[900] Alexander Pogo. Administrative documents. Isis, 40(2):193–196, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227044.

Anonymous:1949:BMb

[901] Anonymous. Back matter. Isis, 40(2):??, May 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227045.

Anonymous:1949:FMc

[902] Anonymous. Front matter. Isis, 40(3):197–198, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227235.

Roberts:1949:WMP

[903] Leo Roberts. Wolfson’s monument to philo. Isis, 40(3):199–213, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227236.

Olmsted:1949:ATA

[904] John W. Olmsted. The “application” of telescopes to astronomical instruments, 1667–1669; a study in historical method. Isis, 40(3):213–225, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227237.

Landsberg:1949:PDP

[905] H. Landsberg. Prelude to the discovery of penicillin. Isis, 40(3):225– 227, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227238.

Sarton:1949:IWD

[906] George Sarton. Incunabula wrongly dated. Fifteen examples with eighteen illustrations. Isis, 40(3):227–240, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227239.

Neugebauer:1949:EHA

[907] O. Neugebauer. The early history of the astrolabe. Studies in ancient astronomy IX. Isis, 40(3):240–256, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227240.

Drabkin:1949:BRB

[908] I. E. Drabkin. Book review: Kommentar zum ersten Buch von Euklids “Elementen.” by Proklus Diadochus; Les Commentaires sur le premier livre des Éléments d’Euclide by Proclus De Lycie. Isis, 40(3):256–257, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227241.

Rosen:1949:BRB

[909] Edward Rosen. Book review: The Sphere of Sacrobosco and Its Commentators by Lynn Thorndike. Isis, 40(3):257–263, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227242.

Sarton:1949:BRBd

[910] George Sarton. Book review: Li livres dou tresor by Brunetto Latini; Francis J. Carmody. Isis, 40(3):263–265, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227243.

Clagett:1949:BRB

[911] Marshall Clagett. Book review: Nicolaus of Autrecourt, a Study in 14th Century Thought by Julius R. Weinberg. Isis, 40(3):265–269, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227244.

Struik:1949:BRB

[912] Dirk J. Struik. Book review: Simon Stevin by E. J. Dijksterhuis. Isis, 40(3):269–270, August 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227245.

Sarton:1949:BRBe

[913] George Sarton. Book review: Mersenne ou la naissance du mécanisme by Robert Lenoble. Isis, 40(3):270–272, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227246.

Graubard:1949:BRBb

[914] Mark Graubard. Book review: Christian Huygens and the Development of Science in the Seventeenth Century by A. E. Bell. Isis, 40(3):272– 273, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227247.

Smith:1949:BRB

[915] Henry Nash Smith. Book review: Pioneer Life in Kentucky by Daniel Drake; Emmet Field Horine. Isis, 40(3):273–274, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227248.

Graubard:1949:BRBc

[916] Mark Graubard. Book review: Mendeleyev, The Story of a Great Scientist by Daniel Q. Posin. Isis, 40(3):274–275, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227249.

Saunders:1949:BRBd

[917] J. B. DeC. M. Saunders. Book review: Sir Frederick Banting by Lloyd Stevenson. Isis, 40(3):275–277, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227250.

Sarton:1949:BRBf

[918] George Sarton. Book review: A Source Book in Greek Science by Morris R. Cohen; I. E. Drabkin. Isis, 40(3):277–278, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227251.

Sarton:1949:BRBg

[919] George Sarton. Book review: Euclide et Ptolémée. Deux stades de l’optique géométrique grecque by Albert Lejeune. Isis, 40(3):278, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227252.

Sarton:1949:BRBh

[920] George Sarton. Book review: Ancient Roman Construction in Italy from the Prehistoric Period to Augustus by Marion Elizabeth Blake; Esther Boise Van Deman. Isis, 40(3):279, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227253.

Sarton:1949:BRBi

[921] George Sarton. Book review: History of Chinese Society. Liao (907– 1125) by Karl A. Wittfogel; Fêng Chiashêng; John De Francis; Esther S. Goldfrank; Lea Kisselgoff; Karl H. Menges. Isis, 40(3):280–281, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227254.

Partington:1949:BRB

[922] J. R. Partington. Book review: The Chemical Arts of Old China by Li Ch’iao-P’ing. Isis, 40(3):281, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227255.

Leake:1949:BRBc

[923] Chauncey D. Leake. Book review: The Life of Science: Essays in the History of Civilization by George Sarton. Isis, 40(3):282–283, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227256.

Brown:1949:BRB

[924] Harcourt Brown. Book review: La Formation de l’Esprit scientifique by Gaston Bachelard. Isis, 40(3):283–285, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227257.

Thomas:1949:BRB

[925] Lewis V. Thomas. Book review: Tarih Boyunca Ilim ve Din. (Science and Religion in the Course of History) by Abdülhak Adnan-adivar. Isis, 40(3):285–286, August 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227258.

Frye:1949:BRB

[926] Richard N. Frye. Book review: Hayatta en Hakikî Mürs.it Ilimdir (The Truest Guide in Life Is Science) by Aydin Sayili. Isis, 40(3):286, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227259.

Graubard:1949:BRBd

[927] Mark Graubard. Book review: Voyages to the Moon by Marjorie Hope Nicholson. Isis, 40(3):286–287, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227260.

Boyer:1949:BRB

[928] Carl B. Boyer. Book review: A Concise History of Mathematics by Dirk J. Struik. Isis, 40(3):287–289, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227261.

Sarton:1949:BRBj

[929] George Sarton. Book review: Outline of the History of Mathematics by Raymond Clare Archibald. Isis, 40(3):289–290, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227262.

Zirkle:1949:BRBe

[930] Conway Zirkle. Book review: The Lungfish, the Dodo, and the Unicorn: An Excursion into Romantic Zoology by Willy Ley. Isis, 40(3):290, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227263.

Zirkle:1949:BRBf

[931] Conway Zirkle. Book review: The Fig by Ira J. Conduit. Isis, 40(3):290– 291, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227264.

Zirkle:1949:BRBg

[932] Conway Zirkle. Book review: The Romance of Coffee: An Outline History of Coffee and Coffee Drinking through a Thousand Years by William H. Ukers. Isis, 40(3):291–292, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227265.

Zirkle:1949:BRBh

[933] Conway Zirkle. Book review: Freedom from Want: A Survey of the Possibilities of Meeting the World’s Food Needs by E. E. DeTurk. Isis, 40(3): 292–293, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227266.

Archibald:1949:BRB

[934] R. C. Archibald. Book review: The Mathematical Basis of the Arts by Joseph Schillinger. Isis, 40(3):293–295, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227267.

Ulich:1949:BRB

[935] Robert Ulich. Book review: Histoire de L’Education dans L’Antiquité by H. I. Marrou. Isis, 40(3):295–296, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227268.

Dufrenoy:1949:BRB

[936] Marie-Louise Dufrenoy. Book review: Les Jésuites et l’éducation de la société française. La géographie des humanistes by François De Dainville. Isis, 40(3):296–297, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227269.

Lenzen:1949:BRBd

[937] V. F. Lenzen. Book review: Existence and Inquiry by Otis Lee. Isis, 40(3):297–298, August 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227270.

Jackman:1949:BRBb

[938] Sydney Jackman. Book review: The City of Reason by Samuel H. Beer. Isis, 40(3):298–299, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227271.

OMalley:1949:BRB

[939] Charles D. O’Malley. Book review: Music and Medicine by Dorothy M. Schullian; Max Schoen. Isis, 40(3):299, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227272.

Urdang:1949:BRB

[940] George Urdang. Book review: An Introduction to the History of Dentistry with Medical & Dental Chronology & Bibliographic Data. Vol. 1 by Bernhard Wolf Weinberger; An Introduction to the History of Dentistry in America. Washington’s need for Medical and Dental Care. Houdon’s Life Mask versus His Portraitures. Vol. 2 by Bernhard Wolf Weinberger. Isis, 40(3):299–301, August 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227273.

Drinker:1949:BRBa

[941] C. K. Drinker. Book review: Doctors of Infamy. The Story of the Nazi Medical Crimes by Alexander Mitscherlich; Fred Mielke; Heinz Norden. Isis, 40(3):301, August 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227274.

Drinker:1949:BRBb

[942] C. K. Drinker. Book review: Acclimatization in the Andes. Historical Confirmations of “Climatic Aggression” in the Development of Andean Man by Carlos Monge; Donald F. Brown. Isis, 40(3):302, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227275.

Drinker:1949:BRBc

[943] C. K. Drinker. Book review: Aviation Medicine in Its Preventive Aspects by John F. Fulton. Isis, 40(3):302–303, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227276.

Cohen:1949:BRBb

[944] I. Bernard Cohen. Book review: Literary History of the United States by Robert E. Spiller; Willard Thorp; Thomas H. Johnson; Henry Seidel Canby. Isis, 40(3):303–304, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227277.

Anonymous:1949:BMc

[945] Anonymous. Back matter. Isis, 40(3):305–308, August 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227278.

Anonymous:1949:FMd

[946] Anonymous. Front matter. Isis, 40(4):309–310, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226363.

Sarton:1949:TPV

[947] George Sarton. Third preface to volume forty: Qualifications of teachers of the history of science (second article). Isis, 40(4):311–313, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226364.

Miles:1949:NSE

[948] Wyndham Miles. Notes on some early chemistry books published in Pennsylvania. Isis, 40(4):313–316, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226365.

Galdston:1949:SFC

[949] Iago Galdston. Sigmund Freud: A critical summary and review. Isis, 40 (4):316–327, November 1949. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226366.

Patterson:1949:HGT

[950] Louise Diehl Patterson. Hooke’s gravitation theory and its influence on Newton. I: Hooke’s gravitation theory. Isis, 40(4):327–341, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226367.

Francon:1949:LMJ

[951] Marcel Françon. Le langage mathématique de J.-J. Rousseau. (French) [The mathematical language of Jean-Jacques Rousseau]. Isis, 40(4):341– 344, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226368.

Szczesniak:1949:NKB

[952] Boleslaw Szczesniak. Note on Kepler’s Tabulae Rudolphinae in the Library of Pei-t’ang in Pekin. Isis, 40(4):344–347, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226369.

Thorndike:1949:CJC

[953] Lynn Thorndike. Concerning John Canonicus. Isis, 40(4):347–349, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226370.

Thorndike:1949:NC

[954] Lynn Thorndike, René Taton, and Edwin G. Beal. Notes and correspondence. Isis, 40(4):349–353, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226371.

Anonymous:1949:Pb

[955] Anonymous. Personalia. Isis, 40(4):353–354, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226372.

Anonymous:1949:Ob

[956] Anonymous. Obituaries. Isis, 40(4):354–355, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226373.

Sarton:1949:SFC

[957] George Sarton and Frances Siegel. Seventy-fourth critical bibliography of the history and philosophy of science and of the history of civilization (to March 1949). Isis, 40(4):356–403, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226374.

Anonymous:1949:MHS

[958] Anonymous. 1949 Meeting of the History of Science Society: Boston, Massachusetts, 28–29 December. Isis, 40(4):403–404, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226375.

Anonymous:1949:SIC

[959] Anonymous. Sixth International Congress for the History of Science, Amsterdam, 14–20 August, 1950. Isis, 40(4):404, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226376.

Anonymous:1949:VI

[960] Anonymous. Volume information. Isis, 40(4):405–409, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226362.

Anonymous:1949:BMd

[961] Anonymous. Back matter. Isis, 40(4):410–412, November 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226377.